### v1

# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

Abstract.
Answering natural language (NL) questions about tables, which is referred to as Tabular Question Answering (TQA), is important because it enables users to extract meaningful insights quickly and efficiently from structured data, bridging the gap between human language and machine-readable formats. Many of these tables originate from web sources or real-world scenarios, necessitating careful data preparation (or data prep for short) to ensure accurate answers. However, unlike traditional data prep, question-aware data prep introduces new requirements, which include tasks such as column augmentation and filtering for given questions, and question-aware value normalization or conversion. Because each of the above tasks is unique, a single model (or agent) may not perform effectively across all scenarios. In this paper, we propose AutoPrep, a large language model (LLM)-based multi-agent framework that leverages the strengths of multiple agents, each specialized in a certain type of data prep, ensuring more accurate and contextually relevant responses. Given an NL question over a table, AutoPrep performs data prep through three key components. Planner: Determines a logical plan, outlining a sequence of high-level operations. Programmer: Translates this logical plan into a physical plan by generating the corresponding low-level code. Executor: Iteratively executes and debugs the generated code to ensure correct outcomes. To support this multi-agent framework, we design a novel Chain-of-Clauses reasoning mechanism for high-level operation suggestion, and a tool-augmented method for low-level code generation. Extensive experiments on real-world TQA datasets demonstrate that AutoPrep can significantly improve the SOTA TQA solutions through question-aware data preparation.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://anonymous.4open.science/r/AutoPrep-Ano.
1. Introduction
Tabular Question Answering (TQA) refers to the tasks of answering natural language (NL) questions based on provided data tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023). TQA empowers non-technical users such as domain scientists to easily analyze tabular data and thus has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022).
(a) WikiTQ
(b) TabFact
Figure 1. An error analysis of LLM-based TQA (using GPT-4) on two well-adopted datasets, with statistics calculated from a sample of 500 instances from each dataset.
(a) Question-Aware Column Augmentation: Necessary information for QA is not explicitly available in tables.
(b) Question-Aware Value Normalization: Heterogeneous data representations or formats in tables, and how to normalize is query-dependent.
(c) Query-Aware Column Filtering: Irrelevant columns increase complexity and cost of LLMs, which also lead to errors in TQA methods.
Figure 2. Examples of Question-Aware Data Preparation for TQA tasks.
Because TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024) naturally rely on large language models (LLMs). These solutions either utilize LLMs as black-boxes or leverage them for code generation. However, question-aware data preparation is not explicitly explored in these approaches, which limits the effectiveness of LLM-based methods.
Figure 1 presents the results of an error analysis of LLM-based code generation (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019). The analysis reveals that 84% and 76% of the errors stem from insufficient data prep at various granularities (i.e., table-level, column-level, and cell-level) in relation to the NL questions. A similar trend is observed when using LLMs as black-boxes, though those results are not displayed here due to space constraints. These findings underscore the necessity for meticulous data prep to ensure accurate answers.
To better understand these cases, let’s illustrate by examples.
Example 0.
Figure 2 shows several examples of question-aware data prep tasks that are needed for TQA.
(a) Question-Aware Column Augmentation. Figure 2(a) illustrates the necessity of augmenting the tables with new columns, as the information that NL questions require is not explicitly available. For example, to answer $Q_{1}$, a new Country column should be created to augment $T_{1}$, e.g., by extracting country information from Cyclist.
(b) Question-Aware Value Normalization: Figure 2(b) highlights the need for normalizing data types or formats in tables (e.g.,ITA” vs. “Italia” in $T_{4}$, and “19-Oct” vs. “9/14” in $T_{6}$). Note that normalizing tables in TQA requires converting data types or formats based on the need of questions, meaning different questions may result in various operations. For example, as $Q_{5}$ asks for summing numbers in $T_{5}$ but the corresponding columns are of string type, we need to remove commas from numbers and convert strings to integers.
(c) Question-Aware Column Filtering. Figure 2(c) highlights the importance of selecting question-relevant columns from tables, as irrelevant columns add complexity and can lead to errors in code generation. For example, the Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, which may mislead the LLM. Additionally, although $T_{8}$ contains 12 columns, only two are relevant to $Q_{8}$. Providing all columns to the LLM can thus result in long-context understanding challenges (Li et al., 2024).
Example 1.1 highlights that different types of data prep operations are needed to satisfy the need of different NL questions.
Given that data prep is a complex challenge without a one-size-fits-all solution, a sensible approach is thus to develop specialized methods w.r.t. each type of data prep operations. These specialized methods, such as data normalization, can involve either LLM-powered code generation or retrieving existing code from data prep libraries. An intelligent module can then automatically determine which operations to perform based on the properties of the question and the corresponding table.
AutoPrep: A Multi-Agent LLM Framework. To address the problem, we propose AutoPrep, a multi-agent LLM framework that automatically prepares tables based on the specific needs of NL questions, eliminating the need for users to specify data prep operations. The core idea is to decompose the data prep task in TQA into sub-tasks, leveraging multiple agents specialized in different areas to ensure accurate and contextually relevant responses.
Given an NL question over a table, AutoPrep performs data prep through the following three core components:
  • Planner: This component is responsible for analyzing the natural language question and the structure of the given table. It generates a logical plan that outlines the sequence of high-level operations required to prepare the data for answering the question. These operations may include tasks such as column augmentation, value normalization, or filtering irrelevant data.
  • Programmer: Once the logical plan is in place, the Programmer translates it into a physical plan by generating low-level executable code. This component is responsible for selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, customizing the code to match the table’s structure and data formats.
  • Executor: The Executor is tasked with running the generated code and ensuring its correct execution. It performs iterative execution, where each operation is applied sequentially to the table, and any issues or errors are detected and debugged in real-time.
Technical Challenges. Two key technical challenges arise in our proposed AutoPrep framework.
(C1) Planner: High-Level Operation Suggestion. Even for the same table, different NL questions may require not only distinct data prep operations but also varying sequences of those operations. Thus, suggesting a tailored sequence of high-level operations that meet the specific needs of these questions is a significant challenge.
(C2) Programmer: Low-Level Code Generation. A given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial. Generating optimized code for each operation that is customized for the tables presents a non-trivial task. For instance, consider the string-to-integer conversion in Example 2.4; failing to account for different value formats can lead to execution errors from a simple type conversion.
Key Ideas. To address the challenge (C1), we propose a Chain-of-Clauses (CoC) based reasoning method for the Planner agent. This method analyzes the NL question and outputs a sequence of high-level desired data prep operations, namely an Analysis Sketch. Based on the Analysis Sketch, the Planner assigns high-level operations to specialized agents for low-level implementation (i.e., code generation).
To tackle challenge (C2), we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions. We establish a search space of API functions for all high-level operations, design an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism.
Contributions. Our contributions are summarized as follows.
(1) We propose AutoPrep, an LLM-based multi-agent framework for automatic data prep in TQA. We formally define the problem and formalizes typical high-level data prep operations (Section 2), and introduce an overview of the AutoPrep framework (Section 3).
(2) We develop effective techniques for our AutoPrep framework. We propose a novel Chain-of-Clauses reasoning mechanism (Section 4) to suggest high-level operations and a tool-augmented method to generate low-level code (Section 5).
(3) We conduct a thorough evaluation on data prep in TQA (Section 6). Extensive experiments show that AutoPrep achieves new SOTA accuracy on well-adopted TQA datasets. We make the code used in our experiments available on an anonymous GitHub repo [^1] https://anonymous.4open.science/r/AutoPrep-Ano.
2. Preliminaries
This section first introduces Tabular Question Answering (TQA) in Section 2.1 and then formalizes the problem of question-aware data prep for TQA in Section 2.2. Then, we analyze the limitation of existing LLM-based methods in Section 2.3.
2.1. Tabular Question Answering
Let $Q$ be a natural language (NL) question, and $T$ a table consisting of $m$ columns (i.e., attributes) ${A_{1},A_{2},\ldots,A_{m}}$ and $n$ rows ${r_{1},r_{2},\ldots,r_{n}}$, where $v_{ij}$ denotes the value in the $i$-th row and $j$-th column of the table. The problem of tabular question answering (TQA) is to generate an answer, denoted as $A$, in response to question $Q$ based on the information in table $T$. By the purposes of the questions, there are two main types of TQA problems: (1) table-based fact verification (Chen et al., 2019; Aly et al., 2021), which determines whether $Q$ can be entailed or refuted by $T$, and (2) table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022), which extracts or reasons the answer to $Q$ from $T$.
Example 0.
Figure 2 provides several examples of TQA. For instance, consider table $T_{1}$, which contains medal information for cyclists from different countries, with two columns: Cyclist and Medal. Given the question $Q_{1}$, “Which country has the most medals?”, the answer $A_{1}$ should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”.
Note that TQA often requires complex reasoning over tables. For instance, to answer question $Q_{2}$, we first need to calculate the “GDP growth rate” for all countries, then sort the countries by growth rate, and finally identify the country with the highest GDP growth rate, i.e., CHN.
2.2. Data Prep for TQA
Unlike traditional data prep, for question-aware data prep for TQA, we need to tailor table $T$ to the specific needs of question $Q$.
Consider Example 2.1: careful data prep, such as extracting countries into a new column and then normalizing them as ISO codes, is crucial for generating an accurate answer. Moreover, it is important to note that different questions may require different data prep tasks, even for the same table. For instance, country extraction may not be necessary if the question is asking which cyclist, instead of country, won the most medals.
Data Prep Operations. To meet the new requirements of data prep for TQA, this paper defines high-level data prep operations (or operations for short) to formalize the question-aware data prep tasks, such as column augmentation, question-aware value normalization or conversion, and column filtering. Formally, an operation, denoted as $o$, encapsulates a specific question-aware data prep task that transforms table $T$ into another table $T^{\prime}$, i.e., $T^{\prime}=o(T)$. Specifically, this paper considers the following three types of operations.
  • Augment: a data prep task that augments a new column for table $T$ based on existing columns, to better answer $Q$. This task typically involves column combinations via arithmetic operations, value extraction and inference, etc.
  • Normalize: a data prep task that normalize types or formats of the values in a column of $T$ based on the needs of $Q$. This task typically involves value representation or format normalization, type conversion, etc.
  • Filter: A data prep task that filters out columns in $T$ that are not relevant to answering question $Q$. This is crucial for handling large tables to address the input token limitations and challenges in long-context understanding of LLMs (Li et al., 2024).
Given a high-level operation $o_{i}$, we define $f_{i}$ as its low-level implementation, either by calling a well-established algorithm from a known Python library or using a customized Python program to meet the requirements of $o_{i}$. For example, to implement the Augment operation $o_{1}$ that augments a Country column in $T_{1}$, a possible implementation $f_{i}$ is to write a customized program by using a specific regular expression as shown in Figure 2(a).
Example 0.
Figure 2 shows examples of question-aware data prep operations for TQA, along with their implementations in Python.
(a) The Augment operation: Figure 2(a) shows three examples of Augment operations, i.e., extracting Country information from column Cyclist in $T_{1}$, computing new column GrowthRate using two columns in $T_{2}$, and inferring a status of IfLost by analyzing the scores in column Result. Note that the corresponding Python implementations are provided alongside the operations.
(b) The Normalization operation: Figure 2(b) shows three examples of Normalize operations, i.e., normalizing country formats in $T_{4}$, converting strings into integers in $T_{5}$, and normalizing value formats in Date in $T_{6}$. We can see that, to implement a specific operation (e.g., in $T_{4}$), we sometimes need to call external tools, like a function clean_country from an external library.
(c) The Filter operation: Figure 2(c) illustrates two examples of Filter operations in tables $T_{5}$ and $T_{6}$. The Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, potentially misleading the LLM. Additionally, although $T_{8}$ has 12 columns, only two are relevant to $Q_{8}$, and providing all columns may cause long-context understanding challenges of LLMs (Li et al., 2024).
Note that the above three types of operations address most of the errors caused by insufficient data prep, as shown in our error analysis in Figure 1. In fact, the set of data prep operations is extensible, in order to support new or domain-specific data prep tasks.
Question-Aware Data Prep for TQA. Given an NL question $Q$ posed over table $T$, the problem of question-aware data prep for TQA is to generate a sequence of high-level operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$ as a logical plan. Then, it generates a physical plan, where each operation $o_{i}$ is implemented by low-level code $f_{i}$, such that these operations transform $T$ into a new table $T^{\prime}$ that meets the needs of $Q$.
Example 0.
To better answer the example question $Q_{1}$ over $T_{1}$ in Figure 2(a), we first generate an Augment operation to transform $T_{1}$ into $T^{\prime}_{1}$ with three columns: Cyclist, Medal, and Country. Next, we generate a Normalize operation, similar to the one in $T_{4}$, to standardize the values in Country into a unified format, producing table $T^{\prime\prime}_{1}$. Finally, we leverage LLMs, either as black-box or for code generation, to compute the answer $A_{1}$ to $Q_{1}$ from the prepared table $T^{\prime\prime}_{1}$.
2.3. Limitations of Existing LLM-based Solutions
Naturally, LLMs can be utilized to support question-aware data prep for TQA. A straightforward way is to prompt an LLM with a task description and several demonstrations (i.e., examples), and requires the LLM to generate data prep code, as shown in Figure 3.
However, as shown in Figure 2, the diversity of NL questions necessitates distinct data prep operations. Next, let’s see how a single LLM is used to solve this problem, by an example.
Example 0.
Consider the example table in Figure 3 with columns Date, Cyclist, Medal and Age, along with question: “Which country has the most medals in total in February”. When using few-shot prompting with a task description and a few demonstrations, this strategy may face limitations in both stages of data prep in TQA.
First, as the number of demonstrations is limited, LLMs may struggle to decide which data prep operations are needed for the NL questions, leading to both false negatives (FNs) and false positives (FPs). For example, numeralizing Age in code block ② is an FP, as it is irrelevant to the question. Conversely, the LLM fails to normalize the Date column before extracting the month in code block ③, which is an FN.
Second, due to input token limitations and challenges in long-context understanding (Li et al., 2024), a single LLM may struggle to generate fine-grained code for each operation. For instance, normalizing Medal in code block ③ does not handle all edge cases (e.g., ‘2*’), and the country extraction in code block ⑤ fails to correctly process ‘Dav.ITA’, which is formatted differently from other values.
Example 2.4 shows that, because each of the above tasks is unique, a single model may not perform well across all scenarios. Thus, given that data prep is a complex challenge without a one-size-fits-all solution, a more effective approach is to develop specialized methods for each type. A centralized module can then determine which operations to perform and assign them accordingly.
Figure 3. An LLM-based solution that prompts an LLM with a task description and several demonstrations (i.e., exemplars), and requires the LLM to generate data prep code.
3. An Overview of AutoPrep
Figure 4. Our proposed Multi-agent Data Prep Framework for TQA (Left), which could be plugged to any other TQA method (Right), such as NL2SQL as shown in the Figure. The proposed framework AutoPrep decomposes the data prep process into two stages: ① The Planning stage and ② the Executing stage.
To address the limitations, we propose AutoPrep, a multi-agent LLM framework that automatically prepares tables for given NL questions. Figure 4 provides an overview of our framework. Given an NL question $Q$ posed over table $T$, AutoPrep decomposes the data prep process into three stages:
  1. (1)
    [Uncaptioned image] Planner Agent: the Planning stage. It guides the LLM to suggest high-level data prep operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$, which are tailored to specific question $Q$,
  2. (2)
    Multiple Programmer Agents (i.e., [Uncaptioned image] Augmenter): the Programming stage. It directs the LLM to generate low-level implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$.
  3. (3)
    An [Uncaptioned image] Executor Agent: the Executing stage. It executes the given low-level code and debugs iteratively to improve the generated code quality.
After that, an [Uncaptioned image] Analyzer agent extracts the answer from the prepared table. This agent can either use LLMs as black-boxes or leverage them for code generation, which is orthogonal to the question-aware data prep problem studied in this paper. For simplicity, we use a Text-to-SQL strategy that translates the question into an SQL query over the prepared table to obtain the final answer, as shown in Figure 4. Note that other strategies could also be used by the agent in a “plug-and-play” manner, which will be discussed in the experiments.
Example 0.
Figure 4 illustrates how AutoPrep supports data prep for question “Which country has the most medals in total in February?” posed over a table with columns Date, Cyclist, Medal and Age.
(a) The Planning stage: The [Uncaptioned image] Planner suggests the following high-level operations to address the specific NL question:
  • $T^{\prime}=$Filter($T$, [Date, Cyclist, Medal]) that filters out column Age, which is irrelevant to the question;
  • $T^{\prime}$[Country]=Augment(“Extract country code”, Cyclist) that extracts the country information from column Cyclist, producing a new Country column, in response to the “which country” part of the question.
  • Normalize(“Case to INT”, Medal) that standardizes the value formats in the Medal column (e.g., removing quotation marks and asterisks) and then converts the strings to integers, as the question requires aggregating the “medals in total”;
  • Normalize(“Format date as %m-%d”, Date) standardizes the values in the Date column into a unified format to support the ‘in February’ condition in the question.
(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e.,, [Uncaptioned image] Augment, [Uncaptioned image] Normalizer and [Uncaptioned image] Filter, corresponding to the Augment, Normalize and Filter operations, respectively. Each specialized Programmer focuses on generating executable code for its assigned operations.
(c) The Executing stage: a Executor agent iteratively refines the generated code if any errors occur.
After these stages, AutoPrep generates a prepared table $T^{*}$, which is then fed into an [Uncaptioned image] Analyzer agent to produce the answer ITA.
The Planner agent and the three Programmer agents (i.e., Augmenter, Normalizer, and Filter) are the core modules of AutoPrep, which will be discussed in more details below.
The Planner Agent. Given a question $Q$ and a table, the Planner agent generates a sequence of high-level operations for the specific question, as illustrated in the previous example. The key challenge here is how to suggest question-aware operations that address specific NL questions. Specifically, even for the same table, different NL questions may require not only different data prep operations but also varying sequences of those operations. For example, the second Normalize operation might not be necessary if “in February” is not part of the question. To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent. This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch. Compared to existing CoT methods (Wei et al., 2022), which break down questions into sub-questions, our approach better captures the relationships between questions and tables. More details of the CoC method are given in Section 4.
The Programmer Agents. Given a high-level operation $o_{i}$, a Programmer agent is responsible for generating low-level code $f_{i}$ to execute the operation. As discussed previous, we develop specialized agents i.e., Augmenter, Normalizer, and Filter for various types of operations. The key challenge is that a given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial. For example, the Augmenter agent may generate an overly generic regular expression that extracts countries based on parentheses. Unfortunately, this code fails to correctly process “Dav. ITA”, which is formatted differently from other values. To tackle this challenge, we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions. We establish a search space of API functions for all high-level operations, create an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism. More details of our tool-augmented approach are discussed in Section 5.
Note that our proposed AutoPrep framework is extensible. When additional question-aware data prep operations are required, more specialized Programmer agents can be designed to handle them. The central Planner agent can then determine which operations should be performed and assign them accordingly.
4. The Planner Agent
Given a question $Q$ posed over a table $T$, Planner aims to generate a logical plan that outlines a sequence of high-level operations needed to prepare the table for answering the question. In this section, we first introduce a direct prompting method in Section 4.1 that better demonstrates the key challenge in designing the Planner agent. Then, to address the challenge, we propose our Chain-of-Clauses (CoC) reasoning method in Section 4.2.
Figure 5. Our proposed approaches to logical plan generation in the Planner agent: (a) A straightforward direct prompting method, and (b) a more effective Chain-of-Clauses (CoC) reasoning method.
4.1. A Direct Prompting Method
The most common way to generate a logical plan is to directly prompt an LLM using a typical in-context learning approach, as shown in Algorithm 1. The inputs are a question $Q$, a table $T$, a set $\Sigma$ of specifications for each operation type, and an LLM $\theta$. Here, each specification $\sigma\in\Sigma$ describes the purpose of an operation type, e.g., “an Augment operation augments a new column for a table based on existing columns, in response to the specific needs of a question”. The output of the algorithm is a set $O$ of high-level operations like those shown in Figure 4, e.g., Normalize(“Cast to INT”, Medal).
Algorithm 1 Direct Prompting for the Planner Agent
1:A question $Q$, a table $T$, an operation specification set $\Sigma$, and a large language model (LLM) $\theta$.
2:// Prepare $k$ in-context exemplars (i.e., demos)
3:$\mathcal{D}=\{(T_{i},Q_{i},\Sigma,O_{i})\}_{i=1}^{k}$
4:// Prompt the LLM for inferring the operations
5:$O\leftarrow\theta(T,Q,\Sigma,\mathcal{D})$
6:return $O$
Specifically, Algorithm 1 first prepares a set $\mathcal{D}$ of $k$ exemplars $\{(T_{i},Q_{i},\Sigma,O_{i})\}_{i=1}^{k}$ that demonstrates to LLM $\theta$ the goal of this logical plan generation task. Then, the algorithm prompts the LLM to infer the operations $O$ for our input table $T$ and question $Q$.
Example 0.
Figure 5(a) shows an example of the aforementioned direct prompting method, which produces two high-level operations, Filter and Normalize. However, this logical plan might not be accurate as discussed as follows.
(a) Incorrect operations: The Filter operation incorrectly filters out the Cyclist column because the column is not explicitly mentioned in the question. This operation is clearly disastrous, as the country information is implicitly embedded in the Cyclist column.
(b) Missing operations: Observing the ground-truth in Figure 4, we see that the Augment operation on Cyclist is not generated, as the column has already been filtered out. Additionally, although the Normalize operation on Medal is generated, the Normalize operation on Date is missing. This is because the phrase “the most” in the question suggests the need for type conversion of Medal, but there are no clues from the question to normalize Date, unless we observe from the table that the values are inconsistent.
Key Challenge. The above example clearly demonstrates that the key challenge in designing the Planner agent, as discussed earlier, is to capture the relationships between different parts of question $Q$ and the columns in table $T$. For instance, as shown in Figure 4, the Normalize operation on the Date column is generated by considering two key factors: (1) the question contains “in February”, and (2) the values in the Date column are inconsistent.
4.2. A Chain-of-Clauses Method
To address this challenge, we propose a Chain-of-Clauses (CoC) reasoning method the decomposes the entire process of logical plan generation into two phrases, as shown in Figure 5(b).
  • The first phase leverages an LLM to generate an SQL-like Analysis Sketch, outlining how the table should be transformed to produce the answer.
  • The second phase iteratively examines different clauses in the Analysis Sketch, e.g., Date LIKE02-%’. For each clause, we associate the corresponding data in $T$ (e.g., values in column Date) with it to prompt the LLM for generating possible high-level operations (e.g., Normalize).
Compared with existing CoT methods (Wei et al., 2022), which simply break down questions into sub-questions, our approach is more effective for logical plan generation due to the following reasons. First, our method decomposes the question into a set of analysis steps over the table, formalized as an SQL-like Analysis Sketch, which simplifies the task of logical plan generation. More importantly, our method jointly considers each analysis step along with the corresponding relevant data to prompt the LLM, effectively capturing the relationships between the question and the table.
Algorithm 2 CoC Reasoning for the Planner Agent
1:A question $Q$, a table $T$, an operation specification set $\Sigma$, and a large language model (LLM) $\theta$.
2:$O\leftarrow\emptyset$
3:// Phase I: Analysis Sketch generation
4:$S\leftarrow\theta(Q,T,\{(Q_{i},T_{i},s_{i})\})$
5:// Phase II: High-level operation generation
6:for each clause $s$ in $S$ do
7:    $T^{s}\leftarrow$ extract_data($T$)
8:    $o\leftarrow\theta(s,T^{s},\Sigma,\{s_{i},T^{s}_{i},\Sigma,o_{i}\})$
9:    $O\leftarrow O\cup\{o\}$
10:end for
11:return $O$
Specifically, Algorithm 2 presents the pseudo-code of our CoC reasoning method, which takes as input a question $Q$, a table $T$, a specification set $\Sigma$, and an LLM $\theta$ to produce an operation set $O$. To this end, we introduce a key concept of SQL-like Analysis Sketch, which is formally defined as follows.
SQL-like Analysis Sketch. An Analysis Sketch $S$ is an SQL-like statement, which can be represented as follows.
SELECT $A~{}|~{}{\tt agg}(A)~{}|~{}f(\{A_{i}\})$ FROM $T$
WHERE ${\tt Pred}(A_{i})$ AND $\ldots$ AND ${\tt Pred}(A_{j})$
GROUP BY $A$ ORDER BY $A$ LIMIT $n$
where ${\tt agg}(A)$ is an aggregation function (e.g., SUM and AVG) over column $A$, ${\tt Pred}(A_{i})$ denotes a predicate (i.e., filtering condition) over column $A_{i}$, such as Date LIKE02-%. Note that $f(\{A_{i}\})$ is a user-defined function (UDF) that maps existing columns $\{A_{i}\}$ in the table into a new column. Figure 5(b) shows an example Analysis Sketch with a UDF function that specifies a new column Country, i.e., $f(T)$ AS Country. Obviously, this UDF is introduced because the Analysis Sketch contains a GROUP BY Country clause.
Given the key concept Analysis Sketch, Algorithm 2 produces a logical plan via two phases, as shown in the following example.
Example 0.
Figure 5(b) shows our proposed CoC reasoning method for the example question and table in Figure 4. Specifically, the method generate a set $O$ of operations via the following two phases.
(a) Phase I - Analysis Sketch Generation: In this phase, the algorithm prompts the LLM $\theta$ with several exemplars $\{(Q_{i},T_{i},s_{i})\}$ to generate an Analysis Sketch $S$, as shown in Figure 5(b).
(b) Phase II - Operation Generation: In this phase, the algorithm iteratively examines the clauses in Analysis Sketch $S$ as follows.
  • $f(T)$ AS Country: this clause and the entire table $T$ are used to prompt the LLM to generate an Augment operation.
  • ${\tt SUM}({\tt Medal})$: this clause and the values in column Medal are used to prompt the LLM to generate a Normalize operation.
  • Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize operation.
Finally, the algorithm prompt the LLM to generate a Filter operation that only selects the columns relevant to the Analysis Sketch.
5. The Programmer & Executor Agents
Figure 6. Our proposed tool-augmented method for physical plan generation and execution.
The Programmer agents translate a high-level logical plan into a physical plan by generating low-level code, which is then passed to an Executor agent for iterative code execution and debugging. A straightforward approach is to prompt an LLM with the logical plan using typical in-context learning with several exemplars, asking the LLM to generate code for each high-level operation in the plan. The code is then executed iteratively, and if any runtime errors occur, the Debugger is prompted with the error messages for debugging.
However, the above method may have limitations, as the generated code may be overly generic and unable to effectively address the heterogeneity challenge of the tables. Specifically, many tables originate from web sources or real-world scenarios, where values have diverse syntactic (e.g., “19-Oct” and “9/14” in $T_{6}$) or semantic formats (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code customized to these tables. For example, consider table $T_{4}$ in 2(b). Given a Normalize operation to standardize country formats, a generic function from an existing Python library may not be sufficient to transform country names (e.g., “Italia”) to their ISO codes (e.g., “ITA”). In such cases, customized functions, such as the clean_country function from an external library dataprep (Peng et al., 2021), are needed to address specific requirements of code generation.
To address these limitations, we introduce a tool-augmented method for the Programmer agents, enhancing the LLM’s code generation capabilities by utilizing pre-defined API functions, referred to as the function pool in this paper. Specifically, we first introduce the key idea of our tool-augmented method in  5.1, then present the function pool in Section 5.2, and finally develop an algorithm for code generation and execution in Section 5.3.
5.1. Tool-Augmented Method: Key Idea
Figure 6 provides an overview of our tool-augmented method for physical plan generation and execution. Given a table $T$ and a set $O={o_{1},\ldots,o_{|O|}}$ of high-level operations, the method iteratively generates and executes low-level code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$.
Specifically, in the $i$-th iteration, given a high-level operation $o_{i}$ (e.g., Augment) and an intermediate table $T_{i}$, the method generates low-level code and executes it to produce an updated table $T_{i+1}$ through the following two steps.
Function Selection and Argument Inference. The method first prompts the LLM to select a specific function from a function pool $\mathcal{F}$ corresponding to the operation type (e.g., $F_{\tt Aug}$) and infer the arguments (e.g., regular expression) for the selected function.
For instance, given the Augment operation over table $T_{i}$ shown in Figure 6, the LLM selects a function from the pool $\mathcal{F}_{\tt Aug}$ designed specifically for Augment, obtaining an extract function with two arguments: column and func. The LLM then generates preliminary code for these arguments, assigning Cyclist to the column argument and generating a lambda function with a specific regular expression for argument func.
Note that, in addition to selecting functions from the corresponding pool, our method may also prompt the LLM to write specific code for an operation if no existing function is suitable to meet the operation’s requirements.
Code Execution and Debugging. Given the function generated in the previous step, the Executor agent then applies the function over table $T_{i}$ and support iterative debugging if there are any runtime errors or incorrect results.
For instance, as shown in Figure 6, the first lambda function for the argument func only extracts countries based on parentheses, which may produce incorrect results for the value “Dav.ITA”, as it is formatted differently from other values. In this case, based on the execution results, the Executor agent will refine the regular expression in the lambda function until no runtime errors or incorrect results occur.
After completing these two steps, the method produces a new intermediate table $T_{i+1}$ and proceeds to the next operation, $o_{i+1}$, i.e., the Normalize operation, as shown in Figure 6.
5.2. Search Space of Function Pools
This section presents a search of function pools for the different types of high-level operations formalized in the current AutoPrep.
Function pool for $\mathtt{Augment}$. We define a specific Programmer agent, i.e., Augmenter, to generate low-level code for the operation type $\mathtt{Augment}$ from the following function pool $\mathcal{F}_{\mathtt{Aug}}$.
  • $\mathtt{extract}$. This function extracts a substring from a column to generate a new column. It has two arguments: ${\tt column}$, representing the name of the source column, and ${\tt func}$, representing the substring extraction function. Using table $T_{1}$ from Figure 2(a) as an example, to extract the country code, we would set ${\tt column}$ to Cyclist and ${\tt func}$ to the lambda function lambda x: re.search(r’(.?)(.*?)( . ∗ ? )’,x).group(1).
  • $\mathtt{calculate}$. This function generates a new column through arithmetic operations of columns such as addition, subtraction, multiplication, or a combination of these. To use this operator, we need to provide two arguments: ${\tt columns}$ and ${\tt func}$. The argument ${\tt columns}$ should be a list of source column names. Similar to $\mathtt{extract}$, the ${\tt func}$ argument is typically a lambda function. However, for $\mathtt{calculate}$, the input to the function should be a dictionary of values rather than a string, since we may perform calculations on multiple columns. For example, given a table with two columns A and B, representing the scores of team A and team B, we need to generate a new column ScoreDifference. In this case, we set ${\tt columns}$ to [’A’, ’B’] and ${\tt func}$ to the specific lambda function lambda x: x[’A’]-x[’B’].
  • $\mathtt{map\mbox{-}to\mbox{-}boolean}$ and $\mathtt{concatenate}$. Both functions generate a new column from multiple columns. The difference is that one generates boolean values, while the other concatenates new strings. Both operators take two arguments, ${\tt columns}$ and ${\tt func}$, similar to those in the $\mathtt{calculate}$ operator.
  • $\mathtt{infer}$. This function directly uses LLMs to deduce values that could not be processed by the functions above. It takes a list of source columns and the failed target values as input, learning from demonstrations of successfully processed values to directly output the target values.
Function pool for $\mathtt{Normalize}$. We define a specific Programmer agent, i.e., Normalizer, to generate low-level code for the operation type $\mathtt{Normalize}$ from the following function pool $\mathcal{F}_{\mathtt{Norm}}$.
  • $\mathtt{to\mbox{-}numerical}$. This function standardizes a column to a numerical type, such as int or float. To use it, two arguments need to be specified: ${\tt column}$, representing the source column, and ${\tt func}$, representing a lambda function. For example, given the table $T_{5}$ in Figure 2(b), we use this function to standardize the Females column to integers, with ${\tt column}$ set to Females and ${\tt func}$ as lambda x: int(x.replace(’,’, ’’)).
  • $\mathtt{format\mbox{-}datetime}$. The function standardizes the format of columns with DateTime values. It has two arguments: ${\tt column}$ and ${\tt format}$. The ${\tt format}$ argument specifies the desired format for standardization. Given the table $T_{6}$ in Figure 2(b), to make the values in the Date column comparable, we use the $\mathtt{format\mbox{-}datetime}$ function with ${\tt column}$ set to Date and ${\tt format}$ set to “%m-%d”.
  • $\mathtt{clean\mbox{-}string}$. To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$. It has two arguments: ${\tt column}$ and ${\tt trans_{d}ict}$. The $\tt trans_{d}ict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$. Given the original table $T_{4}$ in Figure  2(b), we set $trans_{d}ict$ to {’Italia’: ’ITA’}.
  • $\mathtt{infer}$. Similar to the one for the $\mathtt{Augment}$ operation, this function uses LLMs to deduce values that could not be processed by the functions above.
Function pool for $\mathtt{Filter}$. We design the Normalizer agent to utilize a function pool $\mathcal{F}_{\mathtt{Filter}}$ with a single function, $\mathtt{filter\mbox{-}columns}$. This function has one argument, ${\tt rel\_columns}$, which represents a list of question-related column names.
Discussions. It is important to note that the search space is extensible, allowing additional functions or external APIs to be easily incorporated into the function pools. Moreover, AutoPrep can also include external API functions, such as specific data prep functions for various semantic types (e.g., date, name, company) developed in the dataprep library (Peng et al., 2021).
5.3. Code Generation and Execution
Algorithm 3 Low-level Code Generation and Execution
1:A question $Q$, a table $T$, a number $r$, a high-level operation set $O$ and three function pools $\mathcal{F}_{\tt Aug}$, $\mathcal{F}_{\tt Norm}$ and $\mathcal{F}_{\tt Filter}$, an LLM $\theta$
2:$F\leftarrow\varnothing,~{}T_{1}\leftarrow T$
3:for each operation $o_{i}$ in $O$ do
4:    // Step 1: Function selection and argument inference
5:    $\mathcal{F}\leftarrow$ Assign($o_{i}$, [$\mathcal{F}_{\tt Aug}$, $\mathcal{F}_{\tt Norm}$, $\mathcal{F}_{\tt Filter}$])
6:    $f_{i}\leftarrow\theta(o_{i},Q,T,\mathcal{F})$
7:    // Step 2: Code generation and execution
8:    while $f_{i}$ has errors &\&& ${\tt debug\_round}<r$ do
9:         $f_{i}\leftarrow$ Debug($Q,T,f_{i}$)
[…]
11:end for
12:return $T_{|O|+1}$
Algorithm 3 presents the pseudo-code of our tool-augmented approach for code generation and execution. The algorithm takes as input a question $Q$, a table $T$, a set $O$ of high-level operations, three function pools, and an LLM, and produces a prepared table $T_{|O|+1}$. The algorithm iteratively examines the operations in $O$. In each iteration, it first generates a function from the corresponding function pool and refines it through multiple rounds of debugging. The algorithm then applies the generated function to the corresponding intermediate table and proceeds to the next iterations.
6. Experiments
6.1. Experimental Setup
Table 1. Statistics of Datasets.
Dataset # Num. Row. Col. Ans.
WikiTQ $22,033$ 4$\sim$753 3$\sim$25 string / list (3.05%)
TabFact $2,024$ 5$\sim$47 5$\sim$14 true / false (49.60%)
Datasets. We evaluate our proposed framework AutoPrep using two well-adopted benchmarking datasets WikiTQ (Pasupat and Liang, 2015) and TabFact (Chen et al., 2019). The statistics are shown in Table 1.
  • WikiTQ. WikiTableQuestions (WikiTQ) is a dataset designed for QA tasks, containing complex questions annotated by crowd workers based on diverse Wikipedia tables. WikiTQ comprises 17,689 question-answer pairs in the training set and 4,344 in the test set with significant variation in table size. The answers in WikiTQ can be categorized into three forms: (i) string from a cell in the table, (ii) string from analyzing, and (iii) list of cells in the table.
  • TabFact. Table-Fact-Checking (TabFact) dataset provides a collection of Wikipedia tables, along with manually annotated NL statements. One requires to deduce the relations between statements and tables as “true” (a statement is entailed by a table) or “false” (a statement if refuted by a table). To reduce the experimental cost without losing generality, we choose the small test set provided by (Khot et al., 2022) which contains 2024 table-statement pairs. The tables in TabFact are much smaller than those in WikiTQ. The answers are binary with nearly equal proportions.
Evaluation Metrics. We adopt the evaluator in Binder (Cheng et al., 2023) to avoid the situation where program executions are likely to be semantically correct but fail to match the gold answer exactly. For example, when the SQL outputs 1/0 for yes/no questions, the Binder evaluator will conduct a pre-matching check and translate it to boolean values first. It will avoid mismatches of correct prediction and golden answer pairs like 7.45 and 7.4499, 4 years and 4, etc. Afterwards, we adopt accuracy as our evaluation metric.
Baselines. We evaluate AutoPrep using two categories of methods: (a) TQA w/o Data Prep: Zero-Shot QA and Few-Shot QA (Chen, 2022), as well as Text-to-SQL (Rajkumar et al., 2022); (b) TQA with Data Prep: Binder (Cheng et al., 2023) and ReAcTable (Zhang et al., 2024). We further describe them below.
  • Zero-Shot and Few-Shot QA (Chen, 2022). Zero-Shot QA directly asks an LLM to generate the answer based on the serialized table and question. To improve the performance, Few-Shot QA further provides several demonstrations to enhance the In Context Learning ability of the LLM. We implement these two methods with the prompt provided by (Cheng et al., 2023).
  • Text-to-SQL (Rajkumar et al., 2022). Text-to-SQL is a natural code generation method for the TQA tasks. Consider a serialized table and a question, it first generates the SQL program and then executes it to get the final answer from the table for the question.
  • Binder (Cheng et al., 2023). Binder further enhances the Text-to-SQL method by integrating LLMs into SQL programs. It uses LLMs to incorporate external knowledge bases and directly answer questions that are difficult to resolve using SQL alone. We utilize the code provided by (cod, 2023).
  • ReAcTable (Zhang et al., 2024). ReAcTable uses the ReAct paradigm to extract relevant data from the table using Python or SQL code generated by LLMs. Once all relevant data is gathered, it asks the LLMs to predict the answer. We run the original code from (cod, 2024) and kept all settings as default.
Implementation Details. We mainly use DeepSeek (Guo et al., 2024) API [^2] https://platform.deepseek.com/usage model as the foundation model for each LLM-based agent. We also explore other proprietary LLMs, such as GPT-3.5 (Brown et al., 2020), using the GPT-3.5-turbo-0613 version, as adopted by (Wang et al., 2024b). To ensure reproducibility, we set the temperature parameter for all agents to 0.01.
6.2. Data Prep Method Comparison
Table 2. Comparison of AutoPrep with data prep methods for TQA (the best results are bolded, and the second-best results are underlined).
Method DeepSeek GPT3.5
WikiTQ TabFact WikiTQ TabFact
Binder $56.98$ $79.64$ $\underline{56.74}$ $\underline{79.17}$
ReAcTable $\underline{64.25}$- $51.80$ $72.80$
AutoPrep $\mathbf{65.45}$ $\mathbf{86.41}$ (87.01) $\mathbf{63.86}$ $\mathbf{84.29}$
Exp-1: Comparison of AutoPrep with other data prep methods for TQA. We conduct experiments to evaluate AutoPrep against current state-of-the-art data prep methods for TQA on two datasets: WikiTQ and TabFact. Specifically, we compare AutoPrep with Binder and ReAcTable, presenting the results in Table 2. Notably, due to the lack of public scripts for TabFact from ReAcTable (Zhang et al., 2024), we report results only for DeepSeek on WikiTQ and use GPT-3.5 results from their original paper.
The results clearly demonstrate that AutoPrep achieves state-of-the-art performance on both the WikiTQ and TabFact datasets when using DeepSeek and GPT-3.5 as the foundation model. In particular, AutoPrep outperforms Binder by 14.86% on WikiTQ and 8.50% on TabFact with DeepSeek. When using GPT-3.5, the performance improvements are 12.55% and 6.47%, respectively. Furthermore, when compared with ReAcTable, AutoPrep enhances performance by 1.87% on WikiTQ using DeepSeek. Remarkably, with GPT-3.5, AutoPrep surpasses ReAcTable by 23.28% and 15.78% on WikiTQ and TabFact, respectively.
These improvements are attributed to AutoPrep’s comprehensive data prep operations, which are not fully addressed by Binder and ReAcTable. While Binder integrates LLMs to generate new columns to supplement external knowledge, it neglects critical aspects such as column filering and value normalization. In contrast, ReAcTable employs the ReAct data processing paradigm but lacks a specific decomposition of data prep tasks. Our experimental results highlight the importance of thorough data prep in enhancing TQA performance, which is often a significant bottleneck for existing methods.
Additionally, as shown in Table 2, DeepSeek consistently outperforms GPT-3.5 while incurring lower API costs [^3] https://platform.deepseek.com/api-docs/zh-cn/quick_start/pricing/ https://openai.com/api/pricing/. Therefore, we have selected DeepSeek as our default foundation model for subsequent experiments.
6.3. Improvement of Data Prep for TQA
Table 3. Impact of data prep on TQA Methods.
Method  WikiTQ  TabFact
Zero-Shot QA $17.66$ $76.78$
    + AutoPrep $28.22_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 10.56}}$ $78.70_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 1.92}}$
Few-Shot QA $56.05$ $81.97$
    + AutoPrep $56.33_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 0.28}}$ $82.70_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 0.73}}$
Text-to-SQL $51.66$ $68.53$
    + AutoPrep $\mathbf{65.45_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 13.79}}}$ $\mathbf{86.41_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 17.88}}}$
(a) Zero-Shot QA.
(b) Few-Shot QA.
(c) Text-to-SQL.
Figure 7. Performance on Small, Medium, Large tables from WikiTQ without/with data prep of AutoPrep.
Exp-2: Impact of data prep on TQA performance. We evaluate the performance of three baseline TQA methods, namely Zero-Shot QA, Few-Shot QA, and Text-to-SQL with and without the data prep provided by AutoPrep. The results are detailed in Table 3.
As demonstrated, integrating AutoPrep significantly improves the performance of all evaluated methods. Notably, Text-to-SQL experiences the most substantial enhancement, with performance increases of 26.69% on WikiTQ and 26.09% on TabFact. This significant improvement is attributed to Text-to-SQL’s sensitivity to data incompleteness and inconsistency, which can lead to erroneous outcomes when performing operations on improperly formatted data. Thus, data prep operations, such as Augment and Normalize can solve these cases and improve the overall results.
Moreover, Zero-Shot and Few-Shot QA methods also show considerable performance gains. Zero-Shot QA sees an improvement of 59.80% on WikiTQ and 2.50% on TabFact, while Few-Shot QA improves by 0.50% on WikiTQ and 0.89% on TabFact. These improvements are largely due to AutoPrep’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers directly from tables.
To further validate the impact of data prep, we analyze performance across tables of varying sizes. We categorize the tables in WikiTQ and TabFact into small (fewer than 2048 tokens), medium (2048 to 4096 tokens), and large (more than 4096 tokens) categories. While all tables in TabFact are small, WikiTQ includes 4040 small tables, 200 medium tables, and 104 large tables. As shown in Figure 7, AutoPrep achieves significant gains of 5.88% and 23.67% on medium and large tables, respectively. In addition, Few-Shot QA demonstrates limited improvement for small tables, because it has already achieved a high accuracy.
These experimental results clearly show that the integration of AutoPrep enhances the performance of all TQA methods, which demonstrates the critical role of question-aware data preparation in optimizing TQA performance.
6.4. Evaluation on Planner Agent
Table 4. Comparison of different Methods for Planner and Programmer.
Agent Method  WikiTQ  TabFact
Planner Direct Prompting $58.17$ $83.35$
Chain-of-Clauses $65.45$ $86.41$
Programmer Code-generation $62.82$ $81.97$
Tool-augmented $65.45$ $86.41$
Exp-3: How does our proposed Analysis Sketch method in suggesting high-level operations compared with Direct Prompting method? We change the high-level operation suggestion methods in Planner and evaluate the framework variants on WikiTQ and TabFact dataset. The results are reported in Table 4.
As shown in the table, Analysis Sketch based method outperforms Direct Prompting method by 12.52% and 3.67%, which indicates the superiority of our proposed method in generating more accurate high-level operations. Moreover, we find that the performance improvement on WikiTQ is much more significant than that in TabFact. This is because the WikiTQ dataset has relatively large tables and complex questions, which could make the high-level operation suggestion problem more challenging to be solved using direct prompting. In Summary, the experimental results prove that our proposed Analysis Sketch method is more effective in suggesting high-level operations, especially for difficult TQA instances.
6.5. Evaluation on Programmer Agents
Exp-4: How does our proposed tool-augmented method in generating low-level operations compared with existing code generation methods? We change the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default. We evaluate the performance on WikiTQ and TabFact and report the results in Table 4.
As illustrated in Table 4, our proposed tool-augmented method achieves better performance compared with the existing code generation method by an improvement of 4.19% and 5.42%. Considering the high-level operations input to the Programmer agents are the same, we can conclude that selecting function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations. Moreover, when using demonstrations constructed from same table-question pairs for the prompt of programmer agents, Tool-augmented method can save 18.07% input tokens for low-level operation generation. The above experimental results show the superiority of our method in both accuracy and cost.
Table 5. The average debugging round for Code-generation (Code) and Tool-augmented (Tool) Method.
Method  WikiTQ  TabFact
Code $1.35$ $1.28$
Tool $1.0$ $1.0$
Moreover, to compare the debugging method of these two methods, we record the average debugging rounds in Table 5. Our tool-augmented method conduct debugging process with fewer rounds, implying the effectiveness of our proposed debugging method for tool-augmented method.
6.6. Ablation Study
Table 6. Ablation Study of AutoPrep.
Method  WikiTQ  TabFact
AutoPrep $\mathbf{65.45}$ $\mathbf{86.41}$
- Filter $62.78$ $84.98$
  - Augmenter $61.79$ $85.67$
  - Normalizer $62.02$ $83.79$
Exp-5: What is the contribution of each agent in AutoPrep? We conduct ablation experiments to evaluate the contribution of each agent in AutoPrep. Specifically, we ablate each Programmer agent including Filter, Augmenter and Normalizer and compare the performance with AutoPrep.
The results are reported in Table 6. For WikiTQ datasets, without column augmentation, the performance drops the most (5.92%). For TabFact datasets, the Normalizer matters the most with a performance drop by 3.13%. This is because that WikiTQ has more instances requiring string extraction or calculation to generate new columns for answering the question, while column type normalization is a primary issue in TabFact datasets. Moreover, for both datasets, each agent plays an essential role in data preparation for TQA, which brings performance improvement by at least 4.25% and 0.86%.
The experimental results have demonstrated the rationality of the multi-agent design in AutoPrep. By solving three types of data preparation tasks in TQA methods in a specialized manner, the performance will be improved.
6.7. Case Study
(a) Case one
(b) Case two
Figure 8. Two case studies selected from the WikiTQ dataset to qualitatively analyze effectiveness of AutoPrep.
This section presents two illustrative examples selected from the WikiTQ dataset to qualitatively analyze effectiveness of AutoPrep.
As shown in Figure 8(a), to answer the question based on the original table, a Text-to-SQL method tries to filter all records with a score greater than 30. However, due to the limitation of SQL grammar, it can only use the operation Like to extract records where number 31, 32, …, or 39 is in values of column Score. This could miss out values with score greater than 40. Thus, it outputs error answer “3”. AutoPrep improves the performance of TQA methods by solving question-related data prep issues of the table in two stages. Specifically, based on the question in Figure 8(a), we first suggest two required high-level operations, i.e., a $\mathtt{Filter}$ to select the related column Score and a $\mathtt{Augment}$ to generate a new column from Score representing whether the team score is over 30. Then, we select appropriate functions from specific function pool and implement corresponding arguments to implement these high-level operations. Finally, we output a prepared table which only has columns related to the question. By passing the processed table and original question into the Text-to-SQL method, it generate a SQL that can easily extract the correct answer “4”.
Figure 8(b) illustrates a TQA instance with redundancy and inconsistency issues. To answer the question, it requires to first extract records with AirDate previous to “November 15, 2007” and then select records with RankOverall between 30 to 39. However, the values in column AirDate can not be directly assessed for chronological order through string comparison and RankOverall values are not in numerical type. Thus, although the generate SQL query is logically correct, it produces an wrong answer. AutoPrep suggests three high-level operations to solve this. Specifically, AutoPrep adopts an $\mathtt{Filter}$ to extract all question-related columns. And utilize two $\mathtt{Normalize}$ operations to format values of AirDate into sortable strings and cast RankOverall column to the integer type. We select three low-level operations to implement these high-level operations and get a new table with AirDate in %Y-%m-%d and RankOverall in integers. Then, we adopt the Text-to-SQL method on the prepared table, which generates a correct answer “3”.
7. Related Work
Tabular Question Answering. Most of the SOTA solutions for TQA rely on Large Language Models (LLMs) (Brown et al., 2020; Guo et al., 2024), as TQA requires NL understanding and reasoning over tables. The existing solutions can be broadly divided into two categories: end-to-end reasoning (Wei et al., 2022) and code generation (Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024). End-to-end reasoning approaches, which prompt an LLM to directly generate answers through next token prediction (Wei et al., 2022), may struggle when processing large tables due to the input token limitation and long-context understanding challenges (Li et al., 2024). On the other hand, code generation approaches leverage LLMs to generate programming code (e.g., Python) that extracts answers from tables, providing better explainability and achieving higher accuracy on real-world TQA tasks (Cheng et al., 2023; Zhang et al., 2024). For example, Dater (Ye et al., 2023) proposes to decompose questions and use code to deconstruct table contexts, enhancing end-to-end table reasoning with decomposed sub-tables and sub-questions. ReAcTable (Zhang et al., 2024) utilizes the ReAct paradigm (Yao et al., 2023) to generate codes which are executed step-by-step to extract intermediate question-related sub-tables for deducing the answer.
However, current LLM-based TQA methods mainly focus on the reasoning process of answer extraction instead of thorough and comprehensive data prep. While Binder (Cheng et al., 2023) proposes integrating SQL with an LLM-based API to incorporate external knowledge, which may partly address augmentation tasks, it does not account for a wide range of data prep tasks, such as normalization and column filtering. In this paper, we highlight the significance of question-aware data prep for TQA which could be a bottleneck restricting the performance. We also point out primary challenges in question-aware TQA data prep and propose an automatic multi-agent framework to solve them.
Traditional Data Preparation. Data Prep techniques have been widely adapted to a wide range of tasks, such as machine learning (ML) and exploratory data analysis (EDA), as they turn raw data into a format that is ready for downstream tasks. A complete data preparation process usually involves a series of steps, such as data transformation, cleaning, and feature engineering. Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify optimized data prep operation plan. Other approaches, such as Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020), apply meta-learning techniques to discover promising data prep configurations. Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques. Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) have explored the use of reinforcement learning to generate plans autonomously. For human-in-the-loop data prep, systems like DORIAN (Redyuk et al., 2022), Auto-Suggest (Yan and He, 2020), and Auto-Pipeline (Yang et al., 2021) incorporate recommendation-based methods to assist users. HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
In this paper, we propose AutoPrep to generate question-specific data prep operations for TQA tasks. The key difference between these methods and AutoPrep is that the data prep operations generated by AutoPrep are tailored to specific user questions and customized for the tables, as different questions may have different data prep requirements, even for the same table.
LLMs-based Multi-Agent Framework. A multi-agent LLM framework refers to a well-designed hierarchical structure consisting of multiple LLM-based agents and scheduling algorithms (Wang et al., 2024a). Compared with single-agent methods based on prompting techniques, such structures are better suited for handling complex tasks like software development, issue resolution, and code generation (Hong et al., 2023; Qian et al., 2023; Chen et al., 2024; Tao et al., 2024; Ishibashi and Nishimura, 2024; Islam et al., 2024). AutoTQA (Zhu et al., 2024) propose a multi-agent framework supporting Autonomous Tabular Question Answering, which decomposes complex TQA problems step-by-step and design agents to identify and correct shortcomings in the problem-solving process.
While AutoTQA improves table analysisx, this paper focuses on the performance bottleneck caused by data prep issues in TQA. Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
8. Conclusion
In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework designed to support automatic data prep for TQA tasks. AutoPrep consists of three stages: (1) the Planning stage, which suggests high-level data prep operations; (2) the Programming stage, which generates low-level implementations for each high-level operation; and (3) the Executing stage, which runs the Python code and iteratively debugs. We also propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation. Additionally, we introduce an effective method to ensure successful execution of the physical plan. Extensive experiments on WikiTQ and TabFact demonstrate that: (1) AutoPrep achieves state-of-the-art performance in data prep for TQA, and (2) integrating AutoPrep’s data prep improves the overall performance of TQA methods.

### v2

@@
 Abstract.
-Answering natural language (NL) questions about tables, which is referred to as Tabular Question Answering (TQA), is important because it enables users to extract meaningful insights quickly and efficiently from structured data, bridging the gap between human language and machine-readable formats.
-Many of these tables originate from web sources or real-world scenarios, necessitating careful data preparation (or data prep for short) to ensure accurate answers.
-However, unlike traditional data prep, question-aware data prep introduces new requirements, which include tasks such as column augmentation and filtering for given questions, and question-aware value normalization or conversion.
+Answering natural language (NL) questions about tables, known as Tabular Question Answering (TQA), is crucial because it allows users to quickly and efficiently extract meaningful insights from structured data, effectively bridging the gap between human language and machine-readable formats.
+Many of these tables are derived from web sources or real-world scenarios, which require meticulous data preparation (or data prep) to ensure accurate responses.
+However, preparing such tables for NL questions introduces new requirements that extend beyond traditional data preparation.
+This question-aware data preparation involves specific tasks such as column augmentation and filtering tailored to particular questions, as well as question-aware value normalization or conversion, highlighting the need for a more nuanced approach in this context.
 Because each of the above tasks is unique, a single model (or agent) may not perform effectively across all scenarios.
@@
 Programmer: Translates this logical plan into a physical plan by generating the corresponding low-level code.
-Executor: Iteratively executes and debugs the generated code to ensure correct outcomes.
+Executor: Executes the generated code to process the table.
 To support this multi-agent framework, we design a novel Chain-of-Clauses reasoning mechanism for high-level operation suggestion, and a tool-augmented method for low-level code generation.
-Extensive experiments on real-world TQA datasets demonstrate that AutoPrep can significantly improve the SOTA TQA solutions through question-aware data preparation.
+Extensive experiments on real-world TQA datasets demonstrate that AutoPrep can significantly improve the state-of-the-art TQA solutions through question-aware data preparation.
 PVLDB Reference Format:
@@
 PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at https://anonymous.4open.science/r/AutoPrep-Ano.
+The source code, data, and/or other artifacts have been made available at https://github.com/fmh1art/AutoPrep.
 1. Introduction
-Tabular Question Answering (TQA) refers to the tasks of answering natural language (NL) questions based on provided data tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023).
-TQA empowers non-technical users such as domain scientists to easily analyze tabular data and thus has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022).
+Tabular Question Answering (TQA) refers to the task of answering natural language (NL) questions based on provided tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023).
+TQA empowers non-technical users such as domain scientists to easily analyze tabular data and has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022).
+As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024b) rely on large language models (LLMs), either utilizing LLMs as black-boxes or leveraging them for code generation.
 (a) WikiTQ
@@
 (c) Query-Aware Column Filtering: Irrelevant columns increase complexity and cost of LLMs, which also lead to errors in TQA methods.
-Figure 2. Examples of Question-Aware Data Preparation for TQA tasks.
-Because TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024) naturally rely on large language models (LLMs).
-These solutions either utilize LLMs as black-boxes or leverage them for code generation.
-However, question-aware data preparation is not explicitly explored in these approaches, which limits the effectiveness of LLM-based methods.
-Figure 1 presents the results of an error analysis of LLM-based code generation (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019).
-The analysis reveals that 84% and 76% of the errors stem from insufficient data prep at various granularities (i.e., table-level, column-level, and cell-level) in relation to the NL questions.
-A similar trend is observed when using LLMs as black-boxes, though those results are not displayed here due to space constraints.
-These findings underscore the necessity for meticulous data prep to ensure accurate answers.
-To better understand these cases, let’s illustrate by examples.
+Figure 2. Examples of Question-Aware Data Preparation.
+Question-Aware Data Preparation.
+As many tables in TQA originate from web sources or real-world data, they demand meticulous data preparation (or data prep) to produce accurate answers.
+Figure 1 shows an error analysis of an LLM-based approach (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019).
+The results show that 84% and 76% of the errors arise from insufficiently addressing the data problems at various granularities (table-level, column-level, and cell-level) in relation to the NL questions.
+This highlights the critical role of thorough data prep in ensuring accurate responses for TQA.
+Figure 2 shows several typical data problems critical to the quality of TQA, which in turn motivate the need of question-aware data preparation.
 Example 0.
-Figure 2 shows several examples of question-aware data prep tasks that are needed for TQA.
-(a) Question-Aware Column Augmentation.
-Figure 2(a) illustrates the necessity of augmenting the tables with new columns, as the information that NL questions require is not explicitly available.
-For example, to answer $Q_{1}$, a new Country column should be created to augment $T_{1}$, e.g., by extracting country information from Cyclist.
-(b) Question-Aware Value Normalization: Figure 2(b) highlights the need for normalizing data types or formats in tables (e.g.,ITA” vs.
+(a) Column-Level: Missing Column. The example in Figure 2(a) shows the Country column that the NL question $Q_{1}$ requires is not explicitly available.
+In this case, an appropriate data prep could address this problem by extracting country information from Cyclist, i.e., augmenting the tables with a new Country column.
+Clearly, this data augmentation is driven by the specific need of the individual query.
+(b) Cell-Level: Inconsistent Data Types or Formats.
+Figure 2(b) highlights the inconsistency among the data in different columns (e.g.,ITA” vs.
 “Italia” in $T_{4}$, and “19-Oct” vs.
 “9/14” in $T_{6}$).
-Note that normalizing tables in TQA requires converting data types or formats based on the need of questions, meaning different questions may result in various operations.
+This essentially demands data normalization.
+This normalization has to be question-aware, because different questions may require different types of normalization.
 For example, as $Q_{5}$ asks for summing numbers in $T_{5}$ but the corresponding columns are of string type, we need to remove commas from numbers and convert strings to integers.
-(c) Question-Aware Column Filtering.
-Figure 2(c) highlights the importance of selecting question-relevant columns from tables, as irrelevant columns add complexity and can lead to errors in code generation.
-For example, the Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, which may mislead the LLM.
-Additionally, although $T_{8}$ contains 12 columns, only two are relevant to $Q_{8}$.
-Providing all columns to the LLM can thus result in long-context understanding challenges (Li et al., 2024).
-Example 1.1 highlights that different types of data prep operations are needed to satisfy the need of different NL questions.
-Given that data prep is a complex challenge without a one-size-fits-all solution, a sensible approach is thus to develop specialized methods w.r.t.
-each type of data prep operations.
-These specialized methods, such as data normalization, can involve either LLM-powered code generation or retrieving existing code from data prep libraries.
-An intelligent module can then automatically determine which operations to perform based on the properties of the question and the corresponding table.
-AutoPrep: A Multi-Agent LLM Framework.
-To address the problem, we propose AutoPrep, a multi-agent LLM framework that automatically prepares tables based on the specific needs of NL questions, eliminating the need for users to specify data prep operations.
-The core idea is to decompose the data prep task in TQA into sub-tasks, leveraging multiple agents specialized in different areas to ensure accurate and contextually relevant responses.
-Given an NL question over a table, AutoPrep performs data prep through the following three core components:
-  • Planner: This component is responsible for analyzing the natural language question and the structure of the given table.
-It generates a logical plan that outlines the sequence of high-level operations required to prepare the data for answering the question.
-These operations may include tasks such as column augmentation, value normalization, or filtering irrelevant data.
-  • Programmer: Once the logical plan is in place, the Programmer translates it into a physical plan by generating low-level executable code.
-This component is responsible for selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, customizing the code to match the table’s structure and data formats.
-  • Executor: The Executor is tasked with running the generated code and ensuring its correct execution.
-It performs iterative execution, where each operation is applied sequentially to the table, and any issues or errors are detected and debugged in real-time.
-Technical Challenges. Two key technical challenges arise in our proposed AutoPrep framework.
-(C1) Planner: High-Level Operation Suggestion.
-Even for the same table, different NL questions may require not only distinct data prep operations but also varying sequences of those operations.
-Thus, suggesting a tailored sequence of high-level operations that meet the specific needs of these questions is a significant challenge.
-(C2) Programmer: Low-Level Code Generation.
-A given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial.
-Generating optimized code for each operation that is customized for the tables presents a non-trivial task.
-For instance, consider the string-to-integer conversion in Example 2.4; failing to account for different value formats can lead to execution errors from a simple type conversion.
-Key Ideas.
-To address the challenge (C1), we propose a Chain-of-Clauses (CoC) based reasoning method for the Planner agent.
-This method analyzes the NL question and outputs a sequence of high-level desired data prep operations, namely an Analysis Sketch.
-Based on the Analysis Sketch, the Planner assigns high-level operations to specialized agents for low-level implementation (i.e., code generation).
-To tackle challenge (C2), we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions.
-We establish a search space of API functions for all high-level operations, design an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism.
+(c) Table-Level: Irrelevant Columns.
+As shown in Figure 2(c), some columns in the tables are irrelevant to the questions.
+For example, $Q_{7}$ does not need the record column in table $T_{7}$.
+On the other hand, although $T_{8}$ contains 12 columns, only two are relevant to $Q_{8}$.
+Filtering out the columns irrelevant to a question is important to TQA, as these irrelevant columns add complexity, potentially misleading LLMs to generate wrong answers.
+For example, the Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, which may confuse the LLM when answering $Q_{7}$.
+Moreover, this column filtering, avoiding feeding all columns to the LLM, effectively reduces LLM inference cost.
+The above examples demonstrate the urgent need of data prep to ensure the quality of TQA.
+However, blindly applying the existing data prep techniques to address all possible data problems tends to be neither effective nor efficient due to overlooking the specific needs of different NL questions.
+In this work, we thus propose to study a new problem, namely question-aware data preparation.
+Unlike the conventional data prep which prepares tables once at an offline stage to serve many incoming analytics queries (questions), question-aware data prep addresses question specific data problems at online query stage.
+Once receiving a question, it tailors the tables to the specific need of this question, e.g., performing tasks such as question-specific column augmentation and filtering, as well as question-aware value normalization or conversion.
+Key Challenges.
+Building a system to support question-aware data prep is challenging because of the diversity of both the data problems the tables encounter and the questions.
+As illustrated in the above examples, different questions may demand different types of data prep, e.g., augmentation, normalization, or filtering.
+Furthermore, even if facing the same type of data prep problem, different questions may require different ways to handle it.
+For example, the method that normalizes the date to a unified format is clearly different from that of normalizing strings to integers.
+Existing LLM-based TQA solutions do not explicitly consider the question-aware data prep problem, thus suffering from the data issues mentioned above.
+For a detailed analysis, please see Section 2.
+AutoPrep: A Hierarchical, Multi-Stage Approach.
+To address these challenges, we propose AutoPrep, which features two key ideas.
+First, inspired by the design principles of modern DBMS, in particular the distinction between the logical operations and their physical implementations, AutoPrep separates the high level, logical data prep operations and the concrete data prep methods.
+That is, it introduces a planning stage to generate a logical plan for each question.
+The logical plan is composed of a sequence of high-level data prep operations needed by the corresponding questions, including column augmentation, value normalization, or filtering irrelevant data, as shown in Figure 2.
+AutoPrep then introduces another stage to map the selected logical operations to the corresponding physical operations.
+In this way, AutoPrep decomposes a data prep task into smaller sub-tasks which are easier to solve than the original complex task.
+Moreover, it makes AutoPrep extensible, allowing us to easily develop new specialized implementations w.r.t.
+each type of data prep operations, or supporting new operation types.
+Second, unlike conventional DBMS, in AutoPrep the physical operators w.r.t.
+each logical operator are not developed beforehand.
+Instead, given a data prep operation in the logical plan, AutoPrep generates a physical implementation specialized to the question on the fly.
+Taking the question specific need into consideration, this customized implementation thus tends to meet the unique need of different questions.
+Driven by the above insight, we design the system into a multi-stage architecture.
+The Planning Stage: Unlike DBMS where the logical operators are already available beforehand via SQL queries, AutoPrep has to decide on the appropriate logical data prep operations and in turn form the logical plan by analyzing the NL question and the given table.
+This is done during the planning stage.
+The Programming Stage: This stage translates the logical plan into a physical plan by generating low-level executable code, which is responsible for selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, customizing the code to match the table’s structure and data formats, etc.
+The Executing Stage: This stage executes the generated code for each operation and returning any errors encountered to the Programming stage for debugging.
+We implements this AutoPrep architecture with the popular LLM-based Multi-Agent framework, which uses multiple small, independent agents to collaboratively solve complex problems.
+More specifically, we design a Planner agent.
+Corresponding to the Planning Stage, it suggests a tailored sequence of high-level operations that meet the specific needs of these questions, leveraging the semantics understanding and reasoning ability of LLMs.
+The key technical idea behind this Planner agent is a novel Chain-of-Clauses (CoC) reasoning method.
+This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer and guiding the agent’s reasoning based on this sketch.
+Compared with the popular Chain of Thoughts (CoT) methods (Wei et al., 2022), which decompose questions into sub-questions, our approach more effectively captures the relationships between questions and tables.
+AutoPrep also contains a set of Programmer agent, each of which synthesizes a question specific implementation w.r.t.
+one logical data prep operation.
+However, existing LLM-based code synthesis often produces overly generic code that fails to effectively address the heterogeneity challenges of tables.
+For example, values may have diverse syntactic formats (e.g., “19-Oct” and “9/14” in $T_{6}$) or or semantic representations (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code tailored to such variations.
+To address this, we develop a tool-augmented approach that improves the LLM’s code generation capabilities by utilizing predefined API functions.
+This approach effectively resolves the issue of overly generic code by allowing the LLM to leverage specialized functions tailored to variations in table values.
+Moreover, corresponding to the Executing stage, we design an Executor agent that executes the generated code to process the table.
 Contributions. Our contributions are summarized as follows.
-(1) We propose AutoPrep, an LLM-based multi-agent framework for automatic data prep in TQA.
-We formally define the problem and formalizes typical high-level data prep operations (Section 2), and introduce an overview of the AutoPrep framework (Section 3).
-(2) We develop effective techniques for our AutoPrep framework.
-We propose a novel Chain-of-Clauses reasoning mechanism (Section 4) to suggest high-level operations and a tool-augmented method to generate low-level code (Section 5).
+(1) We introduce a novel problem of question-aware data preparation for TQA, which is formally defined in Section 2.
+(2) We propose AutoPrep, an LLM-based multi-agent framework for question-aware data prep (Section 3).
+We develop effective techniques in AutoPrep for the Planner agent (Section 4) and the Programmer agents (Section 5).
 (3) We conduct a thorough evaluation on data prep in TQA (Section 6).
-Extensive experiments show that AutoPrep achieves new SOTA accuracy on well-adopted TQA datasets.
-We make the code used in our experiments available on an anonymous GitHub repo [^1] https://anonymous.4open.science/r/AutoPrep-Ano.
-2. Preliminaries
-This section first introduces Tabular Question Answering (TQA) in Section 2.1 and then formalizes the problem of question-aware data prep for TQA in Section 2.2.
-Then, we analyze the limitation of existing LLM-based methods in Section 2.3.
+Extensive experiments show that AutoPrep achieves new SOTA accuracy, outperforming existing TQA methods without data prep by 12.22 points on WikiTQ and 13.23 points on TabFact, and surpassing TQA methods with data prep by 3.05 points on WikiTQ and 1.96 points on TabFact.
+2. Question-Aware Data Prep for TQA
 2.1. Tabular Question Answering
@@
 Given the question $Q_{1}$, “Which country has the most medals?”, the answer $A_{1}$ should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”.
-Note that TQA often requires complex reasoning over tables.
+TQA often requires complex reasoning over tables.
 For instance, to answer question $Q_{2}$, we first need to calculate the “GDP growth rate” for all countries, then sort the countries by growth rate, and finally identify the country with the highest GDP growth rate, i.e., CHN.
 2.2. Data Prep for TQA
-Unlike traditional data prep, for question-aware data prep for TQA, we need to tailor table $T$ to the specific needs of question $Q$.
-Consider Example 2.1: careful data prep, such as extracting countries into a new column and then normalizing them as ISO codes, is crucial for generating an accurate answer.
-Moreover, it is important to note that different questions may require different data prep tasks, even for the same table.
-For instance, country extraction may not be necessary if the question is asking which cyclist, instead of country, won the most medals.
+Unlike traditional data prep, for question-aware data prep for TQA, we need to tailor table $T$ to specific needs of question $Q$.
 Data Prep Operations.
-To meet the new requirements of data prep for TQA, this paper defines high-level data prep operations (or operations for short) to formalize the question-aware data prep tasks, such as column augmentation, question-aware value normalization or conversion, and column filtering.
+To meet the new requirements of data prep for TQA, this paper defines high-level, logical data prep operations (or operations for short) to formalize the question-aware data prep tasks, such as column augmentation, question-aware value normalization or conversion, and column filtering.
 Formally, an operation, denoted as $o$, encapsulates a specific question-aware data prep task that transforms table $T$ into another table $T^{\prime}$, i.e., $T^{\prime}=o(T)$.
@@
 Given a high-level operation $o_{i}$, we define $f_{i}$ as its low-level implementation, either by calling a well-established algorithm from a known Python library or using a customized Python program to meet the requirements of $o_{i}$.
-For example, to implement the Augment operation $o_{1}$ that augments a Country column in $T_{1}$, a possible implementation $f_{i}$ is to write a customized program by using a specific regular expression as shown in Figure 2(a).
 Example 0.
@@
 Additionally, although $T_{8}$ has 12 columns, only two are relevant to $Q_{8}$, and providing all columns may cause long-context understanding challenges of LLMs (Li et al., 2024).
-Note that the above three types of operations address most of the errors caused by insufficient data prep, as shown in our error analysis in Figure 1.
-In fact, the set of data prep operations is extensible, in order to support new or domain-specific data prep tasks.
+The above three types of operations address most of the errors caused by insufficient data prep, as shown in our error analysis in Figure 1.
+In fact, due to the separation of logical operations and physical implementations in AutoPrep, the set of data prep operations can be easily extended to support new data prep tasks.
 Question-Aware Data Prep for TQA.
-Given an NL question $Q$ posed over table $T$, the problem of question-aware data prep for TQA is to generate a sequence of high-level operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$ as a logical plan.
+Given an NL question $Q$ posed over table $T$, question-aware data prep for TQA is to generate a sequence of high-level operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$ as a logical plan.
 Then, it generates a physical plan, where each operation $o_{i}$ is implemented by low-level code $f_{i}$, such that these operations transform $T$ into a new table $T^{\prime}$ that meets the needs of $Q$.
@@
 2.3. Limitations of Existing LLM-Based Solutions
-Naturally, LLMs can be utilized to support question-aware data prep for TQA.
-A straightforward way is to prompt an LLM with a task description and several demonstrations (i.e., examples), and requires the LLM to generate data prep code, as shown in Figure 3.
-However, as shown in Figure 2, the diversity of NL questions necessitates distinct data prep operations. Next, let’s see how a single LLM is used to solve this problem, by an example.
-Example 0.
-Consider the example table in Figure 3 with columns Date, Cyclist, Medal and Age, along with question: “Which country has the most medals in total in February”.
-When using few-shot prompting with a task description and a few demonstrations, this strategy may face limitations in both stages of data prep in TQA.
-First, as the number of demonstrations is limited, LLMs may struggle to decide which data prep operations are needed for the NL questions, leading to both false negatives (FNs) and false positives (FPs).
-For example, numeralizing Age in code block ② is an FP, as it is irrelevant to the question.
-Conversely, the LLM fails to normalize the Date column before extracting the month in code block ③, which is an FN.
-Second, due to input token limitations and challenges in long-context understanding (Li et al., 2024), a single LLM may struggle to generate fine-grained code for each operation.
-For instance, normalizing Medal in code block ③ does not handle all edge cases (e.g., ‘2*’), and the country extraction in code block ⑤ fails to correctly process ‘Dav.ITA’, which is formatted differently from other values.
-Example 2.4 shows that, because each of the above tasks is unique, a single model may not perform well across all scenarios.
-Thus, given that data prep is a complex challenge without a one-size-fits-all solution, a more effective approach is to develop specialized methods for each type.
-A centralized module can then determine which operations to perform and assign them accordingly.
-Figure 3. An LLM-based solution that prompts an LLM with a task description and several demonstrations (i.e., exemplars), and requires the LLM to generate data prep code.
+Figure 3. An LLM-based method with few-shot prompting for question-aware data prep.
+A straightforward solution to question-aware data prep is to prompt an LLM to prepare tables, leveraging its ability to interpret the specific requirements of NL questions.
+Specifically, given the input token constraints of LLMs, it is often infeasible to supply an entire table and request a fully prepared table as output.
+Therefore, a more effective strategy employs few-shot prompting to generate data prep programs.
+For instance, consider the table in Figure 3 with columns Date, Cyclist, Medal and Age, and question: “Which country has the most medals in total in February”.
+The few-shot prompting strategy prompts an LLM with a task description and a few demonstrations, and requests the LLM to generate Python programs as shown in Figure 3.
+However, this LLM-based solution may encounter the following limitations when performing question-aware data prep for TQA.
+First, given the difficulties of understanding both NL questions and tables, it is challenging to determine which data prep operations are specifically required to meet the needs of each NL question, thus often resulting in false negatives (FNs) and false positives (FPs).
+Take the results in Figure 3 as an example: converting Age to a numerical format in code block ② is an FP, as it is irrelevant to the question.
+Conversely, the method fails to normalize the Date column before extracting the month in code block ③, which constitutes an FN, because it does not recognize the inconsistent Date formats.
+Second, due to input token limitations and challenges in long-context understanding (Li et al., 2024), it is not easy to fully understand all possible issues in a table, and thus may struggle to generate customized programs to correct issues.
+For example, in Figure 3, the normalization of Medal in code block ① overlooks certain corner cases (e.g., “2*”), and the country extraction in code block ⑤ fails to handle “Dav.ITA”, which is formatted differently from other values.
+Recent methods, such as CoTable (Wang et al., 2024b) and ReAcTable (Zhang et al., 2024), can improve few-shot prompting by employing techniques like Chain-of-Thoughts (CoT) and ReAct.
+However, these methods remain insufficient to tackle the challenges, as they combine all diverse tasks, such as determining operations and implementing them at various granularities (table-level, column-level, and cell-level), within a single LLM agent.
+Existing studies (Ishibashi and Nishimura, 2024) have shown that a single LLM agent is often ineffective when tasked with handling a diverse range of operations, due to limited context length in LLMs and decreased inference performance with more input tokens.
+Figure 4. A Multi-agent Data Prep Framework for TQA (Left), which could be plugged to any other TQA methods (Right).
 3. An Overview of AutoPrep
-Figure 4.
-Our proposed Multi-agent Data Prep Framework for TQA (Left), which could be plugged to any other TQA method (Right), such as NL2SQL as shown in the Figure.
-The proposed framework AutoPrep decomposes the data prep process into two stages: ① The Planning stage and ② the Executing stage.
 To address the limitations, we propose AutoPrep, a multi-agent LLM framework that automatically prepares tables for given NL questions.
@@
 It directs the LLM to generate low-level implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$.
+Besides, it is also tasked for code debugging if any execution errors occur.
   3. (3)
-    An [Uncaptioned image] Executor Agent: the Executing stage. It executes the given low-level code and debugs iteratively to improve the generated code quality.
+    An [Uncaptioned image] Executor Agent: the Executing stage. It executes the given low-level code and reports error messages if any code bugs occur.
 After that, an [Uncaptioned image] Analyzer agent extracts the answer from the prepared table.
@@
 (a) The Planning stage: The [Uncaptioned image] Planner suggests the following high-level operations to address the specific NL question:
-  • $T^{\prime}=$Filter($T$, [Date, Cyclist, Medal]) that filters out column Age, which is irrelevant to the question;
   • $T^{\prime}$[Country]=Augment(“Extract country code”, Cyclist) that extracts the country information from column Cyclist, producing a new Country column, in response to the “which country” part of the question.
@@
   • Normalize(“Format date as %m-%d”, Date) standardizes the values in the Date column into a unified format to support the ‘in February’ condition in the question.
+  • $T^{\prime}=$Filter($T$, [Date, Country, Medal]) that filters out column Age, which is irrelevant to the question;
 (b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e.,, [Uncaptioned image] Augment, [Uncaptioned image] Normalizer and [Uncaptioned image] Filter, corresponding to the Augment, Normalize and Filter operations, respectively.
 Each specialized Programmer focuses on generating executable code for its assigned operations.
-(c) The Executing stage: a Executor agent iteratively refines the generated code if any errors occur.
+(c) The Executing stage: an Executor agent iteratively refines the generated code if any error occurs.
 After these stages, AutoPrep generates a prepared table $T^{*}$, which is then fed into an [Uncaptioned image] Analyzer agent to produce the answer ITA.
-The Planner agent and the three Programmer agents (i.e., Augmenter, Normalizer, and Filter) are the core modules of AutoPrep, which will be discussed in more details below.
 The Planner Agent.
-Given a question $Q$ and a table, the Planner agent generates a sequence of high-level operations for the specific question, as illustrated in the previous example.
 The key challenge here is how to suggest question-aware operations that address specific NL questions.
@@
 The Programmer Agents.
-Given a high-level operation $o_{i}$, a Programmer agent is responsible for generating low-level code $f_{i}$ to execute the operation.
-As discussed previous, we develop specialized agents i.e., Augmenter, Normalizer, and Filter for various types of operations.
 The key challenge is that a given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial.
@@
 The central Planner agent can then determine which operations should be performed and assign them accordingly.
-4. The Planner Agent
+4. The Planner
 Given a question $Q$ posed over a table $T$, Planner aims to generate a logical plan that outlines a sequence of high-level operations needed to prepare the table for answering the question.
-In this section, we first introduce a direct prompting method in Section 4.1 that better demonstrates the key challenge in designing the Planner agent.
-Then, to address the challenge, we propose our Chain-of-Clauses (CoC) reasoning method in Section 4.2.
-Figure 5. Our proposed approaches to logical plan generation in the Planner agent: (a) A straightforward direct prompting method, and (b) a more effective Chain-of-Clauses (CoC) reasoning method.
 4.1. A Direct Prompting Method
-The most common way to generate a logical plan is to directly prompt an LLM using a typical in-context learning approach, as shown in Algorithm 1.
+The most common way to generate a logical plan is to directly prompt an LLM using a typical in-context learning approach.
 The inputs are a question $Q$, a table $T$, a set $\Sigma$ of specifications for each operation type, and an LLM $\theta$.
@@
 The output of the algorithm is a set $O$ of high-level operations like those shown in Figure 4, e.g., Normalize(“Cast to INT”, Medal).
-Algorithm 1 Direct Prompting for the Planner Agent
-1:A question $Q$, a table $T$, an operation specification set $\Sigma$, and a large language model (LLM) $\theta$.
-2:// Prepare $k$ in-context exemplars (i.e., demos)
-3:$\mathcal{D}=\{(T_{i},Q_{i},\Sigma,O_{i})\}_{i=1}^{k}$
-4:// Prompt the LLM for inferring the operations
-5:$O\leftarrow\theta(T,Q,\Sigma,\mathcal{D})$
-6:return $O$
-Specifically, Algorithm 1 first prepares a set $\mathcal{D}$ of $k$ exemplars $\{(T_{i},Q_{i},\Sigma,O_{i})\}_{i=1}^{k}$ that demonstrates to LLM $\theta$ the goal of this logical plan generation task.
-Then, the algorithm prompts the LLM to infer the operations $O$ for our input table $T$ and question $Q$.
 Example 0.
@@
 This is because the phrase “the most” in the question suggests the need for type conversion of Medal, but there are no clues from the question to normalize Date, unless we observe from the table that the values are inconsistent.
+Figure 5. Our proposed approaches to logical plan generation in the Planner agent: (a) A straightforward direct prompting method, and (b) a more effective Chain-of-Clauses (CoC) reasoning method.
 Key Challenge.
@@
 4.2. A Chain-of-Clauses Method
-To address this challenge, we propose a Chain-of-Clauses (CoC) reasoning method the decomposes the entire process of logical plan generation into two phrases, as shown in Figure 5(b).
+To address this challenge, we propose a Chain-of-Clauses (CoC) reasoning method that decomposes the entire process of logical plan generation into two phases, as shown in Figure 5(b).
   • The first phase leverages an LLM to generate an SQL-like Analysis Sketch, outlining how the table should be transformed to produce the answer.
@@
 For each clause, we associate the corresponding data in $T$ (e.g., values in column Date) with it to prompt the LLM for generating possible high-level operations (e.g., Normalize).
-Compared with existing CoT methods (Wei et al., 2022), which simply break down questions into sub-questions, our approach is more effective for logical plan generation due to the following reasons.
-First, our method decomposes the question into a set of analysis steps over the table, formalized as an SQL-like Analysis Sketch, which simplifies the task of logical plan generation.
-More importantly, our method jointly considers each analysis step along with the corresponding relevant data to prompt the LLM, effectively capturing the relationships between the question and the table.
-Algorithm 2 CoC Reasoning for the Planner Agent
-1:A question $Q$, a table $T$, an operation specification set $\Sigma$, and a large language model (LLM) $\theta$.
-2:$O\leftarrow\emptyset$
-3:// Phase I: Analysis Sketch generation
-4:$S\leftarrow\theta(Q,T,\{(Q_{i},T_{i},s_{i})\})$
-5:// Phase II: High-level operation generation
-6:for each clause $s$ in $S$ do
-7:    $T^{s}\leftarrow$ extract_data($T$)
-8:    $o\leftarrow\theta(s,T^{s},\Sigma,\{s_{i},T^{s}_{i},\Sigma,o_{i}\})$
-9:    $O\leftarrow O\cup\{o\}$
-10:end for
-11:return $O$
-Specifically, Algorithm 2 presents the pseudo-code of our CoC reasoning method, which takes as input a question $Q$, a table $T$, a specification set $\Sigma$, and an LLM $\theta$ to produce an operation set $O$.
-To this end, we introduce a key concept of SQL-like Analysis Sketch, which is formally defined as follows.
+Compared with existing CoT methods (Wei et al., 2022), which simply break down questions into sub-questions, our approach is more effective for logical plan generation.
+First, our method decomposes the question into a set of analysis steps over table, formalized as an SQL-like Analysis Sketch, which simplifies the task of logical plan generation.
+More importantly, our method jointly considers each analysis step along with the corresponding relevant columns (instead of the whole table) to prompt the LLM, effectively capturing the relationships between the question and the data.
 SQL-like Analysis Sketch. An Analysis Sketch $S$ is an SQL-like statement, which can be represented as follows.
@@
 Note that $f(\{A_{i}\})$ is a user-defined function (UDF) that maps existing columns $\{A_{i}\}$ in the table into a new column.
-Figure 5(b) shows an example Analysis Sketch with a UDF function that specifies a new column Country, i.e., $f(T)$ AS Country.
+Figure 5(b) shows an example Analysis Sketch with a UDF function that specifies a new column Country, i.e., $f({\tt Cyclist})$ AS Country.
 Obviously, this UDF is introduced because the Analysis Sketch contains a GROUP BY Country clause.
-Given the key concept Analysis Sketch, Algorithm 2 produces a logical plan via two phases, as shown in the following example.
 Example 0.
@@
 (b) Phase II - Operation Generation: In this phase, the algorithm iteratively examines the clauses in Analysis Sketch $S$ as follows.
-  • $f(T)$ AS Country: this clause and the entire table $T$ are used to prompt the LLM to generate an Augment operation.
+  • $f({\tt Cyclist})$ AS Country: this clause and relevant columns are used to prompt the LLM to generate an Augment operation.
   • ${\tt SUM}({\tt Medal})$: this clause and the values in column Medal are used to prompt the LLM to generate a Normalize operation.
-  • Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize operation.
-Finally, the algorithm prompt the LLM to generate a Filter operation that only selects the columns relevant to the Analysis Sketch.
-5. The Programmer & Executor Agents
-Figure 6. Our proposed tool-augmented method for physical plan generation and execution.
-The Programmer agents translate a high-level logical plan into a physical plan by generating low-level code, which is then passed to an Executor agent for iterative code execution and debugging.
-A straightforward approach is to prompt an LLM with the logical plan using typical in-context learning with several exemplars, asking the LLM to generate code for each high-level operation in the plan.
-The code is then executed iteratively, and if any runtime errors occur, the Debugger is prompted with the error messages for debugging.
+  • Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize.
+Finally, the algorithm generates a Filter operation that only selects columns relevant to the Analysis Sketch.
+5. The Programmer & Executor
+Programmer agents translate a high-level logical plan into a physical plan by generating low-level code, which is then passed to an Executor agent for code execution and interactive debugging.
+A straightforward approach is to prompt an LLM with the logical plan using in-context learning with several exemplars, asking the LLM to generate code for each high-level operation in the plan.
+The code is then executed iteratively, and if any runtime errors occur, the Programmer is prompted with the error messages for debugging.
 However, the above method may have limitations, as the generated code may be overly generic and unable to effectively address the heterogeneity challenge of the tables.
 Specifically, many tables originate from web sources or real-world scenarios, where values have diverse syntactic (e.g., “19-Oct” and “9/14” in $T_{6}$) or semantic formats (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code customized to these tables.
-For example, consider table $T_{4}$ in 2(b).
+For example, consider table $T_{4}$ in Figure 2(b).
 Given a Normalize operation to standardize country formats, a generic function from an existing Python library may not be sufficient to transform country names (e.g., “Italia”) to their ISO codes (e.g., “ITA”).
@@
 To address these limitations, we introduce a tool-augmented method for the Programmer agents, enhancing the LLM’s code generation capabilities by utilizing pre-defined API functions, referred to as the function pool in this paper.
-Specifically, we first introduce the key idea of our tool-augmented method in  5.1, then present the function pool in Section 5.2, and finally develop an algorithm for code generation and execution in Section 5.3.
 5.1. Tool-Augmented Method: Key Idea
+Figure 6. Our proposed tool-augmented method for physical plan generation and execution.
 Figure 6 provides an overview of our tool-augmented method for physical plan generation and execution.
-Given a table $T$ and a set $O={o_{1},\ldots,o_{|O|}}$ of high-level operations, the method iteratively generates and executes low-level code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$.
+Given a table $T$ and a set $O=\{o_{1},\ldots,o_{|O|}\}$ of high-level operations, the method iteratively generates and executes low-level code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$.
 Specifically, in the $i$-th iteration, given a high-level operation $o_{i}$ (e.g., Augment) and an intermediate table $T_{i}$, the method generates low-level code and executes it to produce an updated table $T_{i+1}$ through the following two steps.
@@
 Code Execution and Debugging.
-Given the function generated in the previous step, the Executor agent then applies the function over table $T_{i}$ and support iterative debugging if there are any runtime errors or incorrect results.
+Given the function generated in the previous step, the Executor agent then applies the function over table $T_{i}$.
+If any bugs occur during execution, it captures and summarizes error messages and returns to corresponding Programmer agent.
 For instance, as shown in Figure 6, the first lambda function for the argument func only extracts countries based on parentheses, which may produce incorrect results for the value “Dav.ITA”, as it is formatted differently from other values.
-In this case, based on the execution results, the Executor agent will refine the regular expression in the lambda function until no runtime errors or incorrect results occur.
-After completing these two steps, the method produces a new intermediate table $T_{i+1}$ and proceeds to the next operation, $o_{i+1}$, i.e., the Normalize operation, as shown in Figure 6.
+In this case, based on the execution results, the Executor agent records the error log and examines relevant data to summarize reasons, which will be passed to the corresponding Programmer to modify the code.
+After the execution and debugging process, the method produces a new intermediate table $T_{i+1}$ and proceeds to the next operation, $o_{i+1}$, i.e., a Normalize operation in Figure 6.
 5.2. Search Space of Function Pools
-This section presents a search of function pools for the different types of high-level operations formalized in the current AutoPrep.
+This section presents a search space of function pools for different types of high-level operations formalized in the current AutoPrep.
 Function pool for $\mathtt{Augment}$.
@@
 It has two arguments: ${\tt column}$, representing the name of the source column, and ${\tt func}$, representing the substring extraction function.
-Using table $T_{1}$ from Figure 2(a) as an example, to extract the country code, we would set ${\tt column}$ to Cyclist and ${\tt func}$ to the lambda function lambda x: re.search(r’(.?)(.*?)( .
+Using table $T_{1}$ from Figure 2(a) as an example, to extract the country code, we would set ${\tt column}$ to Cyclist and ${\tt func}$ to a lambda function lambda x: re.search(r’(.?)(.*?)( .
 ∗ ?
@@
   • $\mathtt{calculate}$.
-This function generates a new column through arithmetic operations of columns such as addition, subtraction, multiplication, or a combination of these.
-To use this operator, we need to provide two arguments: ${\tt columns}$ and ${\tt func}$.
-The argument ${\tt columns}$ should be a list of source column names.
-Similar to $\mathtt{extract}$, the ${\tt func}$ argument is typically a lambda function.
-However, for $\mathtt{calculate}$, the input to the function should be a dictionary of values rather than a string, since we may perform calculations on multiple columns.
-For example, given a table with two columns A and B, representing the scores of team A and team B, we need to generate a new column ScoreDifference.
-In this case, we set ${\tt columns}$ to [’A’, ’B’] and ${\tt func}$ to the specific lambda function lambda x: x[’A’]-x[’B’].
+The function generates a new column through arithmetic operations of existing columns.
+It has two arguments: ${\tt columns}$, representing a list of source columns and ${\tt func}$, representing a lambda function with input to be a dictionary of values, since we may perform calculations on multiple columns.
+For example, to generate a new column GrowthRate for $T_{2}$ in Figure 2(a), we set ${\tt columns}$ to [’2012’, ’2013’] and ${\tt func}$ to a lambda function lambda x: (x[’2013’]-x[’2012’])/x[’2012’].
   • $\mathtt{map\mbox{-}to\mbox{-}boolean}$ and $\mathtt{concatenate}$.
 Both functions generate a new column from multiple columns.
-The difference is that one generates boolean values, while the other concatenates new strings.
+The difference is that one generates boolean values, while the other concatenates strings.
 Both operators take two arguments, ${\tt columns}$ and ${\tt func}$, similar to those in the $\mathtt{calculate}$ operator.
@@
 To use it, two arguments need to be specified: ${\tt column}$, representing the source column, and ${\tt func}$, representing a lambda function.
-For example, given the table $T_{5}$ in Figure 2(b), we use this function to standardize the Females column to integers, with ${\tt column}$ set to Females and ${\tt func}$ as lambda x: int(x.replace(’,’, ’’)).
+For example, given the table $T_{5}$ in Figure 2(b), we use this function to standardize the Females column to integers, with ${\tt column}$ as Females and ${\tt func}$ as lambda x: int(x.replace(’,’, ’’)).
   • $\mathtt{format\mbox{-}datetime}$.
@@
 To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$.
-It has two arguments: ${\tt column}$ and ${\tt trans_{d}ict}$.
-The $\tt trans_{d}ict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$.
-Given the original table $T_{4}$ in Figure  2(b), we set $trans_{d}ict$ to {’Italia’: ’ITA’}.
+It has two arguments: ${\tt column}$ and ${\tt trans\_dict}$.
+The $\tt trans\_dict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$.
+Given the original table $T_{4}$ in Figure  2(b), we set $\tt trans\_dict$ to {’Italia’: ’ITA’}.
   • $\mathtt{infer}$. Similar to the one for the $\mathtt{Augment}$ operation, this function uses LLMs to deduce values that could not be processed by the functions above.
@@
 Discussions.
-It is important to note that the search space is extensible, allowing additional functions or external APIs to be easily incorporated into the function pools.
-Moreover, AutoPrep can also include external API functions, such as specific data prep functions for various semantic types (e.g., date, name, company) developed in the dataprep library (Peng et al., 2021).
-5.3. Code Generation and Execution
-Algorithm 3 Low-level Code Generation and Execution
-1:A question $Q$, a table $T$, a number $r$, a high-level operation set $O$ and three function pools $\mathcal{F}_{\tt Aug}$, $\mathcal{F}_{\tt Norm}$ and $\mathcal{F}_{\tt Filter}$, an LLM $\theta$
-2:$F\leftarrow\varnothing,~{}T_{1}\leftarrow T$
-3:for each operation $o_{i}$ in $O$ do
-4:    // Step 1: Function selection and argument inference
-5:    $\mathcal{F}\leftarrow$ Assign($o_{i}$, [$\mathcal{F}_{\tt Aug}$, $\mathcal{F}_{\tt Norm}$, $\mathcal{F}_{\tt Filter}$])
-6:    $f_{i}\leftarrow\theta(o_{i},Q,T,\mathcal{F})$
-7:    // Step 2: Code generation and execution
-8:    while $f_{i}$ has errors &\&& ${\tt debug\_round}<r$ do
-9:         $f_{i}\leftarrow$ Debug($Q,T,f_{i}$)
-[…]
-11:end for
-12:return $T_{|O|+1}$
-Algorithm 3 presents the pseudo-code of our tool-augmented approach for code generation and execution.
-The algorithm takes as input a question $Q$, a table $T$, a set $O$ of high-level operations, three function pools, and an LLM, and produces a prepared table $T_{|O|+1}$.
-The algorithm iteratively examines the operations in $O$.
-In each iteration, it first generates a function from the corresponding function pool and refines it through multiple rounds of debugging.
-The algorithm then applies the generated function to the corresponding intermediate table and proceeds to the next iterations.
+Our search space is extensible, allowing additional functions or external APIs to be easily incorporated into the function pools, such as specific data prep functions for various semantic types developed in the dataprep library (Peng et al., 2021).
 6. Experiments
 6.1. Experimental Setup
+Datasets.
+We evaluate our proposed framework AutoPrep using two well-adopted benchmarking datasets WikiTQ (Pasupat and Liang, 2015) and TabFact (Chen et al., 2019) and a very recent benchmark TableBench (Wu et al., 2024).
+The statistics of the datasets are shown in Table 1.
+(1) WikiTQ contains complex questions annotated by crowd workers based on diverse Wikipedia tables.
+WikiTQ comprises 17,689 question-answer pairs in the training set and 4,344 in the test set with significant variation in table size.
+The answers in WikiTQ can be categorized into three forms: (i) string from a cell in the table, (ii) string from analyzing, and (iii) list of cells in the table.
+(2) TabFact provides a collection of Wikipedia tables, along with manually annotated NL statements.
+One requires to deduce the relations between statements and tables as “true” (a statement is entailed by a table) or “false” (a statement if refuted by a table).
+To reduce the experimental cost without losing generality, we choose the small test set provided by (Ye et al., 2023) which contains 2024 table-statement pairs.
+The tables in TabFact are much smaller than those in WikiTQ.
+The answers are binary with nearly equal proportions.
+(3) TabBench provides an evaluation of table question answering capabilities within four major categories, including multi-hop fact checking (FC), multi-hop Numerical Reasoning (NR), Trend Forecasting and Chart Generation.
+We evaluate AutoPrep on TabBench-FC and TabBench-NR, where one requires to conduct more complex and multi-hop reasoning over tables for answering a question.
+The answers in TabBench could be a string or list of items.
+Baselines. We consider the following two categories of baselines:
+(1) TQA Methods w/o Data Prep directly generate the answer or utilize programs to extract the answer from the table without considering data prep. We implement four primary TQA baselines.
+End2EndQA (End2End) (Chen, 2022) utilizes the in-context learning abilities of LLMs to generate the answer for TQA task based on the supervision of human-designed demonstrations.
+We implement End2End method with prompt and demonstrations provided by (Cheng et al., 2023).
+Chain-of-Thought (CoT) (Wei et al., 2022) prompts LLMs to generate the reasoning process step-by-step before generating the final answer. We implement CoT with the prompt provided by (Chen, 2022).
+NL2SQL (Rajkumar et al., 2022) first translates the question into a SQL program and then executes it to get the final answer from the table for the question.
+We use the prompt from (Cheng et al., 2023) to implement NL2SQL.
+NL2Py uses Python code to process and reason over the tables.
+To construct the prompt for NL2Py, we use TQA instances in NL2SQL prompt and manually write the Python code to process the table and generate the final answer.
 Table 1. Statistics of Datasets.
-Dataset # Num. Row. Col. Ans.
+Dataset # Rec. # Row. # Col. Ans. Types
 WikiTQ $22,033$ 4$\sim$753 3$\sim$25 string / list (3.05%)
 TabFact $2,024$ 5$\sim$47 5$\sim$14 true / false (49.60%)
-Datasets.
-We evaluate our proposed framework AutoPrep using two well-adopted benchmarking datasets WikiTQ (Pasupat and Liang, 2015) and TabFact (Chen et al., 2019).
-The statistics are shown in Table 1.
-  • WikiTQ.
-WikiTableQuestions (WikiTQ) is a dataset designed for QA tasks, containing complex questions annotated by crowd workers based on diverse Wikipedia tables.
-WikiTQ comprises 17,689 question-answer pairs in the training set and 4,344 in the test set with significant variation in table size.
-The answers in WikiTQ can be categorized into three forms: (i) string from a cell in the table, (ii) string from analyzing, and (iii) list of cells in the table.
-  • TabFact.
-Table-Fact-Checking (TabFact) dataset provides a collection of Wikipedia tables, along with manually annotated NL statements.
-One requires to deduce the relations between statements and tables as “true” (a statement is entailed by a table) or “false” (a statement if refuted by a table).
-To reduce the experimental cost without losing generality, we choose the small test set provided by (Khot et al., 2022) which contains 2024 table-statement pairs.
-The tables in TabFact are much smaller than those in WikiTQ.
-The answers are binary with nearly equal proportions.
+TableBench $886$ 2$\sim$212 2$\sim$20 string / list (31.49%)
+(2) TQA Methods with Data Prep. We also investigate four SOTA TQA methods that consider data prep tasks in their question answering process.
+ICL-Prep.
+uses few-shot in-context learning (ICL) demonstrations to guide LLMs in generating programs for data prep, as shown in Figure 3.
+Subsequently, we employ NL2SQL to extract answers from the cleaned tables.
+Dater (Ye et al., 2023) addresses the TQA task by decomposing the table and question.
+It first selects relevant columns and rows to obtain a sub-table and then decomposes the origin question into sub-questions.
+Dater answers these sub-questions based on the sub-tables to generate the final answer.
+We use code in (cod, 2023b) for implementation.
+Binder (Cheng et al., 2023) enhances the NL2SQL method by integrating LLMs into SQL programs.
+It uses LLMs to incorporate external knowledge bases and directly answer questions that are difficult to resolve using SQL alone.
+We utilize the original code provided by (cod, 2023a).
+ReAcTable (Zhang et al., 2024) uses the ReAct paradigm to extract relevant data from the table using Python or SQL code generated by LLMs.
+Once all relevant data is gathered, it asks the LLMs to predict the answer.
+We run the original code from (cod, 2024b) and keep all settings as default.
+Notice that the original code does not include prompts for TabFact, we generate it based the WikiTQ prompt.
+Chain-of-Table (CoTable) (Wang et al., 2024b) enhances the table reasoning capabilities of LLMs by predefining several common atomic operations (including data prep operations) that can be dynamically selected by the LLM.
+These operations form an “operation chain” that represents the reasoning process over a table and can be executed either via Python code or by prompting the LLM.
+We implement CoTable using the original code from (cod, 2024a).
+Table 2. Improvement of data prep for TQA (the best results are in bold and the second-best are underlined).
+Method DeepSeek GPT3.5 Llama3 QWen2.5
+WikiTQ TabFact WikiTQ TabFact WikiTQ TabFact WikiTQ TabFact
+End2End 56.6556.65{56.65}56.65 81.7781.77{81.77}81.77 $52.56$ 71.5471.54{71.54}71.54 58.7258.72{58.72}58.72 81.2781.27{81.27}81.27 60.0160.01{60.01}60.01 81.1781.17{81.17}81.17
+    + AutoPrep $63.14$ $\uparrow 6.49$ $82.11$ $\uparrow 0.34$ $61.21$ $\uparrow 8.65$ $71.79$ $\uparrow 0.25$ $61.23$ $\uparrow 2.51$ $84.19$ $\uparrow 2.92$ $63.42$ $\uparrow 3.41$ $83.05$ $\uparrow 1.88$
+CoT $54.95$ 82.0282.02{82.02}82.02 53.4853.48{53.48}53.48 $65.37$ $40.75$ 80.9380.93{80.93}80.93 59.6759.67{59.67}59.67 82.3182.31{82.31}82.31
++ AutoPrep $61.12$ $\uparrow 6.17$ $82.26$ $\uparrow 0.24$ $60.01$ $\uparrow 6.53$ $74.36$ $\uparrow 8.99$ $56.01$ $\uparrow 15.26$ $83.65$ $\uparrow 2.72$ $62.02$ $\uparrow 2.35$ $\underline{85.67}$ $\uparrow 3.36$
+NL2Py 59.3559.35{59.35}59.35 $68.13$ 53.5953.59{53.59}53.59 66.1566.15{66.15}66.15 $50.12$ $76.24$ $53.02$ $72.63$
++ AutoPrep $\underline{65.86}$ $\uparrow 6.51$ $\underline{87.35}$ $\uparrow 19.22$ $\underline{64.69}$ $\uparrow 11.1$ $\underline{84.83}$ $\uparrow 18.68$ $\underline{62.55}$ $\uparrow 12.43$ $\underline{85.42}$ $\uparrow 9.18$ $\underline{68.65}$ $\uparrow 15.63$ 85.72 $\uparrow 13.09$
+NL2SQL $52.83$ $70.21$ $52.90$ $64.71$ 51.8051.80{51.80}51.80 $75.15$ $56.86$ $80.09$
++ AutoPrep 66.09 $\uparrow 13.26$ 87.85 $\uparrow 17.64$ 64.75 $\uparrow 11.85$ 84.19 $\uparrow 19.48$ 63.72 $\uparrow 11.92$ 85.72 $\uparrow 10.57$ 68.72 $\uparrow 11.86$ $85.33$ $\uparrow 5.24$
 Evaluation Metrics.
-We adopt the evaluator in Binder (Cheng et al., 2023) to avoid the situation where program executions are likely to be semantically correct but fail to match the gold answer exactly.
+We adopt the evaluator in Binder (Cheng et al., 2023) to avoid the situation where program executions are likely to be semantically correct but fail to match the golden answer exactly.
 For example, when the SQL outputs 1/0 for yes/no questions, the Binder evaluator will conduct a pre-matching check and translate it to boolean values first.
-It will avoid mismatches of correct prediction and golden answer pairs like 7.45 and 7.4499, 4 years and 4, etc.
+It will avoid mismatches of correct prediction and golden answer, such as pairs like 7.45 and 7.4499, 4 years and 4.
 Afterwards, we adopt accuracy as our evaluation metric.
-Baselines.
-We evaluate AutoPrep using two categories of methods: (a) TQA w/o Data Prep: Zero-Shot QA and Few-Shot QA (Chen, 2022), as well as Text-to-SQL (Rajkumar et al., 2022); (b) TQA with Data Prep: Binder (Cheng et al., 2023) and ReAcTable (Zhang et al., 2024).
-We further describe them below.
-  • Zero-Shot and Few-Shot QA (Chen, 2022).
-Zero-Shot QA directly asks an LLM to generate the answer based on the serialized table and question.
-To improve the performance, Few-Shot QA further provides several demonstrations to enhance the In Context Learning ability of the LLM.
-We implement these two methods with the prompt provided by (Cheng et al., 2023).
-  • Text-to-SQL (Rajkumar et al., 2022).
-Text-to-SQL is a natural code generation method for the TQA tasks.
-Consider a serialized table and a question, it first generates the SQL program and then executes it to get the final answer from the table for the question.
-  • Binder (Cheng et al., 2023).
-Binder further enhances the Text-to-SQL method by integrating LLMs into SQL programs.
-It uses LLMs to incorporate external knowledge bases and directly answer questions that are difficult to resolve using SQL alone.
-We utilize the code provided by (cod, 2023).
-  • ReAcTable (Zhang et al., 2024).
-ReAcTable uses the ReAct paradigm to extract relevant data from the table using Python or SQL code generated by LLMs.
-Once all relevant data is gathered, it asks the LLMs to predict the answer.
-We run the original code from (cod, 2024) and kept all settings as default.
-Implementation Details.
-We mainly use DeepSeek (Guo et al., 2024) API [^2] https://platform.deepseek.com/usage model as the foundation model for each LLM-based agent.
-We also explore other proprietary LLMs, such as GPT-3.5 (Brown et al., 2020), using the GPT-3.5-turbo-0613 version, as adopted by (Wang et al., 2024b).
-To ensure reproducibility, we set the temperature parameter for all agents to 0.01.
+Backbone LLMs.
+We evaluate our method using representative LLMs as backbones.
+For closed-source LLMs, we select DeepSeek (Guo et al., 2024) (DeepSeek-V2.5-Chat) and GPT3.5 (Brown et al., 2020) (GPT3.5-Turbo-0613).
+For open-source LLMs, we choose Llama3 (Dubey et al., 2024) (Llama-3.1-70B-Instruct) and QWen2.5 (Yang et al., 2024) (QWen2.5-72B-Instruct) for evaluation.
+Experiment Settings.
+We provide detailed prompts of each component in AutoPrep in our technical report (tec, 2025) due to the space limit.
+For a fair comparison, we set the maximum token input of all methods as 8192.
+Moreover, we set the temperature parameter of all methods to 0.01 for reproducibility.
+6.2. Improvement of Data Prep for TQA
+Exp-1: Impact of question-aware data prep on TQA performance.
+We integrate AutoPrep into our four TQA baselines w/o data prep, in order to investigate the impact of question-aware data prep on TQA performance.
+The results are reported in Table 2.
+As demonstrated, integrating AutoPrep significantly improves the performance of all evaluated methods.
+Notably, NL2SQL shows the most substantial gains, achieving an average accuracy improvement of 12.22 on WikiTQ and 13.23 on TabFact across all LLM backbones.
+Similarly, NL2Py also shows notable improvements in its performance, after being integrated with AutoPrep.
+This significant improvement is attributed to the sensitivity of NL2SQL and NL2Py to data incompleteness and inconsistency, which can lead to erroneous outcomes when performing operations on improperly formatted data.
+Thus, data prep operations, such as Augment and Normalize can solve these cases and improve the overall results.
+Moreover, End2End and CoT methods also show considerable performance gains.
+For example, End2End shows an average accuracy gain of 5.27 and 1.35 on WikiTQ and TabFact respectively.
+These improvements are largely due to AutoPrep ’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers directly from tables.
+Finding 1: The integration of AutoPrep improves the accuracy of TQA methods w/o data prep by 12.22 points on WikiTQ and 13.23 points on TabFact, respectively.
+Since “NL2SQL + AutoPrep” achieves the best accuracy in most cases, we take its results as default for further comparison.
 6.3. Data Prep Method Comparison
-Table 2. Comparison of AutoPrep with data prep methods for TQA (the best results are bolded, and the second-best results are underlined).
-Method DeepSeek GPT3.5
-WikiTQ TabFact WikiTQ TabFact
-Binder $56.98$ $79.64$ $\underline{56.74}$ $\underline{79.17}$
-ReAcTable $\underline{64.25}$- $51.80$ $72.80$
-AutoPrep $\mathbf{65.45}$ $\mathbf{86.41}$ (87.01) $\mathbf{63.86}$ $\mathbf{84.29}$
-Exp-1: Comparison of AutoPrep with other data prep methods for TQA.
-We conduct experiments to evaluate AutoPrep against current state-of-the-art data prep methods for TQA on two datasets: WikiTQ and TabFact.
-Specifically, we compare AutoPrep with Binder and ReAcTable, presenting the results in Table 2.
-Notably, due to the lack of public scripts for TabFact from ReAcTable (Zhang et al., 2024), we report results only for DeepSeek on WikiTQ and use GPT-3.5 results from their original paper.
-The results clearly demonstrate that AutoPrep achieves state-of-the-art performance on both the WikiTQ and TabFact datasets when using DeepSeek and GPT-3.5 as the foundation model.
-In particular, AutoPrep outperforms Binder by 14.86% on WikiTQ and 8.50% on TabFact with DeepSeek.
-When using GPT-3.5, the performance improvements are 12.55% and 6.47%, respectively.
-Furthermore, when compared with ReAcTable, AutoPrep enhances performance by 1.87% on WikiTQ using DeepSeek.
-Remarkably, with GPT-3.5, AutoPrep surpasses ReAcTable by 23.28% and 15.78% on WikiTQ and TabFact, respectively.
-These improvements are attributed to AutoPrep’s comprehensive data prep operations, which are not fully addressed by Binder and ReAcTable.
-While Binder integrates LLMs to generate new columns to supplement external knowledge, it neglects critical aspects such as column filering and value normalization.
-In contrast, ReAcTable employs the ReAct data processing paradigm but lacks a specific decomposition of data prep tasks.
-Our experimental results highlight the importance of thorough data prep in enhancing TQA performance, which is often a significant bottleneck for existing methods.
-Additionally, as shown in Table 2, DeepSeek consistently outperforms GPT-3.5 while incurring lower API costs [^3] https://platform.deepseek.com/api-docs/zh-cn/quick_start/pricing/ https://openai.com/api/pricing/.
-Therefore, we have selected DeepSeek as our default foundation model for subsequent experiments.
-6.3. Improvement of Data Prep for TQA
-Table 3. Impact of data prep on TQA Methods.
-Method  WikiTQ  TabFact
-Zero-Shot QA $17.66$ $76.78$
-    + AutoPrep $28.22_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 10.56}}$ $78.70_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 1.92}}$
-Few-Shot QA $56.05$ $81.97$
-    + AutoPrep $56.33_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 0.28}}$ $82.70_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 0.73}}$
-Text-to-SQL $51.66$ $68.53$
-    + AutoPrep $\mathbf{65.45_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 13.79}}}$ $\mathbf{86.41_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 17.88}}}$
-(a) Zero-Shot QA.
-(b) Few-Shot QA.
-(c) Text-to-SQL.
-Figure 7. Performance on Small, Medium, Large tables from WikiTQ without/with data prep of AutoPrep.
-Exp-2: Impact of data prep on TQA performance.
-We evaluate the performance of three baseline TQA methods, namely Zero-Shot QA, Few-Shot QA, and Text-to-SQL with and without the data prep provided by AutoPrep.
-The results are detailed in Table 3.
-As demonstrated, integrating AutoPrep significantly improves the performance of all evaluated methods.
-Notably, Text-to-SQL experiences the most substantial enhancement, with performance increases of 26.69% on WikiTQ and 26.09% on TabFact.
-This significant improvement is attributed to Text-to-SQL’s sensitivity to data incompleteness and inconsistency, which can lead to erroneous outcomes when performing operations on improperly formatted data.
-Thus, data prep operations, such as Augment and Normalize can solve these cases and improve the overall results.
-Moreover, Zero-Shot and Few-Shot QA methods also show considerable performance gains.
-Zero-Shot QA sees an improvement of 59.80% on WikiTQ and 2.50% on TabFact, while Few-Shot QA improves by 0.50% on WikiTQ and 0.89% on TabFact.
-These improvements are largely due to AutoPrep’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers directly from tables.
-To further validate the impact of data prep, we analyze performance across tables of varying sizes.
-We categorize the tables in WikiTQ and TabFact into small (fewer than 2048 tokens), medium (2048 to 4096 tokens), and large (more than 4096 tokens) categories.
-While all tables in TabFact are small, WikiTQ includes 4040 small tables, 200 medium tables, and 104 large tables.
-As shown in Figure 7, AutoPrep achieves significant gains of 5.88% and 23.67% on medium and large tables, respectively.
-In addition, Few-Shot QA demonstrates limited improvement for small tables, because it has already achieved a high accuracy.
-These experimental results clearly show that the integration of AutoPrep enhances the performance of all TQA methods, which demonstrates the critical role of question-aware data preparation in optimizing TQA performance.
-6.4. Evaluation on Planner Agent
-Table 4. Comparison of different Methods for Planner and Programmer.
-Agent Method  WikiTQ  TabFact
-Planner Direct Prompting $58.17$ $83.35$
-Chain-of-Clauses $65.45$ $86.41$
-Programmer Code-generation $62.82$ $81.97$
-Tool-augmented $65.45$ $86.41$
-Exp-3: How does our proposed Analysis Sketch method in suggesting high-level operations compared with Direct Prompting method?
-We change the high-level operation suggestion methods in Planner and evaluate the framework variants on WikiTQ and TabFact dataset.
+Table 3. Experimental results of AutoPrep and TQA methods with data prep.
+Method DeepSeek GPT3.5 Llama3 QWen2.5
+ WikiTQ  TabFact  WikiTQ  TabFact  WikiTQ  TabFact  WikiTQ  TabFact
+ICL-Prep 56.5456.54{56.54}56.54 $80.53$ $55.71$ $73.91$ $50.05$ $75.20$ $57.00$ $80.14$
+Dater $48.32$ $83.05$ $52.81$ $72.08$ $43.53$ $74.01$ $58.78$ $79.84$
+Binder $56.81$ $82.81$ 56.7456.74{56.74}56.74 79.1779.17{79.17}79.17 $50.51$ $78.16$ $55.43$ 81.7281.72{81.72}81.72
+ReAcTable 64.1364.13{64.13}64.13 85.7185.71{85.71}85.71 $51.80$ $72.80$ 58.0158.01{58.01}58.01 80.0080.00{80.00}80.00 60.1560.15{60.15}60.15 $81.67$
+CoTable $\underline{64.53}$ $\underline{86.22}$ $\underline{59.94}$ $\underline{80.20}$ $\underline{62.22}$ $\underline{85.62}$ $\underline{64.41}$ $\underline{83.20}$
+AutoPrep 66.09 87.85 64.75 84.19 63.72 85.72 68.72 85.33
+Exp-2: Comparison of AutoPrep with previous SOTA TQA methods with data prep.
+We compare AutoPrep with TQA methods that consider data prep tasks in their question answering process, using a single LLM agent.
 The results are reported in Table 3.
-As shown in the table, Analysis Sketch based method outperforms Direct Prompting method by 12.52% and 3.67%, which indicates the superiority of our proposed method in generating more accurate high-level operations.
-Moreover, we find that the performance improvement on WikiTQ is much more significant than that in TabFact.
+The results clearly demonstrate that AutoPrep achieves the new SOTA performance on both the WikiTQ and TabFact datasets across all LLM backbones.
+In particular, AutoPrep outperforms ICL-Prep with a significant accuracy improvement by 11.00 and 8.33 on WikiTQ and TabFact on average, respectively.
+Other baselines, except Dater, generally outperform ICL-Prep by introducing specific technical optimizations to data prep.
+For instance, ReAcTable applies the ReAct paradigm to iteratively generate data prep operations rather than producing all code at once, and it explicitly considers data prep tasks (e.g., filtering and augmentation).
+Despite this, AutoPrep still outperforms the second best method by considerable improvement.
+Specifically, when compared with CoTable, AutoPrep achieves better performance, with an average accuracy improvement by 3.05 and 1.96 on WikiTQ and TabFact, respectively.
+These improvements are primarily attributed to our multi-agent framework, which effectively tackles the question-aware data prep challenge.
+The experimental results indicate that data prep is inherently complex and is hard to be resolved by a single, one-size-fits-all solution; instead, a more effective approach is to develop specialized methods for each type.
+A centralized module can then determine which operations to perform and assign them accordingly.
+Moreover, AutoPrep covers a broader range of data prep operations than these state-of-the-art TQA methods, thereby addressing gaps (e.g., normalization tasks) not fully explored by previous solutions.
+Finding 2: AutoPrep outperforms the the previous SOTA TQA methods with data prep by 3.05 points on WikiTQ and 1.96 points on TabFact, respectively,
+In addition, we find that DeepSeek achieves the better overall performance at lower API costs. Thus, we select DeepSeek as our default backbone LLM for subsequent experiments.
+Exp-3: Evaluation of AutoPrep on tables with various sizes.
+To further investigate AutoPrep, we analyze its performance across tables of varying sizes.
+While all tables in TabFact are small, we categorize the tables in WikiTQ into Small (fewer than 2048 tokens), Medium (2048 to 4096 tokens), and Large (more than 4096 tokens), resulting in 4040 Small tables, 200 Medium tables, and 104 Large tables.
+The results are reported in Figure 7.
+We observe that all baselines exhibit unstable performance, particularly on larger tables.
+For instance, although CoTable achieves the highest average accuracy among previous SOTA methods, it suffers an accuracy drop of 11.78 when reasoning over large tables.
+Likewise, ReAcTable shows relatively stable performance on medium tables (dropping by 2.15), yet its performance on large tables remains unsatisfactory (dropping by 17.53).
+(a) Small
+(b) Medium
+(c) Large
+Figure 7. Comparison of AutoPrep and other data prep methods on tables with different sizes on the WikiTQ dataset.
+In contrast, AutoPrep achieves the highest and most stable performance across tables of varying sizes.
+When processing medium and large tables, the accuracy of AutoPrep drops by 2.29 and 3.79.
+The key to its stability is that it employs specialized agents for each data prep task, mitigating the issue of exceeding prompt length limits.
+Furthermore, each agent generates program-based operators to handle data prep on the entire table, thereby avoiding information loss that occurs when large tables are cropped.
+Finding 3: By employing multiple agents and program-based operations, AutoPrep maintains stable performance as table size grows, ensuring that each data prep task is handled effectively without overwhelming a single model.
+6.4. Evaluation on Generalization
+Exp-4: Evaluating generalization on unseen datasets.
+As discussed previously, we evaluate the generalization capabilities of AutoPrep on two TabBench datasets.
+Specifically, we directly use the designed prompting strategies on the WikiTQ dataset, and examine whether these strategies can be generalized to TabBench.
+As shown in Figure 8, AutoPrep achieves the highest overall accuracy among all methods.
+Specifically, compared with the second best method ReAcTable, AutoPrep improves by 5.28, indicating the strong generalization capability of our method.
+The main reason is that AutoPrep utilizes tool-augmented method for physical plan generation and execution, generalizing well on unseen datasets.
+Finding 4: Utilizing tool-augmented execution enhances the generalization capabilities of AutoPrep on unseen datasets.
+(a) Numerical Reasoning
+(b) Fact Checking
+Figure 8. Evaluating Generalization on the TabBench dataset.
+6.5. Ablation Studies
+Exp-5: Evaluation on the Planner agent.
+We compare two Planner variants with different high-level operation suggestion methods, namely Direct Prompting and our proposed Chain-of-Clauses method, and report the results in Table 4(a).
+We observe that Chain-of-Clauses outperforms Direct Prompting by 7.92, 4.50, 5.27 in accuracy on WikiTQ, TabFact and TabBench respectively.
+This indicates the superiority of our proposed method in generating more accurate high-level operations.
+Moreover, we find that the performance improvement on WikiTQ is more significant than that on other datasets.
 This is because the WikiTQ dataset has relatively large tables and complex questions, which could make the high-level operation suggestion problem more challenging to be solved using Direct Prompting.
-In Summary, the experimental results prove that our proposed Analysis Sketch method is more effective in suggesting high-level operations, especially for difficult TQA instances.
-6.5. Evaluation on Programmer Agents
-Exp-4: How does our proposed tool-augmented method in generating low-level operations compared with existing code generation methods?
-We change the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default.
-We evaluate the performance on WikiTQ and TabFact and report the results in Table 4.
-As illustrated in Table 4, our proposed tool-augmented method achieves better performance compared with the existing code generation method by an improvement of 4.19% and 5.42%.
-Considering the high-level operations input to the Programmer agents are the same, we can conclude that selecting function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations.
-Moreover, when using demonstrations constructed from same table-question pairs for the prompt of programmer agents, Tool-augmented method can save 18.07% input tokens for low-level operation generation.
-The above experimental results show the superiority of our method in both accuracy and cost.
-Table 5. The average debugging round for Code-generation (Code) and Tool-augmented (Tool) Method.
-Method  WikiTQ  TabFact
-Code $1.35$ $1.28$
-Tool $1.0$ $1.0$
-Moreover, to compare the debugging method of these two methods, we record the average debugging rounds in Table 5.
-Our tool-augmented method conduct debugging process with fewer rounds, implying the effectiveness of our proposed debugging method for tool-augmented method.
-6.6. Ablation Study
-Table 6. Ablation Study of AutoPrep.
-Method  WikiTQ  TabFact
-AutoPrep $\mathbf{65.45}$ $\mathbf{86.41}$
-- Filter $62.78$ $84.98$
-  - Augmenter $61.79$ $85.67$
-  - Normalizer $62.02$ $83.79$
-Exp-5: What is the contribution of each agent in AutoPrep?
-We conduct ablation experiments to evaluate the contribution of each agent in AutoPrep.
-Specifically, we ablate each Programmer agent including Filter, Augmenter and Normalizer and compare the performance with AutoPrep.
-The results are reported in Table 6.
-For WikiTQ datasets, without column augmentation, the performance drops the most (5.92%).
-For TabFact datasets, the Normalizer matters the most with a performance drop by 3.13%.
-This is because that WikiTQ has more instances requiring string extraction or calculation to generate new columns for answering the question, while column type normalization is a primary issue in TabFact datasets.
-Moreover, for both datasets, each agent plays an essential role in data preparation for TQA, which brings performance improvement by at least 4.25% and 0.86%.
-The experimental results have demonstrated the rationality of the multi-agent design in AutoPrep.
-By solving three types of data preparation tasks in TQA methods in a specialized manner, the performance will be improved.
-6.7. Case Study
+Finding 5: Our proposed Chain-of-Clauses method is more effective in suggesting high-level operations.
 (a) Case one
 (b) Case two
-Figure 8. Two case studies selected from the WikiTQ dataset to qualitatively analyze effectiveness of AutoPrep.
-This section presents two illustrative examples selected from the WikiTQ dataset to qualitatively analyze effectiveness of AutoPrep.
-As shown in Figure 8(a), to answer the question based on the original table, a Text-to-SQL method tries to filter all records with a score greater than 30.
-However, due to the limitation of SQL grammar, it can only use the operation Like to extract records where number 31, 32, …, or 39 is in values of column Score.
-This could miss out values with score greater than 40.
-Thus, it outputs error answer “3”.
-AutoPrep improves the performance of TQA methods by solving question-related data prep issues of the table in two stages.
-Specifically, based on the question in Figure 8(a), we first suggest two required high-level operations, i.e., a $\mathtt{Filter}$ to select the related column Score and a $\mathtt{Augment}$ to generate a new column from Score representing whether the team score is over 30.
-Then, we select appropriate functions from specific function pool and implement corresponding arguments to implement these high-level operations.
-Finally, we output a prepared table which only has columns related to the question.
-By passing the processed table and original question into the Text-to-SQL method, it generate a SQL that can easily extract the correct answer “4”.
-Figure 8(b) illustrates a TQA instance with redundancy and inconsistency issues.
-To answer the question, it requires to first extract records with AirDate previous to “November 15, 2007” and then select records with RankOverall between 30 to 39.
-However, the values in column AirDate can not be directly assessed for chronological order through string comparison and RankOverall values are not in numerical type.
-Thus, although the generate SQL query is logically correct, it produces an wrong answer.
-AutoPrep suggests three high-level operations to solve this.
-Specifically, AutoPrep adopts an $\mathtt{Filter}$ to extract all question-related columns.
-And utilize two $\mathtt{Normalize}$ operations to format values of AirDate into sortable strings and cast RankOverall column to the integer type.
-We select three low-level operations to implement these high-level operations and get a new table with AirDate in %Y-%m-%d and RankOverall in integers.
-Then, we adopt the Text-to-SQL method on the prepared table, which generates a correct answer “3”.
+Figure 9. Two case studies selected from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
+Exp-6: Evaluation on the Programmer Agents. We change the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default.
+As illustrated in Table 4(b), our proposed tool-augmented method achieves better performance compared with the code generation method.
+Considering the high-level operations input to the Programmer agents are the same, we can conclude that selecting a function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations.
+Moreover, when using demonstrations constructed from same table-question pairs for the prompt of programmer agents, Tool-augmented method can save 18.07% input tokens for low-level operation generation.
+We also record the error ratio of these two methods, as shown in Table 4(b).
+The probability of bugs in our tool-augmented method is greatly reduced (e.g., from 4.51% to 1.50%), demonstrating its effectiveness.
+Finding 6: Our tool-augmented method performs better in both accuracy and cost, with fewer program errors.
+Table 4. Experimental Results of Ablation Studies.
+(a) Evaluation on the Planner agent.
+Method  WikiTQ  TabFact  TabBench
+Direct Prompting $58.17$ $83.35$ $44.83$
+Chain-of-Clauses 66.09 87.85 50.10
+(b) Evaluation on the Programmer Agents.
+Method Metric  WikiTQ  TabFact  TabBench
+Code Generation Acc $\uparrow$ $62.82$ $81.97$ $47.26$
+Err $\downarrow$ $4.51\%$ $4.50\%$ $4.73\%$
+Tool Augmented Acc $\uparrow$ 66.09 87.85 50.10
+Err $\downarrow$ 1.50% 0.05% 1.01%
+(c) Contribution of Each Programmer Agent in AutoPrep.
+Method  WikiTQ  TabFact  TabBench
+AutoPrep 66.09 87.85 50.10
+- Filter $62.78$ ($-3.31$) $84.98$ ($-2.87$) $47.67$ ($-2.43$)
+- Augmenter $61.79$ ${(-4.30)}$ $85.67$ ($-2.18$) $45.44$ ($-4.66$)
+- Normalizer $62.02$ ($-4.07$) $83.79$ ${(-4.06)}$ $41.58$ ${(-8.52)}$
+Exp-7: Contribution of each agent in AutoPrep. We ablate each Programmer agent including Filter, Augmenter and Normalizer and compare the performance with AutoPrep.
+As shown in Table 4(c), for WikiTQ, without column augmentation, the accuracy drops the most (4.30).
+For TabFact and TabBench, the Normalizer matters the most with an accuracy drop by 4.06 and 8.52.
+This is because that WikiTQ has more instances requiring string extraction or calculation to generate new columns for answering the question, while normalization is a primary issue in TabFact and TabBench.
+Moreover, for all datasets, each agent plays an essential role in data preparation for TQA, which brings accuracy improvement by at least 2.43, 2.18 and 4.06.
+Finding 7: Solving three types of data preparation tasks in TQA methods in a specialized manner improves the downstream TQA performance.
+6.6. Case Studies
+This section presents two illustrative examples from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
+To answer the question in Figure 9(a), CoTable generates an operator chain with an “add_col” operator, which is used to generate a new column Tol based on the division results of two existing columns Expense and Percent.
+This operator directly prompts an LLM to output a list containing all values of the new column, which is a challenging task for previous LLMs.
+Thus, CoTable generates a new column with two wrong values which leads to a wrong answer.
+For ReAcTable, although it generates a logically correct Python program, it ignores the types of existing columns which do not support for numerical calculation.
+Thus, ReAcTable also fails to augment the Tol column.
+AutoPrep addresses this issue by first utilizing a Planner agent to generate logical operators specifying data prep requirements, i.e., two $\mathtt{Normalize}$ to normalize column Expense and Percent, a $\mathtt{Augment}$ to generate the overall expenses Tol and a $\mathtt{Filter}$ to select related columns Year and Tol.
+These logical operators are passed to Programmer agents to generate physical operators consisting of a pre-defined Python function.
+For example, we call $\mathtt{to\mbox{-}numerical}$ function two times to cast Expense and Percent to numerical values.
+All physical operators are executed to process the original table and output a prepared table.
+Finally, an Analyzer agent based on NL2SQL method generates a SQL query to extract the correct answer “4214.17”.
+Figure 9(b) illustrates a TQA instance involving large tables and inconsistent values.
+AutoPrep solves this by splitting the TQA question into two phases, i.e., data prep and data analysis.
+It first generates logical operators including two $\mathtt{Normalize}$ and one $\mathtt{Filter}$.
+Next, the Programmer agents implement them with corresponding physical operators, which are executed to produce a prepared table.
+Based on this, the Analyzer agent extracts the final answer “8”.
 7. Related Work
 Tabular Question Answering.
-Most of the SOTA solutions for TQA rely on Large Language Models (LLMs) (Brown et al., 2020; Guo et al., 2024), as TQA requires NL understanding and reasoning over tables.
-The existing solutions can be broadly divided into two categories: end-to-end reasoning (Wei et al., 2022) and code generation (Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024).
-End-to-end reasoning approaches, which prompt an LLM to directly generate answers through next token prediction (Wei et al., 2022), may struggle when processing large tables due to the input token limitation and long-context understanding challenges (Li et al., 2024).
-On the other hand, code generation approaches leverage LLMs to generate programming code (e.g., Python) that extracts answers from tables, providing better explainability and achieving higher accuracy on real-world TQA tasks (Cheng et al., 2023; Zhang et al., 2024).
-For example, Dater (Ye et al., 2023) proposes to decompose questions and use code to deconstruct table contexts, enhancing end-to-end table reasoning with decomposed sub-tables and sub-questions.
-ReAcTable (Zhang et al., 2024) utilizes the ReAct paradigm (Yao et al., 2023) to generate codes which are executed step-by-step to extract intermediate question-related sub-tables for deducing the answer.
-However, current LLM-based TQA methods mainly focus on the reasoning process of answer extraction instead of thorough and comprehensive data prep.
-While Binder (Cheng et al., 2023) proposes integrating SQL with an LLM-based API to incorporate external knowledge, which may partly address augmentation tasks, it does not account for a wide range of data prep tasks, such as normalization and column filtering.
-In this paper, we highlight the significance of question-aware data prep for TQA which could be a bottleneck restricting the performance.
-We also point out primary challenges in question-aware TQA data prep and propose an automatic multi-agent framework to solve them.
+Most of the SOTA solutions for TQA rely on LLMs (Brown et al., 2020; Guo et al., 2024), as TQA requires NL understanding and reasoning over tables.
+Specifically, direct Prompting approaches (Chen, 2022; Wei et al., 2022), which prompt an LLM to directly generate answers through next token prediction, may struggle when processing large tables due to the input token limitation and long-context understanding challenges (Li et al., 2024).
+Code Generation approaches (Rajkumar et al., 2022) leverage LLMs to generate programming code (e.g., Python) that extracts answers from tables, providing better explainability and adaptability.
+Previous TQA methods, like Dater (Ye et al., 2023), Binder (Cheng et al., 2023), CoTable (Wang et al., 2024b) and ReAcTable (Zhang et al., 2024), also consider data prep in their question answering process.
+Specifically, Dater (Ye et al., 2023) prompts LLMs to select relevant columns related to answering the question, targeted at solving filtering tasks.
+Binder (Cheng et al., 2023) proposes to integrate SQL with an LLM-based API to incorporate external knowledge, which may partly address augmentation tasks.
+Similarly, CoTable (Wang et al., 2024b) addresses the filtering tasks and augmentation tasks by designing operators which are implemented by LLM completion.
+ReAcTable (Zhang et al., 2024) uses few-shot demonstrations to instruct LLMs to generate python code or SQL to address the augmentation and filtering tasks.
+However, previous TQA methods have the following limitations when handling data prep.
+First, all these methods do not address the cell-level normalization tasks, which account for the most significant error types, as indicated in Figure 1.
+Second, all these methods utilize a single LLM agent for both data prep and answer reasoning.
+Given that data prep is a complex challenge, these one-size-fits-all solutions still fail to achieve satisfactory performance as demonstrated in our experiments.
 Traditional Data Preparation.
-Data Prep techniques have been widely adapted to a wide range of tasks, such as machine learning (ML) and exploratory data analysis (EDA), as they turn raw data into a format that is ready for downstream tasks.
-A complete data preparation process usually involves a series of steps, such as data transformation, cleaning, and feature engineering.
-Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify optimized data prep operation plan.
-Other approaches, such as Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020), apply meta-learning techniques to discover promising data prep configurations.
+Data prep techniques have been widely adopted in various tasks.
+Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify data prep operations.
+Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020) apply meta-learning to discover promising operations.
 Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques.
-Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) have explored the use of reinforcement learning to generate plans autonomously.
-For human-in-the-loop data prep, systems like DORIAN (Redyuk et al., 2022), Auto-Suggest (Yan and He, 2020), and Auto-Pipeline (Yang et al., 2021) incorporate recommendation-based methods to assist users.
+Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) explore reinforcement learning techniques.
 HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
@@
 Compared with single-agent methods based on prompting techniques, such structures are better suited for handling complex tasks like software development, issue resolution, and code generation (Hong et al., 2023; Qian et al., 2023; Chen et al., 2024; Tao et al., 2024; Ishibashi and Nishimura, 2024; Islam et al., 2024).
-AutoTQA (Zhu et al., 2024) propose a multi-agent framework supporting Autonomous Tabular Question Answering, which decomposes complex TQA problems step-by-step and design agents to identify and correct shortcomings in the problem-solving process.
-While AutoTQA improves table analysisx, this paper focuses on the performance bottleneck caused by data prep issues in TQA.
+AutoTQA (Zhu et al., 2024) propose a multi-agent framework supporting Tabular Question Answering.
+While AutoTQA improves table analysis, this paper focuses on the performance bottleneck caused by data prep issues in TQA.
 Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
@@
 In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework designed to support automatic data prep for TQA tasks.
-AutoPrep consists of three stages: (1) the Planning stage, which suggests high-level data prep operations; (2) the Programming stage, which generates low-level implementations for each high-level operation; and (3) the Executing stage, which runs the Python code and iteratively debugs.
+AutoPrep consists of three stages: (1) the Planning stage, which suggests high-level data prep operations; (2) the Programming stage, which generates low-level implementations for each high-level operation; and (3) the Executing stage, which runs the Python code and report error messages.
 We also propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation.
 Additionally, we introduce an effective method to ensure successful execution of the physical plan.
-Extensive experiments on WikiTQ and TabFact demonstrate that: (1) AutoPrep achieves state-of-the-art performance in data prep for TQA, and (2) integrating AutoPrep’s data prep improves the overall performance of TQA methods.
+Extensive experiments on WikiTQ, TabFact and TabBench demonstrate that: (1) AutoPrep achieves SOTA performance in data prep for TQA, and (2) integrating AutoPrep’s data prep improves the performance of TQA methods.

### v3

@@
 However, preparing such tables for NL questions introduces new requirements that extend beyond traditional data preparation.
-This question-aware data preparation involves specific tasks such as column augmentation and filtering tailored to particular questions, as well as question-aware value normalization or conversion, highlighting the need for a more nuanced approach in this context.
+This question-aware data preparation involves specific tasks such as column derivation and filtering tailored to particular questions, as well as question-aware value normalization or conversion, highlighting the need for a more nuanced approach in this context.
 Because each of the above tasks is unique, a single model (or agent) may not perform effectively across all scenarios.
@@
 TQA empowers non-technical users such as domain scientists to easily analyze tabular data and has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022).
-As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024b) rely on large language models (LLMs), either utilizing LLMs as black-boxes or leveraging them for code generation.
-(a) WikiTQ
-(b) TabFact
-Figure 1. An error analysis of LLM-based TQA (using GPT-4) on two well-adopted datasets, with statistics calculated from a sample of 500 instances from each dataset.
-(a) Question-Aware Column Augmentation: Necessary information for QA is not explicitly available in tables.
-(b) Question-Aware Value Normalization: Heterogeneous data representations or formats in tables, and how to normalize is query-dependent.
-(c) Query-Aware Column Filtering: Irrelevant columns increase complexity and cost of LLMs, which also lead to errors in TQA methods.
-Figure 2. Examples of Question-Aware Data Preparation.
+As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024) rely on large language models (LLMs).
+Figure 1. An error analysis of LLM-based TQA (using GPT-4) on two well-adopted datasets.
+As many tables in TQA originate from web sources or real-world data, they demand meticulous data preparation (or data prep) to produce accurate answers.
+Figure 1 shows an error analysis of an LLM-based approach (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019) (More details of the error analysis can be found in our technical report (tec, 2025)).
+The results indicate that 84% and 76% of the errors stem from inadequately addressing data prep issues, including missing semantics, inconsistent values, and irrelevant columns, as illustrated as follows.
+(a) Missing Semantics: Required semantic information for NL questions is not explicitly present in the tables.
+(b) Value Inconsistency: Heterogeneous data representations or formats across table values, where normalization is question-dependent.
+(c) Irrelevant Columns: Presence of unnecessary columns that increase complexity and induce errors for LLMs.
+Figure 2. Examples of Data Preparation Issues for Table Question Answering.
+(1) Missing Semantics.
+This data prep issue arises when a table lacks the necessary semantics to address the specific requirements of the NL question.
+That is, although some columns in the table may be related to the question, they do not directly provide the required semantic information.
+As shown in Figure 2(a), the semantics needed for the NL questions, such as country and GDP, are not explicitly present in the tables.
+Therefore, to ensure accurate responses in TQA, it is essential to perform column derivation from existing columns, such as extracting country information from text or applying mathematical operations across multiple columns.
+(2) Inconsistent Values.
+This data preparation issue arises when the values within a particular column are inconsistent across different records due to varying data representations (e.g., inconsistent date formats), as illustrated in Figure 2(b).
+This situation necessitates column normalization, which must be question-aware, as different questions may require different types of normalization.
+For instance, as $Q_{5}$ requires comparing numbers in $T_{5}$ but the corresponding columns are of string type, it is necessary to convert the strings in fraction format into numerical values.
+(3) Irrelevant Columns.
+In some tables, although many columns are available, only a few are actually relevant to the NL question.
+For example, as shown in Figure 2(c), $Q_{7}$ does not require the Summary column in table $T_{7}$.
+Similarly, in $T_{8}$, although there are 12 columns, only 2 are needed to answer $Q_{8}$.
+Thus, column filtering, i.e., retaining only the columns directly related to the question, is crucial for TQA, as irrelevant columns add complexity and can mislead LLMs into producing incorrect answers.
+For instance, the Summary column in $T_{7}$ contains an average of 100 words, increasing the difficulty for LLMs to identify the relevant AirDate column.
+This highlights the critical role of thorough data prep in ensuring accurate TQA.
+Note that, although issues like missing values and duplicates are common in real-world data, they are less prevalent than the three highlighted issues in current TQA benchmarks.
+Nevertheless, we recognize their significance and leave the exploration of broader data prep challenges to future work.
 Question-Aware Data Preparation.
-As many tables in TQA originate from web sources or real-world data, they demand meticulous data preparation (or data prep) to produce accurate answers.
-Figure 1 shows an error analysis of an LLM-based approach (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019).
-The results show that 84% and 76% of the errors arise from insufficiently addressing the data problems at various granularities (table-level, column-level, and cell-level) in relation to the NL questions.
-This highlights the critical role of thorough data prep in ensuring accurate responses for TQA.
-Figure 2 shows several typical data problems critical to the quality of TQA, which in turn motivate the need of question-aware data preparation.
-Example 0.
-(a) Column-Level: Missing Column. The example in Figure 2(a) shows the Country column that the NL question $Q_{1}$ requires is not explicitly available.
-In this case, an appropriate data prep could address this problem by extracting country information from Cyclist, i.e., augmenting the tables with a new Country column.
-Clearly, this data augmentation is driven by the specific need of the individual query.
-(b) Cell-Level: Inconsistent Data Types or Formats.
-Figure 2(b) highlights the inconsistency among the data in different columns (e.g.,ITA” vs.
-“Italia” in $T_{4}$, and “19-Oct” vs.
-“9/14” in $T_{6}$).
-This essentially demands data normalization.
-This normalization has to be question-aware, because different questions may require different types of normalization.
-For example, as $Q_{5}$ asks for summing numbers in $T_{5}$ but the corresponding columns are of string type, we need to remove commas from numbers and convert strings to integers.
-(c) Table-Level: Irrelevant Columns.
-As shown in Figure 2(c), some columns in the tables are irrelevant to the questions.
-For example, $Q_{7}$ does not need the record column in table $T_{7}$.
-On the other hand, although $T_{8}$ contains 12 columns, only two are relevant to $Q_{8}$.
-Filtering out the columns irrelevant to a question is important to TQA, as these irrelevant columns add complexity, potentially misleading LLMs to generate wrong answers.
-For example, the Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, which may confuse the LLM when answering $Q_{7}$.
-Moreover, this column filtering, avoiding feeding all columns to the LLM, effectively reduces LLM inference cost.
-The above examples demonstrate the urgent need of data prep to ensure the quality of TQA.
-However, blindly applying the existing data prep techniques to address all possible data problems tends to be neither effective nor efficient due to overlooking the specific needs of different NL questions.
-In this work, we thus propose to study a new problem, namely question-aware data preparation.
-Unlike the conventional data prep which prepares tables once at an offline stage to serve many incoming analytics queries (questions), question-aware data prep addresses question specific data problems at online query stage.
-Once receiving a question, it tailors the tables to the specific need of this question, e.g., performing tasks such as question-specific column augmentation and filtering, as well as question-aware value normalization or conversion.
-Key Challenges.
-Building a system to support question-aware data prep is challenging because of the diversity of both the data problems the tables encounter and the questions.
-As illustrated in the above examples, different questions may demand different types of data prep, e.g., augmentation, normalization, or filtering.
-Furthermore, even if facing the same type of data prep problem, different questions may require different ways to handle it.
+In this paper, we thus propose to study a new problem, namely question-aware data preparation.
+The novelty of this problem lies in the need to deeply understand the semantics of the NL question in order to guide data prep operations over the structured table.
+Unlike traditional data prep, which is performed offline and independent of downstream tasks, question-aware data prep is conducted online and must dynamically tailor the table to the specific demands of the question.
+This introduces new challenges due to the diversity and ambiguity of NL questions and the heterogeneity of tabular data, requiring precise semantic alignment between the question and the structured table.
+For example, as shown in Figure 2(a), given question $Q_{1}$ that requests country-specific information from a text column, table $T_{1}$ needs to be transformed to extract and create a new Country column, which may not be considered in traditional data prep.
+However, building a system to support question-aware data prep is challenging due to the semantic alignment between the table and the NL question.
+Specifically, different questions may demand different types of data prep, e.g., derivation, normalization, or filtering.
+Furthermore, even if facing the same type of data prep issue, different questions may require different ways to handle it.
 For example, the method that normalizes the date to a unified format is clearly different from that of normalizing strings to integers.
-Existing LLM-based TQA solutions do not explicitly consider the question-aware data prep problem, thus suffering from the data issues mentioned above.
-For a detailed analysis, please see Section 2.
+Although LLMs show promise for interpreting semantics in NL questions, simply applying them to address all potential data prep issues has shown ineffective, often leading to false negatives and false positives, as illustrated in Section 2.3.
 AutoPrep: A Hierarchical, Multi-Stage Approach.
-To address these challenges, we propose AutoPrep, which features two key ideas.
-First, inspired by the design principles of modern DBMS, in particular the distinction between the logical operations and their physical implementations, AutoPrep separates the high level, logical data prep operations and the concrete data prep methods.
-That is, it introduces a planning stage to generate a logical plan for each question.
-The logical plan is composed of a sequence of high-level data prep operations needed by the corresponding questions, including column augmentation, value normalization, or filtering irrelevant data, as shown in Figure 2.
-AutoPrep then introduces another stage to map the selected logical operations to the corresponding physical operations.
-In this way, AutoPrep decomposes a data prep task into smaller sub-tasks which are easier to solve than the original complex task.
-Moreover, it makes AutoPrep extensible, allowing us to easily develop new specialized implementations w.r.t.
-each type of data prep operations, or supporting new operation types.
-Second, unlike conventional DBMS, in AutoPrep the physical operators w.r.t.
-each logical operator are not developed beforehand.
-Instead, given a data prep operation in the logical plan, AutoPrep generates a physical implementation specialized to the question on the fly.
-Taking the question specific need into consideration, this customized implementation thus tends to meet the unique need of different questions.
-Driven by the above insight, we design the system into a multi-stage architecture.
-The Planning Stage: Unlike DBMS where the logical operators are already available beforehand via SQL queries, AutoPrep has to decide on the appropriate logical data prep operations and in turn form the logical plan by analyzing the NL question and the given table.
-This is done during the planning stage.
-The Programming Stage: This stage translates the logical plan into a physical plan by generating low-level executable code, which is responsible for selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, customizing the code to match the table’s structure and data formats, etc.
-The Executing Stage: This stage executes the generated code for each operation and returning any errors encountered to the Programming stage for debugging.
-We implements this AutoPrep architecture with the popular LLM-based Multi-Agent framework, which uses multiple small, independent agents to collaboratively solve complex problems.
-More specifically, we design a Planner agent.
-Corresponding to the Planning Stage, it suggests a tailored sequence of high-level operations that meet the specific needs of these questions, leveraging the semantics understanding and reasoning ability of LLMs.
-The key technical idea behind this Planner agent is a novel Chain-of-Clauses (CoC) reasoning method.
-This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer and guiding the agent’s reasoning based on this sketch.
-Compared with the popular Chain of Thoughts (CoT) methods (Wei et al., 2022), which decompose questions into sub-questions, our approach more effectively captures the relationships between questions and tables.
-AutoPrep also contains a set of Programmer agent, each of which synthesizes a question specific implementation w.r.t.
-one logical data prep operation.
-However, existing LLM-based code synthesis often produces overly generic code that fails to effectively address the heterogeneity challenges of tables.
-For example, values may have diverse syntactic formats (e.g., “19-Oct” and “9/14” in $T_{6}$) or or semantic representations (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code tailored to such variations.
-To address this, we develop a tool-augmented approach that improves the LLM’s code generation capabilities by utilizing predefined API functions.
-This approach effectively resolves the issue of overly generic code by allowing the LLM to leverage specialized functions tailored to variations in table values.
-Moreover, corresponding to the Executing stage, we design an Executor agent that executes the generated code to process the table.
+To address the challenges, we propose AutoPrep, which features two key ideas.
+First, drawing inspiration from modern DBMS, particularly the distinction between logical operations and their physical implementations, AutoPrep separates high-level, logical data prep operations from the concrete methods used for execution.
+Specifically, it introduces a planning stage that generates a logical plan for each question, consisting of a sequence of high-level data prep operations tailored to the question’s needs, such as column derivation, normalization and filtering shown in Figure 2.
+In the next stage, AutoPrep maps these logical operations to the corresponding physical implementations.
+This separation allows AutoPrep to break down complex data prep tasks into smaller, more manageable sub-tasks, making the process more easier to solve.
+Moreover, this modular design makes AutoPrep extensible, enabling the development of specialized implementations for each type of data prep operation or the introduction of new operation types.
+Second, unlike conventional DBMSs, AutoPrep does not predefine physical operations for each logical operation.
+Instead, for each data prep operation in the logical plan, AutoPrep generates a physical implementation that is specialized to the specific question on the fly.
+By considering the unique requirements of the NL question, this customized implementation ensures that the data preparation is closely aligned with the needs of different questions.
+Building on the above insights, we design the AutoPrep system with a multi-stage architecture.
+  • The Planning Stage: Unlike traditional DBMS, where logical operations are already available beforehand via SQL queries, AutoPrep has to determine the appropriate logical data prep operations by analyzing the semantic alignment between the table and the NL question.
+This process occurs during the planning stage, where the logical plan is formed.
+  • The Programming Stage: This stage converts the logical plan into a physical plan by generating low-level executable code, selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, and customizing the code to align the table’s structure with the NL question’s semantics.
+  • The Executing Stage: This stage executes the generated code for each operation and returns any errors encountered to the Programming stage for debugging.
+We implement AutoPrep using the popular LLM-based Multi-Agent framework  (Han et al., 2024), which leverages multiple small, independent agents working collaboratively to solve complex problems.
+More specifically, we design a Planner agent, which corresponds to the Planning Stage and suggests a tailored sequence of high-level operations to meet the specific needs of the question, leveraging the semantics understanding and reasoning capabilities of LLMs.
+The core technical idea behind this Planner agent is a novel Chain-of-Clauses (CoC) reasoning method.
+This method translates the NL question into an Analysis Sketch, which outlines how the table should be prepared to produce the answer, guiding the agent’s reasoning based on this sketch.
+Compared to the popular Chain of Thoughts (CoT) methods (Wei et al., 2022), which decompose questions into sub-questions, our approach more effectively captures the semantic relationships between questions and tables.
+AutoPrep also includes a set of Programmer agents, each of which synthesizes a question-specific implementation for a given logical data prep operation.
+However, existing LLM-based code synthesis often generates overly generic code that struggles to effectively address the heterogeneity challenges of tables.
+For instance, values may have diverse syntactic formats (e.g.,September 1” and “11-24” in $T_{6}$) or semantic representations (e.g., “ITA” and “Italia” in $T_{1}$), making it difficult to generate code tailored to these variations.
+To address this, we propose a tool-augmented approach that enhances the LLM’s code generation capabilities by incorporating predefined API functions, which allows the LLM to generate more specialized code that accounts for variations in table values.
+Furthermore, corresponding to the Executing stage, we design an Executor agent that executes the code to process the table.
 Contributions. Our contributions are summarized as follows.
@@
 2.1. Tabular Question Answering
-Let $Q$ be a natural language (NL) question, and $T$ a table consisting of $m$ columns (i.e., attributes) ${A_{1},A_{2},\ldots,A_{m}}$ and $n$ rows ${r_{1},r_{2},\ldots,r_{n}}$, where $v_{ij}$ denotes the value in the $i$-th row and $j$-th column of the table.
-The problem of tabular question answering (TQA) is to generate an answer, denoted as $A$, in response to question $Q$ based on the information in table $T$.
+Let $Q$ be a natural language (NL) question, and $T$ a table consisting of $m$ columns (i.e., attributes) $\{A_{1},A_{2},\ldots,A_{m}\}$ and $n$ rows $\{r_{1},r_{2},\ldots,r_{n}\}$, where $v_{ij}$ denotes the value in the $i$-th row and $j$-th column of the table.
+The problem of tabular question answering (TQA) is to generate an answer $Ans$, in response to question $Q$ based on the information in table $T$.
 By the purposes of the questions, there are two main types of TQA problems: (1) table-based fact verification (Chen et al., 2019; Aly et al., 2021), which determines whether $Q$ can be entailed or refuted by $T$, and (2) table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022), which extracts or reasons the answer to $Q$ from $T$.
@@
 For instance, consider table $T_{1}$, which contains medal information for cyclists from different countries, with two columns: Cyclist and Medal.
-Given the question $Q_{1}$, “Which country has the most medals?”, the answer $A_{1}$ should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”.
+Given the question $Q_{1}$, “Which country has the most medals?”, the answer should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”.
 TQA often requires complex reasoning over tables.
@@
 2.2. Data Prep for TQA
-Unlike traditional data prep, for question-aware data prep for TQA, we need to tailor table $T$ to specific needs of question $Q$.
+In contrast to traditional data prep, question-aware data prep for TQA focuses on adapting the table $T$ to the specific informational needs of a given question $Q$, thereby enhancing the semantic alignment between the structured table and the NL question.
 Data Prep Operations.
-To meet the new requirements of data prep for TQA, this paper defines high-level, logical data prep operations (or operations for short) to formalize the question-aware data prep tasks, such as column augmentation, question-aware value normalization or conversion, and column filtering.
+To meet the new requirements of data prep for TQA, this paper defines high-level, logical data prep operations (or operations for short) to formalize the question-aware data prep tasks.
 Formally, an operation, denoted as $o$, encapsulates a specific question-aware data prep task that transforms table $T$ into another table $T^{\prime}$, i.e., $T^{\prime}=o(T)$.
-Specifically, this paper considers the following three types of operations.
-  • Augment: a data prep task that augments a new column for table $T$ based on existing columns, to better answer $Q$.
-This task typically involves column combinations via arithmetic operations, value extraction and inference, etc.
-  • Normalize: a data prep task that normalize types or formats of the values in a column of $T$ based on the needs of $Q$.
+As shown in Figure 1, the majority of TQA errors arise from inadequately addressing three key data prep issues: missing semantics, inconsistent values, and irrelevant columns.
+To address the challenges, we introduce three types of data prep operations.
+  • Derive: A data prep task that derives a new column for table $T$ from existing columns, aimed at addressing the challenge of missing semantics.
+This task typically involves operations such as combining columns through arithmetic computations, extracting relevant values, etc.
+  • Normalize: a data prep task that normalizes types or formats of the values in a column of $T$ based on the needs of $Q$, aimed at addressing the challenge of inconsistent values.
 This task typically involves value representation or format normalization, type conversion, etc.
-  • Filter: A data prep task that filters out columns in $T$ that are not relevant to answering question $Q$.
+  • Filter: A data prep task that filters out columns in $T$ that are not relevant to answer question $Q$, aimed at addressing the challenge of irrelevant columns.
 This is crucial for handling large tables to address the input token limitations and challenges in long-context understanding of LLMs (Li et al., 2024).
@@
 Figure 2 shows examples of question-aware data prep operations for TQA, along with their implementations in Python.
-(a) The Augment operation: Figure 2(a) shows three examples of Augment operations, i.e., extracting Country information from column Cyclist in $T_{1}$, computing new column GrowthRate using two columns in $T_{2}$, and inferring a status of IfLost by analyzing the scores in column Result.
+(a) The Derive operation: Figure 2(a) shows three examples of the Derive operations, i.e., extracting Country information from column Cyclist in $T_{1}$, computing new column GrowthRate using two columns in $T_{2}$, and inferring a status of IfLost by analyzing the scores in column Result.
 Note that the corresponding Python implementations are provided alongside the operations.
-(b) The Normalization operation: Figure 2(b) shows three examples of Normalize operations, i.e., normalizing country formats in $T_{4}$, converting strings into integers in $T_{5}$, and normalizing value formats in Date in $T_{6}$.
-We can see that, to implement a specific operation (e.g., in $T_{4}$), we sometimes need to call external tools, like a function clean_country from an external library.
-(c) The Filter operation: Figure 2(c) illustrates two examples of Filter operations in tables $T_{5}$ and $T_{6}$.
-The Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, potentially misleading the LLM.
+(b) The Normalize operation: Figure 2(b) shows three examples of Normalize operations, i.e., normalizing the formats of Surface, Time and Date for tables $T_{4}$, $T_{5}$ and $T_{6}$ respectively.
+(c) The Filter operation: Figure 2(c) illustrates two examples of Filter operations in tables $T_{7}$ and $T_{8}$.
+The Summary column in $T_{7}$ contains an average of 100 words, increasing the difficulty for LLMs to identify the relevant AirDate column.
 Additionally, although $T_{8}$ has 12 columns, only two are relevant to $Q_{8}$, and providing all columns may cause long-context understanding challenges of LLMs (Li et al., 2024).
-The above three types of operations address most of the errors caused by insufficient data prep, as shown in our error analysis in Figure 1.
-In fact, due to the separation of logical operations and physical implementations in AutoPrep, the set of data prep operations can be easily extended to support new data prep tasks.
+Remarks.
+While issues such as missing values and duplicates are common in real-world datasets, they are significantly less prevalent than the three highlighted challenges in existing TQA benchmarks.
+Thanks to AutoPrep’s separation of logical operations and physical implementations, the set of data prep operations can be easily extended to accommodate new types of tasks.
+We leave the exploration of broader data prep challenges to future work.
 Question-Aware Data Prep for TQA.
@@
 Then, it generates a physical plan, where each operation $o_{i}$ is implemented by low-level code $f_{i}$, such that these operations transform $T$ into a new table $T^{\prime}$ that meets the needs of $Q$.
-Example 0.
-To better answer the example question $Q_{1}$ over $T_{1}$ in Figure 2(a), we first generate an Augment operation to transform $T_{1}$ into $T^{\prime}_{1}$ with three columns: Cyclist, Medal, and Country.
-Next, we generate a Normalize operation, similar to the one in $T_{4}$, to standardize the values in Country into a unified format, producing table $T^{\prime\prime}_{1}$.
-Finally, we leverage LLMs, either as black-box or for code generation, to compute the answer $A_{1}$ to $Q_{1}$ from the prepared table $T^{\prime\prime}_{1}$.
-2.3. Limitations of Existing LLM-Based Solutions
+2.3. A Straightforward LLM-based Solution
 Figure 3. An LLM-based method with few-shot prompting for question-aware data prep.
 A straightforward solution to question-aware data prep is to prompt an LLM to prepare tables, leveraging its ability to interpret the specific requirements of NL questions.
-Specifically, given the input token constraints of LLMs, it is often infeasible to supply an entire table and request a fully prepared table as output.
-Therefore, a more effective strategy employs few-shot prompting to generate data prep programs.
 For instance, consider the table in Figure 3 with columns Date, Cyclist, Medal and Age, and question: “Which country has the most medals in total in February”.
-The few-shot prompting strategy prompts an LLM with a task description and a few demonstrations, and requests the LLM to generate Python programs as shown in Figure 3.
+A few-shot prompting strategy prompts an LLM with a task description and a few demonstrations, and requests the LLM to generate Python programs as shown in Figure 3.
 However, this LLM-based solution may encounter the following limitations when performing question-aware data prep for TQA.
-First, given the difficulties of understanding both NL questions and tables, it is challenging to determine which data prep operations are specifically required to meet the needs of each NL question, thus often resulting in false negatives (FNs) and false positives (FPs).
-Take the results in Figure 3 as an example: converting Age to a numerical format in code block ② is an FP, as it is irrelevant to the question.
-Conversely, the method fails to normalize the Date column before extracting the month in code block ③, which constitutes an FN, because it does not recognize the inconsistent Date formats.
-Second, due to input token limitations and challenges in long-context understanding (Li et al., 2024), it is not easy to fully understand all possible issues in a table, and thus may struggle to generate customized programs to correct issues.
+First, at the logical-operation level, given the inherent difficulties in understanding both NL questions and tables, it is challenging to accurately identify which data prep operations are specifically required to satisfy the needs of the NL question.
+This often leads to false negatives (FNs) and false positives (FPs).
+For example, as shown in Figure 3, converting the Age column to a numerical format in code block ② is a false positive (FP), as it is irrelevant to the question.
+In contrast, failing to normalize the Date column before extracting the month in code block ③ constitutes a false negative (FN), as the method ignores the inconsistency in Date formats.
+Second, at the physical-operation level, due to input token limitations and challenges in long-context understanding (Li et al., 2024), it is not easy to fully understand all possible issues in a table, and thus may struggle to generate customized programs to correct issues.
 For example, in Figure 3, the normalization of Medal in code block ① overlooks certain corner cases (e.g., “2*”), and the country extraction in code block ⑤ fails to handle “Dav.ITA”, which is formatted differently from other values.
-Recent methods, such as CoTable (Wang et al., 2024b) and ReAcTable (Zhang et al., 2024), can improve few-shot prompting by employing techniques like Chain-of-Thoughts (CoT) and ReAct.
-However, these methods remain insufficient to tackle the challenges, as they combine all diverse tasks, such as determining operations and implementing them at various granularities (table-level, column-level, and cell-level), within a single LLM agent.
+Recent methods, such as CoTable (Wang et al., 2024) and ReAcTable (Zhang et al., 2024), can improve few-shot prompting by employing techniques like Chain-of-Thoughts (CoT) and ReAct.
+However, these methods remain insufficient to tackle the challenges, as they combine all diverse tasks, such as determining operations and implementing them, within a single LLM agent.
 Existing studies (Ishibashi and Nishimura, 2024) have shown that a single LLM agent is often ineffective when tasked with handling a diverse range of operations, due to limited context length in LLMs and decreased inference performance with more input tokens.
-Figure 4. A Multi-agent Data Prep Framework for TQA (Left), which could be plugged to any other TQA methods (Right).
+Figure 4. A Multi-agent Data Prep Framework for TQA (Top), which could be plugged to any other TQA methods (Bottom).
 3. An Overview of AutoPrep
@@
   1. (1)
-    [Uncaptioned image] Planner Agent: the Planning stage. It guides the LLM to suggest high-level data prep operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$, which are tailored to specific question $Q$,
+    Planner Agent: the Planning stage. It guides the LLM to suggest logical data prep operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$, which are tailored to specific question $Q$,
   2. (2)
-    Multiple Programmer Agents (i.e., [Uncaptioned image] Augmenter): the Programming stage.
-It directs the LLM to generate low-level implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$.
+    Multiple Programmer Agents (i.e., Normalize): the Programming stage.
+It directs the LLM to generate physical implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$.
 Besides, it is also tasked for code debugging if any execution errors occur.
   3. (3)
-    An [Uncaptioned image] Executor Agent: the Executing stage. It executes the given low-level code and reports error messages if any code bugs occur.
-After that, an [Uncaptioned image] Analyzer agent extracts the answer from the prepared table.
+    An Executor Agent: the Executing stage. It executes the generated code and reports errors if any bugs occur.
+After that, an Analyzer agent extracts the answer from the prepared table.
 This agent can either use LLMs as black-boxes or leverage them for code generation, which is orthogonal to the question-aware data prep problem studied in this paper.
 For simplicity, we use a Text-to-SQL strategy that translates the question into an SQL query over the prepared table to obtain the final answer, as shown in Figure 4.
-Note that other strategies could also be used by the agent in a “plug-and-play” manner, which will be discussed in the experiments.
+Note that other strategies could also be used by the agent in a “plug-and-play” manner.
 Example 0.
-Figure 4 illustrates how AutoPrep supports data prep for question “Which country has the most medals in total in February?” posed over a table with columns Date, Cyclist, Medal and Age.
-(a) The Planning stage: The [Uncaptioned image] Planner suggests the following high-level operations to address the specific NL question:
-  • $T^{\prime}$[Country]=Augment(“Extract country code”, Cyclist) that extracts the country information from column Cyclist, producing a new Country column, in response to the “which country” part of the question.
-  • Normalize(“Case to INT”, Medal) that standardizes the value formats in the Medal column (e.g., removing quotation marks and asterisks) and then converts the strings to integers, as the question requires aggregating the “medals in total”;
+Figure 4 illustrates how AutoPrep supports data prep for an NL question posed over a table with $4$ columns.
+(a) The Planning stage: The Planner suggests the following high-level operations to address the specific NL question:
+  • $T^{\prime}$[Country]=Derive(“Extract country code”, Cyclist) that extracts the country information from column Cyclist, producing a new Country column, in response to the “which country” part of the question.
+  • Normalize(“Case to INT”, Medal) that standardizes the value formats in the Medal column (e.g., removing quotation marks and asterisks) and then converts the strings to integers, as the question requires “the most medals”;
   • Normalize(“Format date as %m-%d”, Date) standardizes the values in the Date column into a unified format to support the ‘in February’ condition in the question.
-  • $T^{\prime}=$Filter($T$, [Date, Country, Medal]) that filters out column Age, which is irrelevant to the question;
-(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e.,, [Uncaptioned image] Augment, [Uncaptioned image] Normalizer and [Uncaptioned image] Filter, corresponding to the Augment, Normalize and Filter operations, respectively.
+  • $T^{\prime}$ $=$ Filter([Date, Country, Medal], $T$) that filters out column Age, which is irrelevant to the question;
+(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e.,, Derive, Normalize and Filter.
 Each specialized Programmer focuses on generating executable code for its assigned operations.
 (c) The Executing stage: an Executor agent iteratively refines the generated code if any error occurs.
-After these stages, AutoPrep generates a prepared table $T^{*}$, which is then fed into an [Uncaptioned image] Analyzer agent to produce the answer ITA.
+After these stages, AutoPrep generates a prepared table $T^{*}$, which is then fed into an Analyzer agent to produce the answer ITA.
 The Planner Agent.
-The key challenge here is how to suggest question-aware operations that address specific NL questions.
-Specifically, even for the same table, different NL questions may require not only different data prep operations but also varying sequences of those operations.
-For example, the second Normalize operation might not be necessary if “in February” is not part of the question.
+The key challenge is how to suggest a logical plan that address specific NL questions.
+Specifically, even for the same table, different NL questions may require not only different logical operations but also varying sequences of those operations.
 To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent.
-This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch.
+This method translates the NL question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch.
 Compared to existing CoT methods (Wei et al., 2022), which break down questions into sub-questions, our approach better captures the relationships between questions and tables.
-More details of the CoC method are given in Section 4.
+More details of the method are given in Section 4.
 The Programmer Agents.
-The key challenge is that a given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial.
-For example, the Augmenter agent may generate an overly generic regular expression that extracts countries based on parentheses.
+The key challenge is that a given logical operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial.
+For example, the Derive agent may generate an overly generic regular expression that extracts countries based on parentheses.
 Unfortunately, this code fails to correctly process “Dav.
@@
 To tackle this challenge, we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions.
-We establish a search space of API functions for all high-level operations, create an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism.
 More details of our tool-augmented approach are discussed in Section 5.
-Note that our proposed AutoPrep framework is extensible.
+Remarks.
+Our proposed AutoPrep framework is extensible.
 When additional question-aware data prep operations are required, more specialized Programmer agents can be designed to handle them.
@@
 4. The Planner
-Given a question $Q$ posed over a table $T$, Planner aims to generate a logical plan that outlines a sequence of high-level operations needed to prepare the table for answering the question.
 4.1. A Direct Prompting Method
@@
 The inputs are a question $Q$, a table $T$, a set $\Sigma$ of specifications for each operation type, and an LLM $\theta$.
-Here, each specification $\sigma\in\Sigma$ describes the purpose of an operation type, e.g., “an Augment operation augments a new column for a table based on existing columns, in response to the specific needs of a question”.
-The output of the algorithm is a set $O$ of high-level operations like those shown in Figure 4, e.g., Normalize(“Cast to INT”, Medal).
+Here, each specification $\sigma\in\Sigma$ describes the purpose of an operation type, e.g., “an Derive operation creates a new column for a table based on existing columns, in response to the specific needs of a question”.
+The output of the algorithm is a set $O$ of high-level operations
 Example 0.
-Figure 5(a) shows an example of the aforementioned direct prompting method, which produces two high-level operations, Filter and Normalize.
-However, this logical plan might not be accurate as discussed as follows.
-(a) Incorrect operations: The Filter operation incorrectly filters out the Cyclist column because the column is not explicitly mentioned in the question.
-This operation is clearly disastrous, as the country information is implicitly embedded in the Cyclist column.
-(b) Missing operations: Observing the ground-truth in Figure 4, we see that the Augment operation on Cyclist is not generated, as the column has already been filtered out.
-Additionally, although the Normalize operation on Medal is generated, the Normalize operation on Date is missing.
+Figure 5(a) illustrates the direct prompting method, which produces two logical operations, Filter and Normalize. However, this logical plan might not be accurate, as discussed below.
+(a) Incorrect operations: The Filter operation retains the Country column simply because the question mentions “which country.” However, it fails to recognize that the original table does not actually contain a Country column.
+Worse yet, it incorrectly filters out the Cyclist column, merely because Cyclist is not explicitly mentioned in the question.
+This mistake is critical, as the country information is implicitly embedded within the Cyclist values.
+(b) Missing operations: Observing the ground-truth in Figure 4, we see that the Derive operation on Cyclist is not generated, as the column has already been filtered out.
+Moreover, although the Normalize operation on Medal is generated, the Normalize operation on Date is missing.
 This is because the phrase “the most” in the question suggests the need for type conversion of Medal, but there are no clues from the question to normalize Date, unless we observe from the table that the values are inconsistent.
@@
   • The first phase leverages an LLM to generate an SQL-like Analysis Sketch, outlining how the table should be transformed to produce the answer.
-  • The second phase iteratively examines different clauses in the Analysis Sketch, e.g., Date LIKE02-%’.
-For each clause, we associate the corresponding data in $T$ (e.g., values in column Date) with it to prompt the LLM for generating possible high-level operations (e.g., Normalize).
+  • The second phase iteratively examines different clauses in the Analysis Sketch, e.g., ORDER BY Date.
+At first, we use the syntax parser from Binder (Cheng et al., 2023) to extract all UDFs and operation clauses in the sketch and sort them by the execution order.
+Then, for each clause, we associate the corresponding data in $T$ (e.g., values in column Date) with it and provide several in-context learning demos to prompt the LLM for generating possible logical operations (e.g., Normalize).
 Compared with existing CoT methods (Wei et al., 2022), which simply break down questions into sub-questions, our approach is more effective for logical plan generation.
@@
 SQL-like Analysis Sketch. An Analysis Sketch $S$ is an SQL-like statement, which can be represented as follows.
-SELECT $A~{}|~{}{\tt agg}(A)~{}|~{}f(\{A_{i}\})$ FROM $T$
+SELECT $A~{}|~{}{\tt agg}(A)~{}|~{}f(A^{*}_{i},\{A_{i}\})$ FROM $T$
 WHERE ${\tt Pred}(A_{i})$ AND $\ldots$ AND ${\tt Pred}(A_{j})$
@@
 where ${\tt agg}(A)$ is an aggregation function (e.g., SUM and AVG) over column $A$, ${\tt Pred}(A_{i})$ denotes a predicate (i.e., filtering condition) over column $A_{i}$, such as Date LIKE02-%.
-Note that $f(\{A_{i}\})$ is a user-defined function (UDF) that maps existing columns $\{A_{i}\}$ in the table into a new column.
-Figure 5(b) shows an example Analysis Sketch with a UDF function that specifies a new column Country, i.e., $f({\tt Cyclist})$ AS Country.
+Note that $f(A^{*}_{i},\{A_{i}\})$ is a user-defined function (UDF) that maps existing columns $\{A_{i}\}$ in the table into a new column $A^{*}_{i}$.
+Figure 5(b) shows an example Analysis Sketch with a UDF that specifies a new column Country, i.e., $f({\tt Country},{\tt Cyclist})$.
 Obviously, this UDF is introduced because the Analysis Sketch contains a GROUP BY Country clause.
+Logical Operation Generation.
+Given the constructed Analysis Sketch, we orchestrate the logical operations and determine their execution order by translating the sketch into actionable steps.
+Specifically, we leverage the syntax parser from Binder (Cheng et al., 2023) to extract all UDFs and operation clauses from the sketch and sort them according to their intended execution sequence from the parser.
+Based on this parsed order, we then generate the corresponding logical operations, as illustrated below.
 Example 0.
@@
 (b) Phase II - Operation Generation: In this phase, the algorithm iteratively examines the clauses in Analysis Sketch $S$ as follows.
-  • $f({\tt Cyclist})$ AS Country: this clause and relevant columns are used to prompt the LLM to generate an Augment operation.
-  • ${\tt SUM}({\tt Medal})$: this clause and the values in column Medal are used to prompt the LLM to generate a Normalize operation.
-  • Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize.
+(1) $f({\tt Country,Cyclist})$: this clause and relevant columns are used to prompt the LLM to generate an Derive operation.
+(2) ${\tt SUM}({\tt Medal})$: this clause and the values in column Medal are used to prompt the LLM to generate a Normalize operation.
+(3) Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize.
 Finally, the algorithm generates a Filter operation that only selects columns relevant to the Analysis Sketch.
+Note that the Filter operation removes irrelevant columns, whereas the Derive operation creates new columns from existing ones.
 5. The Programmer & Executor
@@
 However, the above method may have limitations, as the generated code may be overly generic and unable to effectively address the heterogeneity challenge of the tables.
-Specifically, many tables originate from web sources or real-world scenarios, where values have diverse syntactic (e.g., “19-Oct” and “9/14” in $T_{6}$) or semantic formats (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code customized to these tables.
-For example, consider table $T_{4}$ in Figure 2(b).
-Given a Normalize operation to standardize country formats, a generic function from an existing Python library may not be sufficient to transform country names (e.g., “Italia”) to their ISO codes (e.g., “ITA”).
-In such cases, customized functions, such as the clean_country function from an external library dataprep (Peng et al., 2021), are needed to address specific requirements of code generation.
+Specifically, many tables originate from web sources or real-world scenarios, where values have diverse syntactic (e.g., “19-Oct” and “9/14”) or semantic formats (e.g., “ITA” and “Italia”), making it difficult to generate code customized to these tables.
+For example, given a Normalize operation to standardize country formats, a generic function from an existing Python library may not be sufficient to transform country names (e.g., “Italia”) to their ISO codes (e.g., “ITA”).
+In such cases, customized functions, such as the clean_country function from an external library dataprep (Peng et al., 2021a), are needed to address specific requirements of code generation.
 To address these limitations, we introduce a tool-augmented method for the Programmer agents, enhancing the LLM’s code generation capabilities by utilizing pre-defined API functions, referred to as the function pool in this paper.
@@
 Figure 6 provides an overview of our tool-augmented method for physical plan generation and execution.
-Given a table $T$ and a set $O=\{o_{1},\ldots,o_{|O|}\}$ of high-level operations, the method iteratively generates and executes low-level code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$.
-Specifically, in the $i$-th iteration, given a high-level operation $o_{i}$ (e.g., Augment) and an intermediate table $T_{i}$, the method generates low-level code and executes it to produce an updated table $T_{i+1}$ through the following two steps.
+Given a table $T$ and a set $O=\{o_{1},\ldots,o_{|O|}\}$ of logical operations, the method iteratively generates and executes physical executable code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$.
+Specifically, in the $i$-th iteration, given a high-level operation $o_{i}$ (e.g., Derive) and an intermediate table $T_{i}$, the method generates physical code and executes it to produce an updated table $T_{i+1}$ through the following two steps.
 Function Selection and Argument Inference.
-The method first prompts the LLM to select a specific function from a function pool $\mathcal{F}$ corresponding to the operation type (e.g., $F_{\tt Aug}$) and infer the arguments (e.g., regular expression) for the selected function.
-For instance, given the Augment operation over table $T_{i}$ shown in Figure 6, the LLM selects a function from the pool $\mathcal{F}_{\tt Aug}$ designed specifically for Augment, obtaining an extract function with two arguments: column and func.
+The method first prompts the LLM to select a specific function from a function pool $\mathcal{F}$ corresponding to the operation type (e.g., $F_{\tt Der}$) and infer the arguments (e.g., regular expression) for the selected function.
+For instance, given the Derive operation over table $T_{i}$ shown in Figure 6, the LLM selects a function from the pool $\mathcal{F}_{\tt Der}$ designed specifically for Derive, obtaining an extract function with two arguments: column and func.
 The LLM then generates preliminary code for these arguments, assigning Cyclist to the column argument and generating a lambda function with a specific regular expression for argument func.
-Note that, in addition to selecting functions from the corresponding pool, our method may also prompt the LLM to write specific code for an operation if no existing function is suitable to meet the operation’s requirements.
+Note that, in addition to selecting functions from the corresponding pool, our method may also prompt the LLM to write specific code for an operation (denoted as $\mathtt{CustomCode}$) if no existing function is suitable to meet the operation’s requirements.
 Code Execution and Debugging.
@@
 This section presents a search space of function pools for different types of high-level operations formalized in the current AutoPrep.
-Function pool for $\mathtt{Augment}$.
-We define a specific Programmer agent, i.e., Augmenter, to generate low-level code for the operation type $\mathtt{Augment}$ from the following function pool $\mathcal{F}_{\mathtt{Aug}}$.
-  • $\mathtt{extract}$.
+Function Pool for $\mathtt{Derive}$.
+We define a specific Programmer agent, i.e., Derive, to generate low-level code for the operation type $\mathtt{Derive}$ from the following function pool $\mathcal{F}_{\mathtt{Der}}$.
+(1) $\mathtt{extract}$.
 This function extracts a substring from a column to generate a new column.
@@
 )’,x).group(1).
-  • $\mathtt{calculate}$.
+(2) $\mathtt{calculate}$.
 The function generates a new column through arithmetic operations of existing columns.
 It has two arguments: ${\tt columns}$, representing a list of source columns and ${\tt func}$, representing a lambda function with input to be a dictionary of values, since we may perform calculations on multiple columns.
-For example, to generate a new column GrowthRate for $T_{2}$ in Figure 2(a), we set ${\tt columns}$ to [’2012’, ’2013’] and ${\tt func}$ to a lambda function lambda x: (x[’2013’]-x[’2012’])/x[’2012’].
-  • $\mathtt{map\mbox{-}to\mbox{-}boolean}$ and $\mathtt{concatenate}$.
+For example, to generate a new column GrowthRate for $T_{2}$ in Figure 2(a), we set ${\tt columns}$ to [‘2012’, ‘2013’] and ${\tt func}$ to a lambda function lambda x: (x[‘2013’]-x[‘2012’])/x[‘2012’].
+(3) $\mathtt{map\mbox{-}to\mbox{-}boolean}$ and $\mathtt{concatenate}$.
 Both functions generate a new column from multiple columns.
@@
 Both operators take two arguments, ${\tt columns}$ and ${\tt func}$, similar to those in the $\mathtt{calculate}$ operator.
-  • $\mathtt{infer}$.
-This function directly uses LLMs to deduce values that could not be processed by the functions above.
-It takes a list of source columns and the failed target values as input, learning from demonstrations of successfully processed values to directly output the target values.
+(4) $\mathtt{infer}$.
+This function uses LLMs to deduce values that could not be processed by the above functions.
+It takes a list of columns and the failed target values as input, learning from demonstrations of successfully processed values to output the target values.
 Function Pool for $\mathtt{Normalize}$.
-We define a specific Programmer agent, i.e., Normalizer, to generate low-level code for the operation type $\mathtt{Normalize}$ from the following function pool $\mathcal{F}_{\mathtt{Norm}}$.
-  • $\mathtt{to\mbox{-}numerical}$.
+We define a specific Programmer agent, i.e., Normalize, to generate low-level code for the operation type $\mathtt{Normalize}$ from the following function pool $\mathcal{F}_{\mathtt{Norm}}$.
+(1) $\mathtt{to\mbox{-}numerical}$.
 This function standardizes a column to a numerical type, such as int or float.
 To use it, two arguments need to be specified: ${\tt column}$, representing the source column, and ${\tt func}$, representing a lambda function.
-For example, given the table $T_{5}$ in Figure 2(b), we use this function to standardize the Females column to integers, with ${\tt column}$ as Females and ${\tt func}$ as lambda x: int(x.replace(’,’, ’’)).
-  • $\mathtt{format\mbox{-}datetime}$.
+For example, we can use this function to standardize the Time column to integers, with ${\tt column}$ as Time and ${\tt func}$ as the lambda function.
+(2) $\mathtt{format\mbox{-}datetime}$.
 The function standardizes the format of columns with DateTime values.
@@
 The ${\tt format}$ argument specifies the desired format for standardization.
-Given the table $T_{6}$ in Figure 2(b), to make the values in the Date column comparable, we use the $\mathtt{format\mbox{-}datetime}$ function with ${\tt column}$ set to Date and ${\tt format}$ set to “%m-%d”.
-  • $\mathtt{clean\mbox{-}string}$.
+For example, to make the values in the Date column comparable, we use the $\mathtt{format\mbox{-}datetime}$ function with ${\tt column}$ set to Date and ${\tt format}$ set to “%Y-%m-%d”.
+(3) $\mathtt{clean\mbox{-}string}$.
 To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$.
@@
 The $\tt trans\_dict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$.
-Given the original table $T_{4}$ in Figure  2(b), we set $\tt trans\_dict$ to {’Italia’: ’ITA’}.
-  • $\mathtt{infer}$. Similar to the one for the $\mathtt{Augment}$ operation, this function uses LLMs to deduce values that could not be processed by the functions above.
-Function pool for $\mathtt{Filter}$.
-We design the Normalizer agent to utilize a function pool $\mathcal{F}_{\mathtt{Filter}}$ with a single function, $\mathtt{filter\mbox{-}columns}$.
+(4) $\mathtt{infer}$. Similar to the one for the $\mathtt{Derive}$ operation, this function uses LLMs to deduce values that could not be processed by the functions above.
+Function Pool for Filter.
+We design the Filter agent to utilize a function pool $\mathcal{F}_{\mathtt{Filter}}$ with a single function, $\mathtt{filter\mbox{-}columns}$.
 This function has one argument, ${\tt rel\_columns}$, which represents a list of question-related column names.
-Discussions.
-Our search space is extensible, allowing additional functions or external APIs to be easily incorporated into the function pools, such as specific data prep functions for various semantic types developed in the dataprep library (Peng et al., 2021).
+Remarks.
+The design of $\mathtt{CustomCode}$ and $\mathtt{infer}$ operations enables AutoPrep to generalize to unseen tasks.
+For example, when encountering missing or clearly erroneous values, AutoPrep can directly prompt an LLM to infer plausible replacements.
+Moreover, our search space is extensible: additional functions or external APIs can be easily integrated into the function pool, including specialized data preparation functions for various semantic types, such as those provided by the dataprep library (Peng et al., 2021a).
 6. Experiments
 6.1. Experimental Setup
-Datasets.
-We evaluate our proposed framework AutoPrep using two well-adopted benchmarking datasets WikiTQ (Pasupat and Liang, 2015) and TabFact (Chen et al., 2019) and a very recent benchmark TableBench (Wu et al., 2024).
-The statistics of the datasets are shown in Table 1.
-(1) WikiTQ contains complex questions annotated by crowd workers based on diverse Wikipedia tables.
+Datasets. We evaluate AutoPrep using the following well-adopted datasets. The statistics of the datasets are shown in Table 1.
+(1) WikiTQ (Pasupat and Liang, 2015) contains complex questions annotated by crowd workers based on diverse Wikipedia tables.
 WikiTQ comprises 17,689 question-answer pairs in the training set and 4,344 in the test set with significant variation in table size.
-The answers in WikiTQ can be categorized into three forms: (i) string from a cell in the table, (ii) string from analyzing, and (iii) list of cells in the table.
-(2) TabFact provides a collection of Wikipedia tables, along with manually annotated NL statements.
+(2) TabFact (Chen et al., 2019) provides a collection of Wikipedia tables, along with manually annotated NL statements.
 One requires to deduce the relations between statements and tables as “true” (a statement is entailed by a table) or “false” (a statement if refuted by a table).
@@
 The answers are binary with nearly equal proportions.
-(3) TabBench provides an evaluation of table question answering capabilities within four major categories, including multi-hop fact checking (FC), multi-hop Numerical Reasoning (NR), Trend Forecasting and Chart Generation.
-We evaluate AutoPrep on TabBench-FC and TabBench-NR, where one requires to conduct more complex and multi-hop reasoning over tables for answering a question.
-The answers in TabBench could be a string or list of items.
-Baselines. We consider the following two categories of baselines:
+(3) TabBench (Wu et al., 2024) provides an evaluation of TQA capabilities within four major categories, including multi-hop fact checking (FC), multi-hop Numerical Reasoning (NR), Trend Forecasting and Chart Generation.
+We evaluate the generalization of AutoPrep on FC and NR, where one requires to conduct more complex and multi-hop reasoning over tables for answering a question.
+Specifically, we directly use the prompts in WikiTQ to evaluate the generalization capabilities of AutoPrep on TabBench.
+Baselines. We consider the following three categories of baselines:
 (1) TQA Methods w/o Data Prep directly generate the answer or utilize programs to extract the answer from the table without considering data prep. We implement four primary TQA baselines.
@@
 Chain-of-Thought (CoT) (Wei et al., 2022) prompts LLMs to generate the reasoning process step-by-step before generating the final answer. We implement CoT with the prompt provided by (Chen, 2022).
-NL2SQL (Rajkumar et al., 2022) first translates the question into a SQL program and then executes it to get the final answer from the table for the question.
+NL2SQL (Rajkumar et al., 2022) first translates the question into an SQL program and then executes it to get the final answer from the table for the question.
 We use the prompt from (Cheng et al., 2023) to implement NL2SQL.
@@
 TableBench $886$ 2$\sim$212 2$\sim$20 string / list (31.49%)
-(2) TQA Methods with Data Prep. We also investigate four SOTA TQA methods that consider data prep tasks in their question answering process.
-ICL-Prep.
-uses few-shot in-context learning (ICL) demonstrations to guide LLMs in generating programs for data prep, as shown in Figure 3.
-Subsequently, we employ NL2SQL to extract answers from the cleaned tables.
+(2) Data Prep Baselines. We consider the following data prep baselines. For each baseline, we first perform data prep and then use the above NL2SQL to extract answers from the prepared tables.
+Offline DataPrep (Off-Prep) performs offline data prep operations for all tables in TQA.
+To this end, we have surveyed and consolidated the offline data prep operations, such as data cleaning, value normalization and column renaming, adopted by current SOTA TQA methods (Yao et al., 2022; Cheng et al., 2023; Khot et al., 2022; Wang et al., 2024; Zhu et al., 2024) to construct a comprehensive offline data prep pipeline, by utilizing the Pandas library (McKinney, 2010) and the popular DataPrep toolkit (Peng et al., 2021b).
+ICL-Prep uses few-shot ICL demonstrations to guide LLMs in generating Python programs for data prep, as shown in Figure 3.
+(3) SOTA TQA Methods with Data Prep. We investigate four SOTA TQA methods considering data prep tasks for TQA task.
 Dater (Ye et al., 2023) addresses the TQA task by decomposing the table and question.
@@
 We utilize the original code provided by (cod, 2023a).
+AutoTQA (Zhu et al., 2024) uses a multi-agent framework for TQA.
+Since the official code for AutoTQA is not publicly available, we reimplement it under the guidance of the authors (see our repository (https://github.com/fmh1art/AutoPrep/src/model/autotqa).
+Given the high time and API costs, we evaluate AutoTQA on a sampled subset of 500 instances from each of WikiTQ and TabFact.
 ReAcTable (Zhang et al., 2024) uses the ReAct paradigm to extract relevant data from the table using Python or SQL code generated by LLMs.
@@
 Notice that the original code does not include prompts for TabFact, we generate it based the WikiTQ prompt.
-Chain-of-Table (CoTable) (Wang et al., 2024b) enhances the table reasoning capabilities of LLMs by predefining several common atomic operations (including data prep operations) that can be dynamically selected by the LLM.
+Chain-of-Table (CoTable) (Wang et al., 2024) enhances the table reasoning capabilities of LLMs by predefining several common atomic operations (including data prep operations) that can be dynamically selected by the LLM.
 These operations form an “operation chain” that represents the reasoning process over a table and can be executed either via Python code or by prompting the LLM.
@@
 + AutoPrep 66.09 $\uparrow 13.26$ 87.85 $\uparrow 17.64$ 64.75 $\uparrow 11.85$ 84.19 $\uparrow 19.48$ 63.72 $\uparrow 11.92$ 85.72 $\uparrow 10.57$ 68.72 $\uparrow 11.86$ $85.33$ $\uparrow 5.24$
-Evaluation Metrics.
-We adopt the evaluator in Binder (Cheng et al., 2023) to avoid the situation where program executions are likely to be semantically correct but fail to match the golden answer exactly.
-For example, when the SQL outputs 1/0 for yes/no questions, the Binder evaluator will conduct a pre-matching check and translate it to boolean values first.
-It will avoid mismatches of correct prediction and golden answer, such as pairs like 7.45 and 7.4499, 4 years and 4.
-Afterwards, we adopt accuracy as our evaluation metric.
+Evaluation Metrics. We consider both accuracy and cost.
+Accuracy. We adopt the evaluator from Binder (Cheng et al., 2023) to address cases where program executions are semantically correct but do not exactly match the golden answers.
+Cost.
+We measure both the time and API cost for all methods.
+For time cost, we ensure a stable network environment and record the end-to-end processing time for each method on a single TQA instance.
+For API cost, we follow the official pricing guidelines (https://api-docs.deepseek.com/quick_start/pricing/) and calculate the cost based on the default LLM backbones used.
 Backbone LLMs.
-We evaluate our method using representative LLMs as backbones.
+We evaluate our methods using representative LLMs as backbones.
 For closed-source LLMs, we select DeepSeek (Guo et al., 2024) (DeepSeek-V2.5-Chat) and GPT3.5 (Brown et al., 2020) (GPT3.5-Turbo-0613).
@@
 We provide detailed prompts of each component in AutoPrep in our technical report (tec, 2025) due to the space limit.
-For a fair comparison, we set the maximum token input of all methods as 8192.
+Moreover, for fair comparison, we set the maximum token input of all methods as 8192.
 Moreover, we set the temperature parameter of all methods to 0.01 for reproducibility.
 6.2. Improvement of Data Prep for TQA
-Exp-1: Impact of question-aware data prep on TQA performance.
-We integrate AutoPrep into our four TQA baselines w/o data prep, in order to investigate the impact of question-aware data prep on TQA performance.
-The results are reported in Table 2.
+Exp-1: Impact of question-aware data prep on TQA performance. We integrate AutoPrep into our four TQA baselines w/o data prep, and report the results in Table 2.
 As demonstrated, integrating AutoPrep significantly improves the performance of all evaluated methods.
@@
 This significant improvement is attributed to the sensitivity of NL2SQL and NL2Py to data incompleteness and inconsistency, which can lead to erroneous outcomes when performing operations on improperly formatted data.
-Thus, data prep operations, such as Augment and Normalize can solve these cases and improve the overall results.
+Thus, data prep operations, such as Derive and Normalize can solve these cases and improve the overall results.
 Moreover, End2End and CoT methods also show considerable performance gains.
-For example, End2End shows an average accuracy gain of 5.27 and 1.35 on WikiTQ and TabFact respectively.
-These improvements are largely due to AutoPrep ’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers directly from tables.
-Finding 1: The integration of AutoPrep improves the accuracy of TQA methods w/o data prep by 12.22 points on WikiTQ and 13.23 points on TabFact, respectively.
+These improvements are largely due to AutoPrep ’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers from tables.
 Since “NL2SQL + AutoPrep” achieves the best accuracy in most cases, we take its results as default for further comparison.
@@
  WikiTQ  TabFact  WikiTQ  TabFact  WikiTQ  TabFact  WikiTQ  TabFact
+Off-Prep 55.3255.32{55.32}55.32 $81.67$ $56.86$ $\underline{81.52}$ $53.02$ $75.40$ $58.01$ $82.31$
 ICL-Prep 56.5456.54{56.54}56.54 $80.53$ $55.71$ $73.91$ $50.05$ $75.20$ $57.00$ $80.14$
@@
 Binder $56.81$ $82.81$ 56.7456.74{56.74}56.74 79.1779.17{79.17}79.17 $50.51$ $78.16$ $55.43$ 81.7281.72{81.72}81.72
+AutoTQA* 60.8060.80{60.80}60.80 $84.40$ $58.40$ $80.60$ $58.80$ $82.40$ $62.40$ 83.0083.00{83.00}83.00
 ReAcTable 64.1364.13{64.13}64.13 85.7185.71{85.71}85.71 $51.80$ $72.80$ 58.0158.01{58.01}58.01 80.0080.00{80.00}80.00 60.1560.15{60.15}60.15 $81.67$
-CoTable $\underline{64.53}$ $\underline{86.22}$ $\underline{59.94}$ $\underline{80.20}$ $\underline{62.22}$ $\underline{85.62}$ $\underline{64.41}$ $\underline{83.20}$
+CoTable $\underline{64.53}$ $\underline{86.22}$ $\underline{59.94}$ 80.2080.20{80.20}80.20 $\underline{62.22}$ $\underline{85.62}$ $\underline{64.41}$ $\underline{83.20}$
 AutoPrep 66.09 87.85 64.75 84.19 63.72 85.72 68.72 85.33
-Exp-2: Comparison of AutoPrep with previous SOTA TQA methods with data prep.
-We compare AutoPrep with TQA methods that consider data prep tasks in their question answering process, using a single LLM agent.
-The results are reported in Table 3.
-The results clearly demonstrate that AutoPrep achieves the new SOTA performance on both the WikiTQ and TabFact datasets across all LLM backbones.
-In particular, AutoPrep outperforms ICL-Prep with a significant accuracy improvement by 11.00 and 8.33 on WikiTQ and TabFact on average, respectively.
-Other baselines, except Dater, generally outperform ICL-Prep by introducing specific technical optimizations to data prep.
-For instance, ReAcTable applies the ReAct paradigm to iteratively generate data prep operations rather than producing all code at once, and it explicitly considers data prep tasks (e.g., filtering and augmentation).
-Despite this, AutoPrep still outperforms the second best method by considerable improvement.
-Specifically, when compared with CoTable, AutoPrep achieves better performance, with an average accuracy improvement by 3.05 and 1.96 on WikiTQ and TabFact, respectively.
-These improvements are primarily attributed to our multi-agent framework, which effectively tackles the question-aware data prep challenge.
-The experimental results indicate that data prep is inherently complex and is hard to be resolved by a single, one-size-fits-all solution; instead, a more effective approach is to develop specialized methods for each type.
-A centralized module can then determine which operations to perform and assign them accordingly.
-Moreover, AutoPrep covers a broader range of data prep operations than these state-of-the-art TQA methods, thereby addressing gaps (e.g., normalization tasks) not fully explored by previous solutions.
-Finding 2: AutoPrep outperforms the the previous SOTA TQA methods with data prep by 3.05 points on WikiTQ and 1.96 points on TabFact, respectively,
-In addition, we find that DeepSeek achieves the better overall performance at lower API costs. Thus, we select DeepSeek as our default backbone LLM for subsequent experiments.
-Exp-3: Evaluation of AutoPrep on tables with various sizes.
-To further investigate AutoPrep, we analyze its performance across tables of varying sizes.
-While all tables in TabFact are small, we categorize the tables in WikiTQ into Small (fewer than 2048 tokens), Medium (2048 to 4096 tokens), and Large (more than 4096 tokens), resulting in 4040 Small tables, 200 Medium tables, and 104 Large tables.
-The results are reported in Figure 7.
-We observe that all baselines exhibit unstable performance, particularly on larger tables.
-For instance, although CoTable achieves the highest average accuracy among previous SOTA methods, it suffers an accuracy drop of 11.78 when reasoning over large tables.
-Likewise, ReAcTable shows relatively stable performance on medium tables (dropping by 2.15), yet its performance on large tables remains unsatisfactory (dropping by 17.53).
-(a) Small
-(b) Medium
-(c) Large
-Figure 7. Comparison of AutoPrep and other data prep methods on tables with different sizes on the WikiTQ dataset.
-In contrast, AutoPrep achieves the highest and most stable performance across tables of varying sizes.
-When processing medium and large tables, the accuracy of AutoPrep drops by 2.29 and 3.79.
-The key to its stability is that it employs specialized agents for each data prep task, mitigating the issue of exceeding prompt length limits.
-Furthermore, each agent generates program-based operators to handle data prep on the entire table, thereby avoiding information loss that occurs when large tables are cropped.
-Finding 3: By employing multiple agents and program-based operations, AutoPrep maintains stable performance as table size grows, ensuring that each data prep task is handled effectively without overwhelming a single model.
-6.4. Evaluation on Generalization
-Exp-4: Evaluating generalization on unseen datasets.
-As discussed previously, we evaluate the generalization capabilities of AutoPrep on two TabBench datasets.
+Exp-2: Comparison of AutoPrep with data prep baselines.
+We compare AutoPrep against two representative baselines: a traditional offline data prep method Off-Prep and an LLM-based in-context learning data prep method ICL-Prep.
+As shown in Table 3, AutoPrep consistently outperforms both baselines on WikiTQ and TabFact.
+Specifically, AutoPrep improves over Off-Prep by 10.02 and 5.55 on average, respectively.
+The performance gap can be attributed to the fact that Off-Prep lacks question guidance, making it ineffective in addressing question-specific issues such as missing semantics and irrelevant columns.
+Moreover, predefined normalization routines struggle with table heterogeneity, whereas AutoPrep can generate specialized code for such cases via LLMs.
+Similarly, AutoPrep outperforms ICL-Prep by 11.00 and 8.33 on average on the two datasets.
+This highlights that a multi-agent framework is essential for producing comprehensive data prep requirements and precise, executable programs.
+Exp-3: Comparison of AutoPrep with previous SOTA TQA methods with data prep.
+We compare AutoPrep with TQA methods that integrate data prep tasks in their question-answering process.
+The results, shown in Table 3, highlight that AutoPrep sets a new SOTA performance on both the WikiTQ and TabFact datasets across LLM backbones.
+Although CoTable achieves the best overall performance among existing SOTA methods, AutoPrep outperforms it with an impressive average improvement of 3.05 on WikiTQ and 1.96 on TabFact.
+These improvements are largely attributed to our multi-agent framework, which effectively addresses the question-aware data prep challenges.
+Furthermore, when compared to AutoTQA, AutoPrep shows significant gains of 5.81 on WikiTQ and 2.77 on TabFact.
+This is because AutoTQA lacks comprehensive data prep, leading to execution errors or incorrect answers.
+The results demonstrate that data prep is inherently complex and cannot be solved with a one-size-fits-all solution.
+Instead, a more effective strategy involves specialized LLM-based agents for each type of data prep task, coordinated by a centralized planning agent.
+Moreover, AutoPrep covers a broader range of data prep operations, filling gaps (e.g., Derive and Normalize) that previous solutions have not fully addressed.
+Moreover, we also evaluate AutoPrep on tables with various sizes.
+We find that, by employing multiple agents and program-based operations, AutoPrep maintains stable performance as table size grows, ensuring that each data prep task is handled effectively without overwhelming a single model.
+More details on the experimental results and analysis can be found in our technical report (tec, 2025).
+As DeepSeek achieves the best accuracy at lower cost, we select it as our default backbone LLM for subsequent experiments.
+6.4. Evaluation on Efficiency
+Figure 7.
+Efficiency evaluation.
+The $x$-axis and $y$-axis represent the time cost (in seconds) and monetary cost (in dollars) for processing a single instance, respectively.
+The color intensity of each scatter point reflects the overall accuracy.
+Exp-4: How efficient is AutoPrep compared with other methods?
+To evaluate the efficiency of AutoPrep, we compare it with other TQA methods in terms of time cost, monetary cost, and accuracy.
+The results for both small and large tables are shown in Figure 7.
+As shown, AutoPrep achieves the best overall performance with insignificant time and monetary costs.
+Specifically, its cost is lower than all methods except Off-Prep and ICL-Prep.
+However, AutoPrep significantly outperforms these two baselines in accuracy, achieving improvements of at least 9.36 and 5.77 points on small and large tables, respectively.
+Furthermore, as table size increases, AutoPrep maintains a much more stable time and cost overhead compared to other state-of-the-art TQA methods.
+This efficiency stems from the program-based data prep mechanism of AutoPrep, which avoids high latency and API expenses associated with directly processing tables via LLMs.
+For instance, in the column derivation task on Table $T_{1}$ (Figure 2(a)), CoTable invokes LLMs to generate a full list of country codes.
+In contrast, AutoPrep only generates a physical operator $\mathtt{extract}$ and executes it to produce the column, which results in significantly lower time and monetary costs, particularly on large tables.
+6.5. In-Depth Exploration of System Capabilities
+Exp-5: Evaluating generalization on unseen datasets.
+As discussed in Section 6.1, we evaluate the generalization capabilities of AutoPrep on the TabBench datasets.
 Specifically, we directly use the designed prompting strategies on the WikiTQ dataset, and examine whether these strategies can be generalized to TabBench.
 As shown in Figure 8, AutoPrep achieves the highest overall accuracy among all methods.
-Specifically, compared with the second best method ReAcTable, AutoPrep improves by 5.28, indicating the strong generalization capability of our method.
+Specifically, compared with the second best method ReAcTable, AutoPrep improves by 5.28, indicating the strong generalization capabilities of our method.
 The main reason is that AutoPrep utilizes tool-augmented method for physical plan generation and execution, generalizing well on unseen datasets.
-Finding 4: Utilizing tool-augmented execution enhances the generalization capabilities of AutoPrep on unseen datasets.
-(a) Numerical Reasoning
-(b) Fact Checking
+Table 4. Evaluating Extensibility for New Operation Types.
+Method TransTQ WikiTQ
+ origin  after
+ICL-Prep 58.92 56.54 55.00
+ReAcTable 51.04 64.13 63.03
+AutoTQA 59.75 60.80 61.20
+AutoPrep 68.88 66.09 66.28
+Exp-6: Evaluating extensibility on datasets requiring new logical operation types.
+To further evaluate the extensibility of AutoPrep, we evaluate it on a more complex dataset involving logical operations not originally supported.
+Since no existing TQA dataset presents sufficiently complex data quality issues, we construct a new dataset—TransTQ with the assistance of LLMs, which is available on our GitHub repository.
+Specifically, we prompt an LLM to identify tables from WikiTQ where inverse transformation operations (e.g., pivot and stack (Li et al., 2023)) can be applied.
+These operations are then executed to generate non-relational table formats, introducing new challenges for TQA on table transformation.
+As shown in Table 4, without explicit guidance for handling table transformations, all baseline TQA methods exhibit poor accuracy.
+In contrast, after incorporating a new logical operation Transform, AutoPrep achieves the best performance, outperforming the second-best method by 9.31 points.
+Moreover, we also validate that adding this new operation does not degrade the performance of AutoPrep on the original WikiTQ dataset, confirming that its multi-agent architecture is well extensible to support the seamless integration of new data prep functionalities.
 Figure 8. Evaluating generalization on the TabBench dataset.
 6.6. Ablation Studies
-Exp-5: Evaluation on the Planner agent.
+Exp-7: Evaluation on the Planner agent.
 We compare two Planner variants with different high-level operation suggestion methods, namely Direct Prompting and our proposed Chain-of-Clauses method, and report the results in Table 5(a).
 We observe that Chain-of-Clauses outperforms Direct Prompting by 7.92, 4.50, 5.27 in accuracy on WikiTQ, TabFact and TabBench respectively.
-This indicates the superiority of our proposed method in generating more accurate high-level operations.
+This indicates the superiority of our proposed method in generating more accurate logical operations.
 Moreover, we find that the performance improvement on WikiTQ is more significant than that on other datasets.
-This is because the WikiTQ dataset has relatively large tables and complex questions, which could make the high-level operation suggestion problem more challenging to be solved using Direct Prompting.
-Finding 5: Our proposed Chain-of-Clauses method is more effective in suggesting high-level operations.
-(a) Case one
-(b) Case two
-Figure 9. Two case studies selected from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
-Exp-6: Evaluation on the Programmer Agents. We change the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default.
+This is because the WikiTQ dataset has relatively large tables and complex questions, which could make the logical operation suggestion problem more challenging to be solved using Direct Prompting.
+Exp-8: Evaluation on the Programmer Agents. We evaluate the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default.
 As illustrated in Table 5(b), our proposed tool-augmented method achieves better performance compared with the code generation method.
-Considering the high-level operations input to the Programmer agents are the same, we can conclude that selecting a function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations.
+Considering the logical operations input to the Programmer agents are the same, we can conclude that selecting a function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations.
 Moreover, when using demonstrations constructed from same table-question pairs for the prompt of programmer agents, Tool-augmented method can save 18.07% input tokens for low-level operation generation.
@@
 The probability of bugs in our tool-augmented method is greatly reduced (e.g., from 4.51% to 1.50%), demonstrating its effectiveness.
-Finding 6: Our tool-augmented method performs better in both accuracy and cost, with fewer program errors.
 Table 5. Experimental Results of Ablation Studies.
@@
 - Filter $62.78$ ($-3.31$) $84.98$ ($-2.87$) $47.67$ ($-2.43$)
-- Augmenter $61.79$ ${(-4.30)}$ $85.67$ ($-2.18$) $45.44$ ($-4.66$)
-- Normalizer $62.02$ ($-4.07$) $83.79$ ${(-4.06)}$ $41.58$ ${(-8.52)}$
-Exp-7: Contribution of each agent in AutoPrep. We ablate each Programmer agent including Filter, Augmenter and Normalizer and compare the performance with AutoPrep.
-As shown in Table 4(c), for WikiTQ, without column augmentation, the accuracy drops the most (4.30).
-For TabFact and TabBench, the Normalizer matters the most with an accuracy drop by 4.06 and 8.52.
+- Derive $61.79$ ${(-4.30)}$ $85.67$ ($-2.18$) $45.44$ ($-4.66$)
+- Normalize $62.02$ ($-4.07$) $83.79$ ${(-4.06)}$ $41.58$ ${(-8.52)}$
+Exp-9: Contribution of each agent in AutoPrep. We ablate each Programmer agent including Filter, Derive and Normalize and compare the performance with AutoPrep.
+As shown in Table 5(c), for WikiTQ, without column derivation, the accuracy drops the most (4.30).
+For TabFact and TabBench, the Normalize matters the most with an accuracy drop by 4.06 and 8.52.
 This is because that WikiTQ has more instances requiring string extraction or calculation to generate new columns for answering the question, while normalization is a primary issue in TabFact and TabBench.
 Moreover, for all datasets, each agent plays an essential role in data preparation for TQA, which brings accuracy improvement by at least 2.43, 2.18 and 4.06.
-Finding 7: Solving three types of data preparation tasks in TQA methods in a specialized manner improves the downstream TQA performance.
-6.6. Case Studies
+7. Related Work
+Tabular Question Answering.
+Most of the SOTA solutions for TQA rely on LLMs (Brown et al., 2020; Guo et al., 2024), as TQA requires NL understanding and reasoning over tables.
+There are two types of methods for TQA named Direct Prompting (Chen, 2022; Wei et al., 2022) and Code Generation (Rajkumar et al., 2022).
+Previous TQA methods, like Dater (Ye et al., 2023), Binder (Cheng et al., 2023), CoTable (Wang et al., 2024) and ReAcTable (Zhang et al., 2024) also consider data prep in their question answering process.
+Specifically, Dater (Ye et al., 2023) prompts LLMs to select relevant columns related to answering the question, targeted at solving filtering tasks.
+Binder (Cheng et al., 2023) proposes to integrate SQL with an LLM-based API to incorporate external knowledge, which may partly address derivation tasks.
+Similarly, CoTable (Wang et al., 2024) addresses the filtering tasks and derivation tasks by designing operators which are implemented by LLM completion.
+ReAcTable (Zhang et al., 2024) uses few-shot demonstrations to instruct LLMs to generate python code or SQL to address the derivation and filtering tasks.
+However, previous TQA methods have the following limitations when handling data prep.
+First, all these methods do not address the column normalization tasks, which account for the most significant error types, as indicated in Figure 1.
+Second, all these methods utilize a single LLM agent for both data prep and answer reasoning.
+Given that data prep is a complex challenge, these one-size-fits-all solutions still fail to achieve satisfactory performance as demonstrated in our experiments.
+Traditional Data Preparation Data prep techniques are widely used across various tasks, including training machine learning models for data analytics.
+Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify data prep operations.
+Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020) apply meta-learning to discover promising operations.
+Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques.
+Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) explore reinforcement learning techniques.
+HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
+In this paper, we propose AutoPrep to generate question-aware data prep operations for TQA tasks.
+The key difference between AutoPrep and traditional data prep methods is that AutoPrep requires semantic alignment between the table and the NL question.
+Traditional data prep methods typically operate at an offline stage, independent of any specific downstream question or query.
+In contrast, question-aware data prep studied in AutoPrep tailors the tables to the specific needs of the question during the online stage, directly addressing the challenge of aligning the table’s structure with the NL question’s semantics.
+LLMs-based Multi-Agent Framework.
+A multi-agent LLM framework refers to a well-designed hierarchical structure consisting of multiple LLM-based agents and scheduling algorithms (Han et al., 2024).
+Compared with single-agent methods based on prompting techniques, such structures are better suited for handling complex tasks like software development, issue resolution, and code generation (Hong et al., 2023; Qian et al., 2023; Chen et al., 2024; Tao et al., 2024; Ishibashi and Nishimura, 2024; Islam et al., 2024).
+The structure of multi-agent frameworks can be categorized into equi-level (Terekhov et al., 2023), hierarchical (Ahilan and Dayan, 2019; Harris et al., 2023), and nested (Chan et al., 2023) structures.
+AutoTQA (Zhu et al., 2024) proposes a multi-agent framework with hierarchical structure for supporting Tabular Question Answering.
+We adopt the hierarchical structure because our design introduces a Planner agent that decomposes the overall data preparation task into three distinct sub-tasks, each handled by specialized agents.
+This top-down coordination naturally aligns with the principles of the hierarchical multi-agent framework.
+Moreover, while AutoTQA focus on improving table analysis, this paper focuses on the performance bottleneck caused by data prep issues in TQA.
+Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
+8. Conclusion and Future work
+In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework designed to support automatic data prep for TQA tasks.
+AutoPrep consists of three stages: (1) the Planning stage, which suggests logical data prep operations; (2) the Programming stage, which generates physical implementations for each logical operation; and (3) the Executing stage, which executes the Python code and reports error messages.
+We propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation.
+Moreover, we introduce an effective method to ensure successful execution of the physical plan.
+Extensive experiments on WikiTQ, TabFact and TabBench demonstrate the superiority of AutoPrep.
+For future work, we identify three promising directions.
+First, we aim to extend the system’s data prep capabilities by integrating a broader set of operations to handle more complex issues such as missing values, duplicates, etc.
+Second, AutoPrep currently focuses on single-table question answering; extending its capabilities to support multi-table TQA is essential for tackling more realistic and complex scenarios.
+The third direction is to enable question-aware data preparation over enterprise datasets, which are much more complex than existing TQA benchmarks.
+Appendix A Details of the Error Analysis
+This section presents details of the error analysis shown in Figure 1, which motivates us to study question-aware data prep.
+TQA Method: NL2SQL Based on GPT-4.
+We adopt NL2SQL (Rajkumar et al., 2022), a representative TQA baseline that does not perform data prep, as introduced in Section 6.1.
+Specifically, given a natural language question $Q$ and a table $T$, we prompt GPT-4 to generate an SQL query ${\tt SQL}$ using a few-shot in-context learning strategy.
+The query is then executed on $T$ to obtain the final answer $Ans$ for $Q$.
+Error Categorization.
+We recruite three graduate students from our university’s database group to annotate error categories for a randomly sampled set of 500 instances from WikiTQ and TabFact.
+All annotators have a good understanding of the TQA task and are proficient in SQL syntax.
+We instruct the annotators to first identify errors stemming from insufficient data prep, assigning any remaining errors to a default category labeled Other.
+For the data prep-related errors, after careful investigation, we further categorize them into the following three types: (1) Missing Semantics: The table lacks necessary semantic information required to accurately answer the NL question.
+(2) Inconsistent Values: Values within a column are represented inconsistently across different records.
+(3) Irrelevant Columns: The table includes many columns, but only a small subset are actually relevant to the NL question.
+In-depth Analysis of the “Other” Category.
+We report the detailed sub-categories within the “Other” category in Figure 9.
+As shown, for WikiTQ, more than half of the cases stem from issues inherent to the dataset itself.
+Notably, 40% of the errors are due to limitations in the evaluation function, where semantically equivalent but non-identical responses are marked as incorrect, a problem also highlighted in (Cheng et al., 2023).
+Additionally, 20% of the errors arise from formatting issues in the model’s responses, where the generated SQL logic is correct but the output includes extraneous values from irrelevant columns.
+On the other hand, for the TabFact dataset, most errors are attributed to model reasoning failures.
+Over 50% result from the model’s inability to generate logically complete and accurate SQL queries.
+Furthermore, more than 35% of the errors are due to response formatting issues, where the model returns ambiguous values instead of definitive true/false answers.
+Appendix B More Experiments
+B.1. Evaluation on Various Table Sizes
+Exp: Evaluation of AutoPrep on tables with various sizes.
+To further investigate AutoPrep, we analyze its performance across tables of varying sizes.
+While all tables in TabFact are small, we categorize the tables in WikiTQ into Small (fewer than 2048 tokens), Medium (2048 to 4096 tokens), and Large (more than 4096 tokens), resulting in 4040 Small tables, 200 Medium tables, and 104 Large tables.
+The results are reported in Figure 10.
+We observe that all baselines exhibit unstable performance, particularly on larger tables.
+For instance, although CoTable achieves the highest average accuracy among previous SOTA methods, it suffers an accuracy drop of 11.78 when reasoning over large tables.
+Likewise, ReAcTable shows relatively stable performance on medium tables (dropping by 2.15), yet its performance on large tables remains unsatisfactory (dropping by 17.53).
+Figure 10. Accuracy on different size of tables in WikiTQ.
+In contrast, AutoPrep achieves the highest and most stable performance across tables of varying sizes.
+When processing medium and large tables, the accuracy of AutoPrep drops by 2.29 and 3.79.
+The key to its stability is that it employs specialized agents for each data prep task, mitigating the issue of exceeding prompt length limits.
+Furthermore, each agent generates program-based operators to handle data prep on the entire table, thereby avoiding information loss that occurs when large tables are cropped.
+Finding: By employing multiple agents and program-based operations, AutoPrep maintains stable performance as table size grows, ensuring that each data prep task is handled effectively without overwhelming a single model.
+B.2. Evaluation on Various Question Difficulties
+Table 6. In-depth Analysis of AutoPrep and all baselines when varying question difficulties on WikiTQ and TabFact.
+Method WikiTQ TabFact
+Simple Complex Simple Complex
+End2End 57.8057.80{57.80}57.80 55.6855.68{55.68}55.68 $86.51$ $77.07$
+CoT $56.85$ $55.33$ 84.9284.92{84.92}84.92 79.1379.13{79.13}79.13
+NL2Py 61.8761.87{61.87}61.87 57.2157.21{57.21}57.21 $70.83$ $65.45$
+NL2SQL $57.02$ $49.13$ $70.83$ $69.59$
+ICL-Prep 59.5159.51{59.51}59.51 54.0254.02{54.02}54.02 $82.64$ 78.4478.44{78.44}78.44
+Dater $51.33$ $45.77$ $88.39$ $77.76$
+Binder $62.02$ $52.04$ $89.29$ $76.38$
+ReAcTable 66.9866.98{66.98}66.98 61.7161.71{61.71}61.71 90.4890.48{90.48}90.48 79.7279.72{79.72}79.72
+CoTable 66.8866.88{66.88}66.88 62.5362.53{62.53}62.53 90.5890.58{90.58}90.58 81.8981.89{81.89}81.89
+AP + End2End $66.58$ ($+8.78$) $60.23$ ($+4.55$) $\underline{87.20}$ ($+0.69$) $77.07$ ($+0.00$)
++ CoT $64.12$ ($+7.27$) $58.57$ ($+5.24$) $84.52$ ($-0.40$) $80.02$ ($+0.89$)
++ NL2Py $\underline{68.09}$ ($+6.22$) 63.97 ($+6.76$) 90.77 ($+19.94$) $\underline{83.96}$ ($+18.51$)
++ NL2SQL 69.17 ($+11.97$) $\underline{63.50}$ ($+14.37$) 90.77 ($+19.94$) 84.94 ($+15.35$)
+Also, following previous work (Ye et al., 2023; Wang et al., 2024), we compare AutoPrep and all baseline models on questions with different difficulties.
+Exp: How does AutoPrep and other baselines perform on question with different difficulties?
+We divide the TabFact dataset based on officially provided question difficulty labels.
+Since WikiTQ does not provide difficulty labels, we divide the dataset based on question length following previous work (Ye et al., 2023).
+As shown in Table 6, CoTable achieves the best performance among all baselines on both Simple and Complex questions, indicating its abilities of complex reasoning abilities over tables.
+While AutoPrep still outperforms CoTable, especially on Complex questions.
+Specifically, AutoPrep outperforms CoTable by 0.21% on Simple questions and 3.72% on Complex questions.
+Moreover, notice that trivial performance improvement is achieved by AutoPrep for Direct Prompting methods.
+Especially for CoT on Simple questions, the performance of CoT has been dropped by 0.47% when integrated with AutoPrep.
+This is because tables of TabFact are all small tables, data prep operations of AutoPrep fail to bring obvious benefit for Direct Prompting methods.
+While for Code Generation methods, data quality is a critical problem.
+Thus, NL2Code and NL2SQL have achieved obvious performance gains on prepared tables.
+B.3. Impact of LLM-Based $\mathtt{CustomCode}$
+Table 7. Analyzing LLM generated code for logical operator.
+Method  WikiTQ
+$\mathtt{CustomCode}$ ratio ${1.08\%}$
+$\mathtt{CustomCode}$ accuracy ${55.32\%}$
+AutoPrep w. $\mathtt{CustomCode}$ 66.09
+AutoPrep w.o. $\mathtt{CustomCode}$ 65.6365.63{65.63}65.63
+Exp : How does LLM-based $\mathtt{CustomCode}$ help AutoPrep generalize to cases requiring new physical operations?
+As discussed in Section 5.1, we prompt the LLM to write code to implement the logical operation if no suitable function in the function pool can meet the requirement, which we denote as $\mathtt{CustomCode}$.
+We quantify the usage frequency of $\mathtt{CustomCode}$ and evaluate whether $\mathtt{CustomCode}$ produces the correct code on WikiTQ.
+We also evaluate the performance of AutoPrep without $\mathtt{CustomCode}$ design.
+As shown in Table 7, The results show that $\mathtt{CustomCode}$ is invoked in 1.08% of cases, achieving an accuracy of 55.32%.
+Notably, removing $\mathtt{CustomCode}$ leads to a performance drop of 0.46 points (from 66.09 to 65.63).
+These results indicate that, despite its relatively low invocation frequency and moderate accuracy, $\mathtt{CustomCode}$ improves generalization of AutoPrep, particularly in handling new or unforeseen physical operations.
+Finding : $\mathtt{CustomCode}$ improves the generalization of AutoPrep on cases requiring new physical operations.
+B.4. Analysis on Other Data Quality Issues
+We have conducted a new experiment to analyze the impact of missing values and duplication in TQA tasks on the WikiTQ dataset.
+Specifically, for duplication, we first examine all tables containing duplicated records using a heuristic method from (Fan et al., 2024) on the WikiTQ dataset, and only obtain 24 such instances.
+Then, we manually remove the duplicate records and re-evaluate AutoPrep ’s performance on the updated dataset.
+The results remain unchanged (accuracy: 70.83% for both the original and de-duplicated datasets), suggesting that duplication is not a critical issue on existing TQA benchmarking datasets.
+Similarly, for missing values, we prompt LLMs to impute the missing entries based on the available data for all tables and then re-evaluate AutoPrep on the imputated dataset.
+The results also show insignificant difference (65.00% w/o imputation vs.
+64.33% with imputation), indicating that missing values have limited impact.
+These above results suggest that such data preparation issues are uncommon and have little effect on existing TQA benchmarking datasets.
+This is largely due to the way of constructing the datasets.
+Specifically, during dataset construction (Chen et al., 2019; Pasupat and Liang, 2015; Wu et al., 2024), annotators are instructed to first generate natural language questions based on a given table and then derive the corresponding answers directly from the same table.
+This construction paradigm implicitly ensures data completeness and consistency, as annotators tend to avoid crafting questions around missing or duplicating information.
+As a result, common data quality issues such as missing values and duplicate records are naturally excluded from the dataset, making them less relevant in the context of existing TQA tasks.
+Nevertheless, we acknowledge that issues such as missing values and duplicates are common in real-world datasets, we leave the exploration of these broader data prep challenges to future work.
+Finding : While issues such as missing values and duplicates are common in real-world datasets, they are significantly less prevalent than the three highlighted challenges in existing TQA benchmarks.
+Appendix C A Running Example of AUTOPREP
+(a) Case one
+(b) Case two
+Figure 11. Two case studies selected from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
 This section presents two illustrative examples from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
@@
 For ReAcTable, although it generates a logically correct Python program, it ignores the types of existing columns which do not support for numerical calculation.
-Thus, ReAcTable also fails to augment the Tol column.
-AutoPrep addresses this issue by first utilizing a Planner agent to generate logical operators specifying data prep requirements, i.e., two $\mathtt{Normalize}$ to normalize column Expense and Percent, a $\mathtt{Augment}$ to generate the overall expenses Tol and a $\mathtt{Filter}$ to select related columns Year and Tol.
+Thus, ReAcTable also fails to derive the Tol column.
+AutoPrep addresses this issue by first utilizing a Planner agent to generate logical operators specifying data prep requirements, i.e., two $\mathtt{Normalize}$ to normalize column Expense and Percent, a $\mathtt{Derive}$ to generate the overall expenses Tol and a $\mathtt{Filter}$ to select related columns Year and Tol.
 These logical operators are passed to Programmer agents to generate physical operators consisting of a pre-defined Python function.
@@
 Based on this, the Analyzer agent extracts the final answer “8”.
-7. Related Work
-Tabular Question Answering.
-Most of the SOTA solutions for TQA rely on LLMs (Brown et al., 2020; Guo et al., 2024), as TQA requires NL understanding and reasoning over tables.
-Specifically, direct Prompting approaches (Chen, 2022; Wei et al., 2022), which prompt an LLM to directly generate answers through next token prediction, may struggle when processing large tables due to the input token limitation and long-context understanding challenges (Li et al., 2024).
-Code Generation approaches (Rajkumar et al., 2022) leverage LLMs to generate programming code (e.g., Python) that extracts answers from tables, providing better explainability and adaptability.
-Previous TQA methods, like Dater (Ye et al., 2023), Binder (Cheng et al., 2023), CoTable (Wang et al., 2024b) and ReAcTable (Zhang et al., 2024), also consider data prep in their question answering process.
-Specifically, Dater (Ye et al., 2023) prompts LLMs to select relevant columns related to answering the question, targeted at solving filtering tasks.
-Binder (Cheng et al., 2023) proposes to integrate SQL with an LLM-based API to incorporate external knowledge, which may partly address augmentation tasks.
-Similarly, CoTable (Wang et al., 2024b) addresses the filtering tasks and augmentation tasks by designing operators which are implemented by LLM completion.
-ReAcTable (Zhang et al., 2024) uses few-shot demonstrations to instruct LLMs to generate python code or SQL to address the augmentation and filtering tasks.
-However, previous TQA methods have the following limitations when handling data prep.
-First, all these methods do not address the cell-level normalization tasks, which account for the most significant error types, as indicated in Figure 1.
-Second, all these methods utilize a single LLM agent for both data prep and answer reasoning.
-Given that data prep is a complex challenge, these one-size-fits-all solutions still fail to achieve satisfactory performance as demonstrated in our experiments.
-Traditional Data Preparation.
-Data prep techniques have been widely adopted in various tasks.
-Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify data prep operations.
-Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020) apply meta-learning to discover promising operations.
-Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques.
-Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) explore reinforcement learning techniques.
-HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
-In this paper, we propose AutoPrep to generate question-specific data prep operations for TQA tasks.
-The key difference between these methods and AutoPrep is that the data prep operations generated by AutoPrep are tailored to specific user questions and customized for the tables, as different questions may have different data prep requirements, even for the same table.
-LLMs-based Multi-Agent Framework.
-A multi-agent LLM framework refers to a well-designed hierarchical structure consisting of multiple LLM-based agents and scheduling algorithms (Wang et al., 2024a).
-Compared with single-agent methods based on prompting techniques, such structures are better suited for handling complex tasks like software development, issue resolution, and code generation (Hong et al., 2023; Qian et al., 2023; Chen et al., 2024; Tao et al., 2024; Ishibashi and Nishimura, 2024; Islam et al., 2024).
-AutoTQA (Zhu et al., 2024) propose a multi-agent framework supporting Tabular Question Answering.
-While AutoTQA improves table analysis, this paper focuses on the performance bottleneck caused by data prep issues in TQA.
-Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
-8. Conclusion
-In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework designed to support automatic data prep for TQA tasks.
-AutoPrep consists of three stages: (1) the Planning stage, which suggests high-level data prep operations; (2) the Programming stage, which generates low-level implementations for each high-level operation; and (3) the Executing stage, which runs the Python code and report error messages.
-We also propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation.
-Additionally, we introduce an effective method to ensure successful execution of the physical plan.
-Extensive experiments on WikiTQ, TabFact and TabBench demonstrate that: (1) AutoPrep achieves SOTA performance in data prep for TQA, and (2) integrating AutoPrep’s data prep improves the performance of TQA methods.
+Appendix D Prompts in AutoPrep
+The representative demonstration examples and prompts used by AutoPrep are presented.
+D.1. Prompts in Planner Agent
+Figure 12. Prompt of Planner on generating Analysis Sketch on WikiTQ Dataset.
+Figure 13. Prompt of Planner on generating Analysis Sketch on TabFact Dataset.
+Figure 14. Prompt of Planner on generating requirement for Derive Operator.
+Figure 15. Prompt of Planner on generating requirement for Normalize Operator.
+The Planner aims at generating logical operators in two stages: (1) generate the QA-Sketch. (2) generate specific requirement for Normalize and Derive.
+In the first stage, AutoPrep analyzes the table and the question to generate the reasoning process of question answering over table with the assistance of pseudo codes.
+In this paper, we define the grammar of the pseudo code as SQL with a UDF grammar.
+The prompts for WikITQ and TabFact datasets are shown in Figure 12 and Figure 13.
+Next, based on the generated QA-Sketch, we extract each operation clause and provide the related column data in the clause to query LLMs to generate specific requirement of the logical operator.
+The prompts of generating specific requirement for each type of logical operator are shown in Figure 14 and Figure 15.
+Notice that the Filter operator is generated by analyzing the QA-Sketch and the arguments of Derive operator, which does not need to query LLMs.
+D.2. Prompts in Programmer Agents
+Figure 16. Prompt of Derive.
+Figure 17. Prompt of Normalize.
+Figure 18. Prompt of Imputater.
+Figure 19. Prompt of LLM Coder.
+Given logical operators with specific requirement and related columns, the programmer agents generate physical operators to implement the logical operators.
+The prompts of Derive and Normalize are shown in Figure 16 and Figure 17.
+In addition, if any logical operators can not be implemented with functions in the function pool, we call LLMs to generate Python codes.
+The prompt of generating Python code is shown in Figure 19.
+Finally, if any values can not be derived or normalized, we call an Imputater agent to direct infer the correct values.
+The prompt is shown in Figure 18.
+D.3. Prompts in other baselines
+Figure 20. Prompt of End2End method on WikiTQ dataset.
+Figure 21. Prompt of End2End method on TabFact dataset.
+Figure 22. Prompt of Chain-of-Thought TQA method on WikiTQ dataset.
+Figure 23. Prompt of Chain-of-Thought TQA method on TabFact dataset.
+Figure 24. Prompt of NL2SQL on WikiTQ dataset.
+Figure 25. Prompt of NL2SQL on TabFact dataset.
+Figure 26. Prompt of ICL-Prep on WikiTQ dataset.
+Figure 27. Prompt of ICL-Prep on TabFact dataset.
+Figure 28. Prompt of NL2Py on WikiTQ dataset.
+Figure 29. Prompt of NL2Py on TabFact dataset.
+We also provide the prompts of implemented baselines, including End2End (Figure 20 and Figure 21), CoT (Figure 22 and Figure 23), NL2SQL (Figure 24 and Figure 25), NL2Py (Figure 28 and Figure 29) and ICL-Prep (Figure 26 and Figure 27).
+Some of the prompts are from existing works (Cheng et al., 2023; Chen, 2022) as described in the experimental setup.

### v4

@@
 PVLDB Reference Format:
-PVLDB, 14(1): XXX-XXX, 2020.
+PVLDB, 18(10): XXX-XXX, 2025.
 doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
@@
 Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
+Proceedings of the VLDB Endowment, Vol. 18, No. 10 ISSN 2150-8097.
 doi:XX.XX/XXX.XX
 PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at https://github.com/fmh1art/AutoPrep.
+The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/ruc-datalab/AutoPrep.
 1. Introduction
-Tabular Question Answering (TQA) refers to the task of answering natural language (NL) questions based on provided tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023).
-TQA empowers non-technical users such as domain scientists to easily analyze tabular data and has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022).
-As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024) rely on large language models (LLMs).
+Tabular Question Answering (TQA) refers to the task of answering natural language (NL) questions based on provided tables (Lu et al., 2025; Cheng et al., 2022, 2023; Pal et al., 2023).
+TQA empowers non-technical users such as domain scientists to easily analyze tabular data and has a wide range of applications, including table-based fact verification (Chen et al., 2019; Gu et al., 2022a, b) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022).
+As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024) mainly rely on large language models (LLMs).
 Figure 1. An error analysis of LLM-based TQA (using GPT-4) on two well-adopted datasets.
@@
 Thus, column filtering, i.e., retaining only the columns directly related to the question, is crucial for TQA, as irrelevant columns add complexity and can mislead LLMs into producing incorrect answers.
-For instance, the Summary column in $T_{7}$ contains an average of 100 words, increasing the difficulty for LLMs to identify the relevant AirDate column.
 This highlights the critical role of thorough data prep in ensuring accurate TQA.
-Note that, although issues like missing values and duplicates are common in real-world data, they are less prevalent than the three highlighted issues in current TQA benchmarks.
+Note that, although other issues like missing values and duplicates are common in real-world data, they are less prevalent than the three highlighted issues in current TQA benchmarks.
 Nevertheless, we recognize their significance and leave the exploration of broader data prep challenges to future work.
@@
 In the next stage, AutoPrep maps these logical operations to the corresponding physical implementations.
-This separation allows AutoPrep to break down complex data prep tasks into smaller, more manageable sub-tasks, making the process more easier to solve.
+This separation allows AutoPrep to break down complex data prep tasks into smaller, more manageable sub-tasks, making the process easier to solve.
 Moreover, this modular design makes AutoPrep extensible, enabling the development of specialized implementations for each type of data prep operation or the introduction of new operation types.
@@
 Figure 2 provides several examples of TQA.
-For instance, consider table $T_{1}$, which contains medal information for cyclists from different countries, with two columns: Cyclist and Medal.
+Consider table $T_{1}$, which contains medal information for cyclists from different countries, with two columns: Cyclist and Medal.
 Given the question $Q_{1}$, “Which country has the most medals?”, the answer should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”.
-TQA often requires complex reasoning over tables.
+TQA often requires reasoning over tables.
 For instance, to answer question $Q_{2}$, we first need to calculate the “GDP growth rate” for all countries, then sort the countries by growth rate, and finally identify the country with the highest GDP growth rate, i.e., CHN.
@@
 (a) The Derive operation: Figure 2(a) shows three examples of the Derive operations, i.e., extracting Country information from column Cyclist in $T_{1}$, computing new column GrowthRate using two columns in $T_{2}$, and inferring a status of IfLost by analyzing the scores in column Result.
-Note that the corresponding Python implementations are provided alongside the operations.
 (b) The Normalize operation: Figure 2(b) shows three examples of Normalize operations, i.e., normalizing the formats of Surface, Time and Date for tables $T_{4}$, $T_{5}$ and $T_{6}$ respectively.
@@
 First, at the logical-operation level, given the inherent difficulties in understanding both NL questions and tables, it is challenging to accurately identify which data prep operations are specifically required to satisfy the needs of the NL question.
-This often leads to false negatives (FNs) and false positives (FPs).
-For example, as shown in Figure 3, converting the Age column to a numerical format in code block ② is a false positive (FP), as it is irrelevant to the question.
-In contrast, failing to normalize the Date column before extracting the month in code block ③ constitutes a false negative (FN), as the method ignores the inconsistency in Date formats.
+This often leads to false negatives and false positives.
+For example, as shown in Figure 3, converting the Age column to a numerical format in code block ② is a false positive, as it is irrelevant to the question.
+In contrast, failing to normalize the Date column before extracting the month in code block ③ constitutes a false negative, as the method ignores the inconsistency in Date formats.
 Second, at the physical-operation level, due to input token limitations and challenges in long-context understanding (Li et al., 2024), it is not easy to fully understand all possible issues in a table, and thus may struggle to generate customized programs to correct issues.
@@
   2. (2)
-    Multiple Programmer Agents (i.e., Normalize): the Programming stage.
+    Multiple Programmer Agents (e.g., Normalize): the Programming stage.
 It directs the LLM to generate physical implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$.
@@
   • Normalize(“Case to INT”, Medal) that standardizes the value formats in the Medal column (e.g., removing quotation marks and asterisks) and then converts the strings to integers, as the question requires “the most medals”;
-  • Normalize(“Format date as %m-%d”, Date) standardizes the values in the Date column into a unified format to support the ‘in February’ condition in the question.
+  • Normalize(“Format date as %m-%d”, Date) standardizes the values in the Date column into a unified format to support the “in February” condition in the question.
   • $T^{\prime}$ $=$ Filter([Date, Country, Medal], $T$) that filters out column Age, which is irrelevant to the question;
-(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e.,, Derive, Normalize and Filter.
+(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e., Derive, Normalize and Filter.
 Each specialized Programmer focuses on generating executable code for its assigned operations.
@@
 The key challenge is how to suggest a logical plan that address specific NL questions.
-Specifically, even for the same table, different NL questions may require not only different logical operations but also varying sequences of those operations.
+Even for the same table, different NL questions may require not only different logical operations but also varying sequences of those operations.
 To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent.
 This method translates the NL question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch.
-Compared to existing CoT methods (Wei et al., 2022), which break down questions into sub-questions, our approach better captures the relationships between questions and tables.
 More details of the method are given in Section 4.
@@
 This mistake is critical, as the country information is implicitly embedded within the Cyclist values.
-(b) Missing operations: Observing the ground-truth in Figure 4, we see that the Derive operation on Cyclist is not generated, as the column has already been filtered out.
+(b) Missing operations: Observing the ground-truth in Figure 4, we can see that the Derive operation on Cyclist is not generated, as the column has already been filtered out.
 Moreover, although the Normalize operation on Medal is generated, the Normalize operation on Date is missing.
@@
 Figure 5. Our proposed approaches to logical plan generation in the Planner agent: (a) A straightforward direct prompting method, and (b) a more effective Chain-of-Clauses (CoC) reasoning method.
-Key Challenge.
 The above example clearly demonstrates that the key challenge in designing the Planner agent, as discussed earlier, is to capture the relationships between different parts of question $Q$ and the columns in table $T$.
@@
 (3) $\mathtt{clean\mbox{-}string}$.
-To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$.
-It has two arguments: ${\tt column}$ and ${\tt trans\_dict}$.
-The $\tt trans\_dict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$.
-(4) $\mathtt{infer}$. Similar to the one for the $\mathtt{Derive}$ operation, this function uses LLMs to deduce values that could not be processed by the functions above.
+To clean string values in tables, we define function $\mathtt{clean\mbox{-}string}$ with two arguments: ${\tt column}$ and ${\tt trans\_dict}$.
+The $\tt trans\_dict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ is replaced by $v$.
+(4) $\mathtt{infer}$. This function uses LLMs to deduce values that could not be processed by the functions above.
 Function Pool for Filter.
@@
 Moreover, End2End and CoT methods also show considerable performance gains.
-These improvements are largely due to AutoPrep ’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers from tables.
+These improvements are largely due to the filtering mechanism of AutoPrep, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers from tables.
 Since “NL2SQL + AutoPrep” achieves the best accuracy in most cases, we take its results as default for further comparison.
@@
 To further evaluate the extensibility of AutoPrep, we evaluate it on a more complex dataset involving logical operations not originally supported.
-Since no existing TQA dataset presents sufficiently complex data quality issues, we construct a new dataset—TransTQ with the assistance of LLMs, which is available on our GitHub repository.
+Since no existing TQA dataset presents sufficiently complex data quality issues, we construct a new dataset TransTQ with the assistance of LLMs, which is available on our GitHub repository.
 Specifically, we prompt an LLM to identify tables from WikiTQ where inverse transformation operations (e.g., pivot and stack (Li et al., 2023)) can be applied.
@@
 ReAcTable (Zhang et al., 2024) uses few-shot demonstrations to instruct LLMs to generate python code or SQL to address the derivation and filtering tasks.
-However, previous TQA methods have the following limitations when handling data prep.
-First, all these methods do not address the column normalization tasks, which account for the most significant error types, as indicated in Figure 1.
-Second, all these methods utilize a single LLM agent for both data prep and answer reasoning.
-Given that data prep is a complex challenge, these one-size-fits-all solutions still fail to achieve satisfactory performance as demonstrated in our experiments.
-Traditional Data Preparation Data prep techniques are widely used across various tasks, including training machine learning models for data analytics.
-Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify data prep operations.
+However, previous TQA methods do not address the column normalization tasks, which account for the most significant error types, as indicated in Figure 1.
+Second, the methods utilize a single LLM agent for both data prep and answer reasoning.
+Given that data prep is a complex challenge, these one-size-fits-all solutions may not achieve satisfactory performance.
+Traditional Data Preparation.
+Data prep techniques are widely used across various tasks (Chai et al., 2023; Fan et al., 2024; Tu et al., 2022).
+For training machine learning models for data analytics, Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify data prep operations.
 Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020) apply meta-learning to discover promising operations.
 Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques.
-Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) explore reinforcement learning techniques.
+Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020; Zhang et al., 2025) explore reinforcement learning techniques.
 HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
-In this paper, we propose AutoPrep to generate question-aware data prep operations for TQA tasks.
-The key difference between AutoPrep and traditional data prep methods is that AutoPrep requires semantic alignment between the table and the NL question.
+We propose AutoPrep to generate question-aware data prep operations for TQA tasks.
 Traditional data prep methods typically operate at an offline stage, independent of any specific downstream question or query.
@@
 This top-down coordination naturally aligns with the principles of the hierarchical multi-agent framework.
-Moreover, while AutoTQA focus on improving table analysis, this paper focuses on the performance bottleneck caused by data prep issues in TQA.
+Moreover, while AutoTQA focuses on improving table analysis, this paper focuses on the performance bottleneck caused by data prep issues in TQA.
 Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
 8. Conclusion and Future work
-In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework designed to support automatic data prep for TQA tasks.
-AutoPrep consists of three stages: (1) the Planning stage, which suggests logical data prep operations; (2) the Programming stage, which generates physical implementations for each logical operation; and (3) the Executing stage, which executes the Python code and reports error messages.
+In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework to support data prep for TQA tasks.
+AutoPrep consists of three stages: (1) the Planning stage, which suggests logical data prep operations, (2) the Programming stage, which generates physical implementations for each logical operation, and (3) the Executing stage, which executes the Python code and reports error messages.
 We propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation.
-Moreover, we introduce an effective method to ensure successful execution of the physical plan.
-Extensive experiments on WikiTQ, TabFact and TabBench demonstrate the superiority of AutoPrep.
+Extensive experiments on real datasets demonstrate the superiority of AutoPrep.
 For future work, we identify three promising directions.
 First, we aim to extend the system’s data prep capabilities by integrating a broader set of operations to handle more complex issues such as missing values, duplicates, etc.
-Second, AutoPrep currently focuses on single-table question answering; extending its capabilities to support multi-table TQA is essential for tackling more realistic and complex scenarios.
-The third direction is to enable question-aware data preparation over enterprise datasets, which are much more complex than existing TQA benchmarks.
-Appendix A Details of the Error Analysis
-This section presents details of the error analysis shown in Figure 1, which motivates us to study question-aware data prep.
-TQA Method: NL2SQL Based on GPT-4.
-We adopt NL2SQL (Rajkumar et al., 2022), a representative TQA baseline that does not perform data prep, as introduced in Section 6.1.
-Specifically, given a natural language question $Q$ and a table $T$, we prompt GPT-4 to generate an SQL query ${\tt SQL}$ using a few-shot in-context learning strategy.
-The query is then executed on $T$ to obtain the final answer $Ans$ for $Q$.
-Error Categorization.
-We recruite three graduate students from our university’s database group to annotate error categories for a randomly sampled set of 500 instances from WikiTQ and TabFact.
-All annotators have a good understanding of the TQA task and are proficient in SQL syntax.
-We instruct the annotators to first identify errors stemming from insufficient data prep, assigning any remaining errors to a default category labeled Other.
-For the data prep-related errors, after careful investigation, we further categorize them into the following three types: (1) Missing Semantics: The table lacks necessary semantic information required to accurately answer the NL question.
-(2) Inconsistent Values: Values within a column are represented inconsistently across different records.
-(3) Irrelevant Columns: The table includes many columns, but only a small subset are actually relevant to the NL question.
-In-depth Analysis of the “Other” Category.
-We report the detailed sub-categories within the “Other” category in Figure 9.
-As shown, for WikiTQ, more than half of the cases stem from issues inherent to the dataset itself.
-Notably, 40% of the errors are due to limitations in the evaluation function, where semantically equivalent but non-identical responses are marked as incorrect, a problem also highlighted in (Cheng et al., 2023).
-Additionally, 20% of the errors arise from formatting issues in the model’s responses, where the generated SQL logic is correct but the output includes extraneous values from irrelevant columns.
-On the other hand, for the TabFact dataset, most errors are attributed to model reasoning failures.
-Over 50% result from the model’s inability to generate logically complete and accurate SQL queries.
-Furthermore, more than 35% of the errors are due to response formatting issues, where the model returns ambiguous values instead of definitive true/false answers.
-Appendix B More Experiments
-B.1. Evaluation on Various Table Sizes
-Exp: Evaluation of AutoPrep on tables with various sizes.
-To further investigate AutoPrep, we analyze its performance across tables of varying sizes.
-While all tables in TabFact are small, we categorize the tables in WikiTQ into Small (fewer than 2048 tokens), Medium (2048 to 4096 tokens), and Large (more than 4096 tokens), resulting in 4040 Small tables, 200 Medium tables, and 104 Large tables.
-The results are reported in Figure 10.
-We observe that all baselines exhibit unstable performance, particularly on larger tables.
-For instance, although CoTable achieves the highest average accuracy among previous SOTA methods, it suffers an accuracy drop of 11.78 when reasoning over large tables.
-Likewise, ReAcTable shows relatively stable performance on medium tables (dropping by 2.15), yet its performance on large tables remains unsatisfactory (dropping by 17.53).
-Figure 10. Accuracy on different size of tables in WikiTQ.
-In contrast, AutoPrep achieves the highest and most stable performance across tables of varying sizes.
-When processing medium and large tables, the accuracy of AutoPrep drops by 2.29 and 3.79.
-The key to its stability is that it employs specialized agents for each data prep task, mitigating the issue of exceeding prompt length limits.
-Furthermore, each agent generates program-based operators to handle data prep on the entire table, thereby avoiding information loss that occurs when large tables are cropped.
-Finding: By employing multiple agents and program-based operations, AutoPrep maintains stable performance as table size grows, ensuring that each data prep task is handled effectively without overwhelming a single model.
-B.2. Evaluation on Various Question Difficulties
-Table 6. In-depth Analysis of AutoPrep and all baselines when varying question difficulties on WikiTQ and TabFact.
-Method WikiTQ TabFact
-Simple Complex Simple Complex
-End2End 57.8057.80{57.80}57.80 55.6855.68{55.68}55.68 $86.51$ $77.07$
-CoT $56.85$ $55.33$ 84.9284.92{84.92}84.92 79.1379.13{79.13}79.13
-NL2Py 61.8761.87{61.87}61.87 57.2157.21{57.21}57.21 $70.83$ $65.45$
-NL2SQL $57.02$ $49.13$ $70.83$ $69.59$
-ICL-Prep 59.5159.51{59.51}59.51 54.0254.02{54.02}54.02 $82.64$ 78.4478.44{78.44}78.44
-Dater $51.33$ $45.77$ $88.39$ $77.76$
-Binder $62.02$ $52.04$ $89.29$ $76.38$
-ReAcTable 66.9866.98{66.98}66.98 61.7161.71{61.71}61.71 90.4890.48{90.48}90.48 79.7279.72{79.72}79.72
-CoTable 66.8866.88{66.88}66.88 62.5362.53{62.53}62.53 90.5890.58{90.58}90.58 81.8981.89{81.89}81.89
-AP + End2End $66.58$ ($+8.78$) $60.23$ ($+4.55$) $\underline{87.20}$ ($+0.69$) $77.07$ ($+0.00$)
-+ CoT $64.12$ ($+7.27$) $58.57$ ($+5.24$) $84.52$ ($-0.40$) $80.02$ ($+0.89$)
-+ NL2Py $\underline{68.09}$ ($+6.22$) 63.97 ($+6.76$) 90.77 ($+19.94$) $\underline{83.96}$ ($+18.51$)
-+ NL2SQL 69.17 ($+11.97$) $\underline{63.50}$ ($+14.37$) 90.77 ($+19.94$) 84.94 ($+15.35$)
-Also, following previous work (Ye et al., 2023; Wang et al., 2024), we compare AutoPrep and all baseline models on questions with different difficulties.
-Exp: How does AutoPrep and other baselines perform on question with different difficulties?
-We divide the TabFact dataset based on officially provided question difficulty labels.
-Since WikiTQ does not provide difficulty labels, we divide the dataset based on question length following previous work (Ye et al., 2023).
-As shown in Table 6, CoTable achieves the best performance among all baselines on both Simple and Complex questions, indicating its abilities of complex reasoning abilities over tables.
-While AutoPrep still outperforms CoTable, especially on Complex questions.
-Specifically, AutoPrep outperforms CoTable by 0.21% on Simple questions and 3.72% on Complex questions.
-Moreover, notice that trivial performance improvement is achieved by AutoPrep for Direct Prompting methods.
-Especially for CoT on Simple questions, the performance of CoT has been dropped by 0.47% when integrated with AutoPrep.
-This is because tables of TabFact are all small tables, data prep operations of AutoPrep fail to bring obvious benefit for Direct Prompting methods.
-While for Code Generation methods, data quality is a critical problem.
-Thus, NL2Code and NL2SQL have achieved obvious performance gains on prepared tables.
-B.3. Impact of LLM-Based $\mathtt{CustomCode}$
-Table 7. Analyzing LLM generated code for logical operator.
-Method  WikiTQ
-$\mathtt{CustomCode}$ ratio ${1.08\%}$
-$\mathtt{CustomCode}$ accuracy ${55.32\%}$
-AutoPrep w. $\mathtt{CustomCode}$ 66.09
-AutoPrep w.o. $\mathtt{CustomCode}$ 65.6365.63{65.63}65.63
-Exp : How does LLM-based $\mathtt{CustomCode}$ help AutoPrep generalize to cases requiring new physical operations?
-As discussed in Section 5.1, we prompt the LLM to write code to implement the logical operation if no suitable function in the function pool can meet the requirement, which we denote as $\mathtt{CustomCode}$.
-We quantify the usage frequency of $\mathtt{CustomCode}$ and evaluate whether $\mathtt{CustomCode}$ produces the correct code on WikiTQ.
-We also evaluate the performance of AutoPrep without $\mathtt{CustomCode}$ design.
-As shown in Table 7, The results show that $\mathtt{CustomCode}$ is invoked in 1.08% of cases, achieving an accuracy of 55.32%.
-Notably, removing $\mathtt{CustomCode}$ leads to a performance drop of 0.46 points (from 66.09 to 65.63).
-These results indicate that, despite its relatively low invocation frequency and moderate accuracy, $\mathtt{CustomCode}$ improves generalization of AutoPrep, particularly in handling new or unforeseen physical operations.
-Finding : $\mathtt{CustomCode}$ improves the generalization of AutoPrep on cases requiring new physical operations.
-B.4. Analysis on Other Data Quality Issues
-We have conducted a new experiment to analyze the impact of missing values and duplication in TQA tasks on the WikiTQ dataset.
-Specifically, for duplication, we first examine all tables containing duplicated records using a heuristic method from (Fan et al., 2024) on the WikiTQ dataset, and only obtain 24 such instances.
-Then, we manually remove the duplicate records and re-evaluate AutoPrep ’s performance on the updated dataset.
-The results remain unchanged (accuracy: 70.83% for both the original and de-duplicated datasets), suggesting that duplication is not a critical issue on existing TQA benchmarking datasets.
-Similarly, for missing values, we prompt LLMs to impute the missing entries based on the available data for all tables and then re-evaluate AutoPrep on the imputated dataset.
-The results also show insignificant difference (65.00% w/o imputation vs.
-64.33% with imputation), indicating that missing values have limited impact.
-These above results suggest that such data preparation issues are uncommon and have little effect on existing TQA benchmarking datasets.
-This is largely due to the way of constructing the datasets.
-Specifically, during dataset construction (Chen et al., 2019; Pasupat and Liang, 2015; Wu et al., 2024), annotators are instructed to first generate natural language questions based on a given table and then derive the corresponding answers directly from the same table.
-This construction paradigm implicitly ensures data completeness and consistency, as annotators tend to avoid crafting questions around missing or duplicating information.
-As a result, common data quality issues such as missing values and duplicate records are naturally excluded from the dataset, making them less relevant in the context of existing TQA tasks.
-Nevertheless, we acknowledge that issues such as missing values and duplicates are common in real-world datasets, we leave the exploration of these broader data prep challenges to future work.
-Finding : While issues such as missing values and duplicates are common in real-world datasets, they are significantly less prevalent than the three highlighted challenges in existing TQA benchmarks.
-Appendix C A Running Example of AUTOPREP
-(a) Case one
-(b) Case two
-Figure 11. Two case studies selected from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
-This section presents two illustrative examples from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
-To answer the question in Figure 11(a), CoTable generates an operator chain with an “add_col” operator, which is used to generate a new column Tol based on the division results of two existing columns Expense and Percent.
-This operator directly prompts an LLM to output a list containing all values of the new column, which is a challenging task for previous LLMs.
-Thus, CoTable generates a new column with two wrong values which leads to a wrong answer.
-For ReAcTable, although it generates a logically correct Python program, it ignores the types of existing columns which do not support for numerical calculation.
-Thus, ReAcTable also fails to derive the Tol column.
-AutoPrep addresses this issue by first utilizing a Planner agent to generate logical operators specifying data prep requirements, i.e., two $\mathtt{Normalize}$ to normalize column Expense and Percent, a $\mathtt{Derive}$ to generate the overall expenses Tol and a $\mathtt{Filter}$ to select related columns Year and Tol.
-These logical operators are passed to Programmer agents to generate physical operators consisting of a pre-defined Python function.
-For example, we call $\mathtt{to\mbox{-}numerical}$ function two times to cast Expense and Percent to numerical values.
-All physical operators are executed to process the original table and output a prepared table.
-Finally, an Analyzer agent based on NL2SQL method generates a SQL query to extract the correct answer “4214.17”.
-Figure 11(b) illustrates a TQA instance involving large tables and inconsistent values.
-AutoPrep solves this by splitting the TQA question into two phases, i.e., data prep and data analysis.
-It first generates logical operators including two $\mathtt{Normalize}$ and one $\mathtt{Filter}$.
-Next, the Programmer agents implement them with corresponding physical operators, which are executed to produce a prepared table.
-Based on this, the Analyzer agent extracts the final answer “8”.
-Appendix D Prompts in AutoPrep
-The representative demonstration examples and prompts used by AutoPrep are presented.
-D.1. Prompts in Planner Agent
-Figure 12. Prompt of Planner on generating Analysis Sketch on WikiTQ Dataset.
-Figure 13. Prompt of Planner on generating Analysis Sketch on TabFact Dataset.
-Figure 14. Prompt of Planner on generating requirement for Derive Operator.
-Figure 15. Prompt of Planner on generating requirement for Normalize Operator.
-The Planner aims at generating logical operators in two stages: (1) generate the QA-Sketch. (2) generate specific requirement for Normalize and Derive.
-In the first stage, AutoPrep analyzes the table and the question to generate the reasoning process of question answering over table with the assistance of pseudo codes.
-In this paper, we define the grammar of the pseudo code as SQL with a UDF grammar.
-The prompts for WikITQ and TabFact datasets are shown in Figure 12 and Figure 13.
-Next, based on the generated QA-Sketch, we extract each operation clause and provide the related column data in the clause to query LLMs to generate specific requirement of the logical operator.
-The prompts of generating specific requirement for each type of logical operator are shown in Figure 14 and Figure 15.
-Notice that the Filter operator is generated by analyzing the QA-Sketch and the arguments of Derive operator, which does not need to query LLMs.
-D.2. Prompts in Programmer Agents
-Figure 16. Prompt of Derive.
-Figure 17. Prompt of Normalize.
-Figure 18. Prompt of Imputater.
-Figure 19. Prompt of LLM Coder.
-Given logical operators with specific requirement and related columns, the programmer agents generate physical operators to implement the logical operators.
-The prompts of Derive and Normalize are shown in Figure 16 and Figure 17.
-In addition, if any logical operators can not be implemented with functions in the function pool, we call LLMs to generate Python codes.
-The prompt of generating Python code is shown in Figure 19.
-Finally, if any values can not be derived or normalized, we call an Imputater agent to direct infer the correct values.
-The prompt is shown in Figure 18.
-D.3. Prompts in other baselines
-Figure 20. Prompt of End2End method on WikiTQ dataset.
-Figure 21. Prompt of End2End method on TabFact dataset.
-Figure 22. Prompt of Chain-of-Thought TQA method on WikiTQ dataset.
-Figure 23. Prompt of Chain-of-Thought TQA method on TabFact dataset.
-Figure 24. Prompt of NL2SQL on WikiTQ dataset.
-Figure 25. Prompt of NL2SQL on TabFact dataset.
-Figure 26. Prompt of ICL-Prep on WikiTQ dataset.
-Figure 27. Prompt of ICL-Prep on TabFact dataset.
-Figure 28. Prompt of NL2Py on WikiTQ dataset.
-Figure 29. Prompt of NL2Py on TabFact dataset.
-We also provide the prompts of implemented baselines, including End2End (Figure 20 and Figure 21), CoT (Figure 22 and Figure 23), NL2SQL (Figure 24 and Figure 25), NL2Py (Figure 28 and Figure 29) and ICL-Prep (Figure 26 and Figure 27).
-Some of the prompts are from existing works (Cheng et al., 2023; Chen, 2022) as described in the experimental setup.
+Second, AutoPrep currently focuses on single-table question answering, and extending its capabilities to support multi-table TQA is essential for tackling more realistic and complex scenarios.
+The third direction is to enable question-aware data preparation over enterprise datasets, which are more complex than existing TQA benchmarks.
+Acknowledgment
+This paper was partly supported by the NSF of China (62436010, 62441230, 62525202 and 62232009), National Key R&D Program of China (2023YFB4503600), NSF DBI-2327954, Guangdong Provincial Project 2023CX10X008 and Amazon Research Awards.